### v1

# Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis

Abstract.
Animal pose estimation (APE) aims to locate the animal body parts using a diverse array of sensor and modality inputs, which is crucial for research across neuroscience, biomechanics, and veterinary medicine. By evaluating 178 papers since 2013, APE methods are categorised by sensor and modality types, learning paradigms, experimental setup, and application domains, presenting detailed analyses of current trends, challenges, and future directions in single- and multi-modality APE systems. The analysis also highlights the transition between human and animal pose estimation. Additionally, 2D and 3D APE datasets and evaluation metrics based on different sensors and modalities are provided. A regularly updated project page is provided in this GitHub link.
Survey of animal pose estimation, human pose estimation, 2D and 3D pose estimation, unimodal and multi-modal pose estimation, deep learning-based pose estimation, pose estimation datasets, pose estimation metrics
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; 2024; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: General and reference Surveys and overviews
1. Introduction
Animal pose estimation (APE) identifies the spatial configuration of an animal’s body parts using a diverse array of sensor and modality inputs, which is important in numerous disciplines (Karashchuk et al., 2021). For instance, neuroscientists rely on accurate measurements of animal movement to correlate with brain dynamics (Mathis et al., 2018; Krakauer et al., 2017; Pereira et al., 2020; Luxem et al., 2022). Biomechanists analyse the movement of specific animal body structures to comprehend biomechanical phenomena, such as mechanical properties, stress-strain relationships, and functionality of different body structures (Patel and Braae, 2013; Joska et al., 2021; Patel and Braae, 2014). In addition, ecologists monitor wildlife behaviours to develop more effective conservation strategies. Similarly, ethologists examine detailed animal gestures and postures to study intra- and inter-species communication and societal structures (Zuffi et al., 2019a; Ng et al., 2022a). Moreover, veterinary practitioners employ precise APE techniques to diagnose health conditions (Farahnakian et al., 2021; Hu et al., 2023) and assess recovery post-treatment, thereby mitigating potential disorders (Fang et al., 2021).
In this paper, we analyse the current state of deep learning-based APE in both single- and multi-modality approaches. We discuss the developments in single-sensor and single-modality methods, explore the challenges they face, and subsequently study the potential of multi-sensor and multi-modal systems to overcome these limitations and further enhance the accuracy and robustness of APE. To provide context, we first overview the extensive research conducted in Human Pose Estimation (HPE).
1.1. Human Pose Estimation
The advancements in HPE provides insights to develop the APE methods due to the skeleton similarity between humans and animals. Therefore, we first provide an overview of HPE to help contextualize the current state and future potential of APE.
The field of HPE has rapidly evolved from manual feature engineering to automated feature learning over the past decade, due to the developments in deep learning approaches. Modern HPE methods are mainly RGB-based and predominantly employ convolutional neural networks (CNNs). In the early stages, models like VGG (Simonyan and Zisserman, 2014), ResNet-50 (He et al., 2016), ResNet-101 (He et al., 2016), and ResNet-152 (He et al., 2016) were employed as backbones for HPE from images or videos. The stacked hourglass network (Newell et al., 2016) then presented a novel architecture characterized by repeated bottom-up and top-down processing. This design aimed to capture and consolidate features at multiple resolutions, enabling precise localization of keypoints across scales. Subsequent innovations, such as the High-Resolution Network (HRNet) (Wang et al., 2020) and its variants (Cheng et al., 2020; Yu et al., 2021a), have further advanced spatial information preservation by maintaining high-resolution representations throughout the entire network. Recently, transformer-based methods (Yang et al., 2021; Yuan et al., 2021) have integrated attention mechanisms to capture long-range dependencies and spatial correlations among joints and appearances, expanding the possibilities of HPE. Beyond vision-based HPE, alternative sensors like inertial measurement units (IMUs) (Huang et al., 2020), millimeter-wave (mmWave) radar (Sengupta et al., 2020; Xue et al., 2021; Xie et al., 2023) and thermal cameras (Smith et al., 2023; Chen et al., 2020) have been explored, tailoring HPE to diverse use cases. Additionally, multi-sensor and multi-modal datasets (Ma et al., 2024; Yan et al., 2023; An et al., 2022) and approaches (Fürst et al., 2021; Zheng et al., 2022; An et al., 2022) have emerged as an innovative step over unimodal HPE to overcome individual sensor and modality limitations, offering a synergistic fusion of data that promises increased robustness and precision, especially in scenarios marked by occlusions, changing lighting conditions, and dynamic environments.
1.2. Definition of Animal Pose Estimation
APE, as discussed in this review, is defined as the task of identifying and localizing body joints of terrestrial mammals using a diverse array of sensor inputs. This task results in the reconstruction of the animal’s pose, depicted as either a skeletal representation marked by keypoints (LABEL:fig:Skeleton_Pose_Representations) or as a comprehensive body mesh (LABEL:fig:Body_Mesh_Reconstructions).
Figure 2. Transfering human pose estimation knowledge to proximal animal classes (e.g. Chimpanzees) for their pose estimation (Sanakoyeu et al., 2020).
[…]
Chiampanzee and Human
Exploring APE through various sensory and modality inputs is important as animals may reside in diverse habitats which require different sensors for data collection. Most current APE methodologies are RGB-based, which predict the articulated joint positions from images or videos. However, RGB-based methods fail in many scenarios, such as nocturnal monitoring. For example, rodents are more active in the dark, and standard RGB cameras become inadequate due to poor lighting conditions. Consequently, infrared sensors have emerged as suitable tools for capturing the nocturnal activities of animals (Wang et al., 2018; Patel et al., 2023). Additionally, mmWave radar can detect movements but suffers from the scarcity of its generated point clouds; LiDAR adds value by providing depth information. Inertial sensors (Patel et al., 2017) can also track movement and orientation, offering invaluable insights but requiring on-animal attachment. In specific scenarios, even auditory cues (Li et al., 2022) can indicate animal movement. This paper also explores the potential of using multiple sensors (i.e. RGB + acoustic) to improve APE performance.
There are two representations for the final estimated animal poses regardless of our sensors or modalities of the APE input data. The first encompasses either two-dimensional (2D) or three-dimensional (3D) keypoint pose representations (Karashchuk et al., 2021; Li and Lee, 2023; Nath et al., 2019; Desai et al., 2023) (LABEL:fig:Skeleton_Pose_Representations). In this category, the animal’s anatomy is represented as a tree of interconnected keypoints. These keypoints present animal body joints, and their interconnections reflect the natural relationship between these joints, providing a skeletal representation of the animal’s posture. The second method, body mesh reconstruction (Zuffi et al., 2017; Rüegg et al., 2022; Wu et al., 2023b, a; Yang et al., 2022a), presents a 3D view of the animal’s morphology (LABEL:fig:Body_Mesh_Reconstructions). Within this framework, we differentiate between model-based and model-free reconstructions. The model-based technique (e.g. SMAL (Zuffi et al., 2017)) utilizes supervision from the predefined template shape of the creature learned from 3D toy animal scans. This template is then refined by optimizing the shape and pose parameters. In particular, the shape parameters capture variations in body metrics (e.g. proportions, height, and weight), while the pose parameters address the observable deformations due to movement. The model-based approach can generate detailed models, but collecting the requisite data for new types of animals can be time-intensive, restricting their applicability to rare species. In contrast, the model-free approach operates without any supervision from the predefined template shape of the creature (e.g. MagicPony (Wu et al., 2023b), DOVE (Wu et al., 2023a) and 3D-Fauna (Li et al., 2024b)), which are more capable when estimating the animal poses for the diverse animal species.
The transition from HPE to APE builds on HPE frameworks. Due to skeletal similarities between humans and certain animals like quadrupeds, HPE models for skeletal keypoint detection (Wang et al., 2020; Insafutdinov et al., 2016; Deb et al., 2023) and body mesh reconstruction (Loper et al., 2023) can be pertinent to APE, with modifications to cater to distinct anatomical and dynamic variations across animal species (Li and Lee, 2023; Rashid et al., 2017; Zuffi et al., 2017; Rüegg et al., 2022). One notable example is (Sanakoyeu et al., 2020), which involves transferring human DensePose knowledge to proximal animal classes (fig. 2). This approach leverages the DensePose-COCO dataset, which contains dense human body annotations, and adapts it to animal species like chimpanzees. The transfer is facilitated by establishing correspondences between the 3D models of humans and animals using semantic alignment, so that the learned dense pose mappings for humans can be applied to the animal class. This enables pose extraction without dense animal-specific annotations. In addition, innovations in APE that allow adaptation to unseen animals and environmental changes can also improve the robustness of HPE, particularly in challenging scenarios like emergency response, where the surrounding environment changes rapidly. However, bridging these two fields requires adaptation of knowledge and consideration of the distinct challenges and ethical issues in human and animal studies, which we will discuss next.
1.3. Challenges in Human Pose Estimation and Animal Pose Estimation
HPE and APE encounter several common challenges that can impact the accuracy of associated algorithms. For example, humans and animals may appear differently due to their age, body composition, and clothing. Furthermore, we can see that a subject’s perceived pose, whether human or animal, can significantly alter depending on the viewpoint. Another shared challenge is occlusion. Self-occlusions always exist, and certain body parts may be obscured by external objects (e.g. cars for humans, trees for wild animals). Environmental factors can be another factor. For example, on a rainy day, it is harder to identify the humans or animals’ poses in an outdoor scenario, this further introduces ambiguities. Lastly, a busy or cluttered background can make separating the subject from its environment difficult.
There are also unique challenges in APE. The animal kingdom has a much greater morphological diversity and inter/intra-species variations than humans. Different animal species exhibit significant domain discrepancies in physical characteristics, body structures, and postures. Even within the same species, there can be a large variety in body size, shape, texture, motion, and posture, presenting difficulties in creating universally applicable models (Rüegg et al., 2022). For example, the pelage of porcupines can vary a lot, and this is often more pronounced and thicker in animals than in humans’ hair, which can make tracking skeleton segments tricky. Furthermore, many animals exhibit natural camouflage, travel in herds, interact frequently, or share similar appearances within their species (Zuffi et al., 2019a). These characteristics make it hard to distinguish individual animals and accurately estimate their poses as they blend into the background or appear similar to one another.
Environmental factors also play a pronounced role. Vegetation or terrain features in the natural environments may obstruct the view of animals. One example can be found in the AcinoSet dataset (Joska et al., 2021), where the cheetahs’ paws were excluded as the grass occluded them in most of the videos, making it difficult to estimate animal poses accurately. Moreover, changing lighting conditions can also heavily affect an animal’s appearance, especially when the sun’s position is moving in outdoor environments.
There are also many observational constraints in APE. For example, we often need to observe wildlife from afar; if we approach animals too closely, it may disturb their natural behaviour. This distant viewing distance may lead to reduced resolution and detail. In addition, animals can exhibit rapid and unpredictable movements (e.g. the high-speed pursuit of a cheetah); thus, we may need a stable camera platform and high-speed capture techniques to avoid camera shake and motion blur. Also, nocturnal species may require specialized sensors like infrared or thermal sensors to capture their movements and poses.
Moreover, there are very limited annotated APE datasets (Yu et al., 2021b). Unlike HPE, which predominantly leans on supervised (Wang et al., 2020; Cheng et al., 2020) and semi-supervised learning methodologies (Ukita and Uematsu, 2018; Pavllo et al., 2019; Mitra et al., 2020) sustained by large-scale human pose datasets such as MSCOCO (Lin et al., 2014a) and MPII (Andriluka et al., 2014), APE is hindered by a lack of such datasets, especially for endangered species and wildlife (Zuffi et al., 2019a). Also, it is often impractical to place animals in controlled environments to obtain ground truth poses, although zoos provide a semi-structured environment towards data collection. Therefore, unsupervised approaches that do not rely on labelled data have become popular in APE (Wu et al., 2023b; Li and Lee, 2021b).
1.4. Applications of Animal Pose Estimation
HPE has gained significant attention recently due to its applications in many domains, such as human-computer interaction, sports analytics, and healthcare. On the other hand, APE research has not been as prominent due to limited funding and methods. Despite this, APE is critically essential in various applications, as explained below and summarised in table 1.
1.4.1. Tracking
APE has been applied in various tracking contexts; for example, in the CATER (Haalck et al., 2023) framework, APE is combined with environment reconstruction to track ants in complex natural environments, providing fine-scale motion trajectories even in visually challenging conditions. The SMART-BARN (Nagy et al., 2023) system uses APE to track multiple animal species simultaneously in a large 3D arena and achieve accurate tracking and behavioural analysis in indoor and naturalistic environments. In the FaceMap (Syeda et al., 2024) framework, APE is used for tracking the orofacial movements of mice, enabling detailed analysis of their facial expressions and related neural activities. For honey bees, APE tracked an entire colony over extended periods, capturing complex social behaviours and interactions without needing physical tags, enhancing our understanding of colony dynamics and individual behaviours (Bozek et al., 2021).
Table 1. Table listing various applications of animal pose estimation. Behav. denotes Behaviour Analysis.
Sl Year APE Paper Animal Classes Domain
1 2022 I-MuPPET (Urs Waldmann and Goldlücke, 2022) Pigeon Tracking
2 2023 Bozek et al. (Bozek et al., 2021) Honey Bee Tracking
3 2023 CATER (Haalck et al., 2023) Desert Ants Tracking
4 2023 SMART-BARN (Nagy et al., 2023) Insects, Birds, Tracking
Mammals
5 2023 3D-POP (Naik et al., 2023) Birds Tracking
6 2023 Facemap (Syeda et al., 2024) Mice Tracking
7 2023 DeepWild (Wiltshire et al., 2023) Chimpanzees, Tracking
Bonobos
8 2024 3D-MuPPET (Urs Waldmann and Kano, 2024) Pigeon Tracking
9 2023 Stathopoulos et al. (Stathopoulos et al., 2023) Sheep, Cows, Bear Articulation
Giraffe, Horses
10 2022 Sheppard et al. (Sheppard et al., 2022) Mouse Behav.
11 2023 Voloh et al. (Voloh et al., 2023) Macaque Behav.
12 2023 BARN (Yang et al., 2023a) Macaque Behav.
13 2023 Koger et al. (Koger et al., 2023) Gelada Monkeys Behav.
14 2023 AmadeusGPT (Ye et al., 2023) Mammals Behav.
15 2024 Testard et al. (Testard et al., 2024) Macaque Behav.
1.4.2. Articulation
The authors of (Stathopoulos et al., 2023) explore how APE can be used to understand animal articulation by predicting their 3D shapes from monocular images. They demonstrate that training a 2D keypoint estimator on a small set of labelled images (50-150) can generate pseudo-labels on a more extensive set of unlabeled web images. Then, these pseudo-labels are used to train models for 3D shape reconstruction of various articulated quadrupeds.
1.4.3. Behaviour Analysis
In agriculture, animal behaviour analysis based on APE plays an important role by enabling effective livestock management and health surveillance, which in turn provides continuous monitoring of movement and also early detection of animal’s illness, thereby ensuring optimal animal welfare (Fang et al., 2021; Russello et al., 2022; Farahnakian et al., 2021; Gong et al., 2022; Doornweerd et al., 2021; Fang et al., 2022). Recently, more APE methods have been applied to study the behaviour of more species, such as macaques (Voloh et al., 2023; Yang et al., 2023a; Testard et al., 2024), Gelada monkeys (Koger et al., 2023), and mice (Sheppard et al., 2022). Since 2023, there has been a notable shift towards integrating language prompts with APE in behaviour analysis research, exemplified by developing models like AmadeusGPT (Ye et al., 2023). This combination has opened up new possibilities for more comprehensive interpretations of animal behaviour.
1.5. Scope of this Survey
We aim to provide a systematic review of APE, focusing on developments in deep learning-based methodologies for both 2D and 3D scenarios. It categorizes existing APE methods according to their sensor and modality inputs rather than limiting the scope to monocular images or videos.
Several existing reviews have been reported on HPE and APE separately. Prior surveys (Zheng et al., 2020; Ji and Liu, 2009; Sarafianos et al., 2016; Khan and Wan, 2018; Gong et al., 2016; Liu et al., 2022a; Song et al., 2021; Dubey and Dixit, 2023; Wang et al., 2021b; Moeslund and Granum, 2001; Moeslund et al., 2006; Poppe, 2007) have primarily concentrated on vision-based human motion capture, encompassing aspects of pose estimation, motion tracking, and action recognition. While these studies have significantly contributed to HPE, there remains a conspicuous absence of in-depth comparison and analysis between HPE and APE techniques. Reference (Jiang et al., 2022) discusses vision-based APE between 2013 and 2021, yet it falls short of offering an exhaustive discussion on the existing challenges and future directions in APE. Another limitation is that only a limited number of reviews (Zheng et al., 2020) discuss HPE or APE from sources other than vision-based inputs. This paper aims to bridge these gaps in the literature by offering a detailed analysis that extends beyond the scope of prior surveys and reviews:
  1. (1)
    We meticulously reviewed 178 papers up to 2024, categorizing them based on different parameters such as their output forms (e.g. 2D vs. 3D pose estimation; keypoint vs. mesh pose representation), input sensor and modality types (monocular inputs vs. other sensory inputs; single-modal APE vs. multi-modal APE), learning paradigms (e.g. heatmap-based methods vs. regression methods; top-down methods vs. bottom-up methods; supervised vs. semi-supervised vs. self-supervised vs. unsupervised APE methods), experimental setup (e.g. single-camera-view vs. multiple-camera-views approaches; indoor vs. in-the-wild setup; the range of the sensors).
  2. (2)
    We summarised and discussed datasets and evaluation metrics for both 2D and 3D APEs not only based on RGB cameras but also include other sensors and modalities such as motion capture (MoCap) systems, RGBD, LiDAR, infrared, 3D Scanner, IMU, acoustic and language cues.
  3. (3)
    We delved into the potential and limitations of current multi-sensor and multi-modal APE, highlighting how these approaches can complement or enhance unimodal APE methodologies.
  4. (4)
    We presented an extensive comparison between HPE and APE. This includes the transferability of knowledge and techniques between these two fields and how innovations in APE can reciprocally enrich HPE and the broader machine learning paradigm. It also explained the trend towards self-supervised and unsupervised methods for APE and compared the advantages and disadvantages of the existing supervised methods.
  5. (5)
    We also provided an overview of a wide array of applications of APE, ranging from wildlife conservation, biomechanics and neuroscience to 3D animal asset articulation, illustrating the practical implications and the interdisciplinary nature of APE.
  6. (6)
    Finally, we thoroughly discussed the key challenges in APE and pointed towards potential future research avenues to enhance performance and applicability.
Table 2. RGB-based APE Dataset (Images-based). Column N indicates no. of animals: ”M” denotes multiple animals’ pose is annotated in one frame, while ”S” denotes a single animal’s pose is annotated. Column I/W is ”indoor” [Uncaptioned image] or ”in-the-wild” [Uncaptioned image] set-up.
Sl Year Datasets & Size Animal Classes Source/Experiment N I/W Annotation
set-up
RGB-based APE Dataset - Images-based
1 2011 Stanford Dogs (Khosla et al., 2011) Dogs (120 breeds) a. ImageNet data N/A [Uncaptioned image] Class labels,
20k images b. Single view bounding box
2 2014 COCO (Lin et al., 2014b) Cat, Dog, Horse, Bear, a. Online Repositories N/A [Uncaptioned image] & [Uncaptioned image] Bounding box, class
1k images for animals Sheep, Zebra, Giraffe b. Single view label, segmentation
3 2018 Animals-10 (C., 2018) Dog, Cat, Horse, a. Google Images N/A [Uncaptioned image] & [Uncaptioned image] N/A
28k images Sheep, Cow, Elephant b. Single view
4 2019 Animal Pose (Cao et al., 2019) Dogs, Cats, Cows, a. Animals-10 (C., 2018) & S [Uncaptioned image] & [Uncaptioned image] 20 keypoints (2D pose),
4k images Horses, Sheep, COCO (Lin et al., 2014b) bounding box
& other 7 categories b. Single view
5 2019 Grevy’s Zebra (Zuffi et al., 2019b) Synthetic Zebra a. Computer generated S [Uncaptioned image] 28 keypoints (2D pose),
12,850 images b. Single view 3D mesh
6 2020 Synthetic Animal (Mu et al., 2020) Synthetic Hound, a. COCO2017 (Lin et al., 2014b) S [Uncaptioned image] & [Uncaptioned image] 18 keypoints (2D &
10k images Horse, Sheep, b. Single view 3D pose), depth maps,
Elephant, Tiger segmentation
7 2020 StanfordExtra (Biggs et al., 2020) Dogs a. Stanford Dogs (Khosla et al., 2011) S [Uncaptioned image] 24 keypoints (2D pose),
12k images b. Single view segmentation
8 2021 AP-10K (Yu et al., 2021b) 23 animal families, a. Online Repositories [Uncaptioned image] 17 keypoints (2D pose)
10,015 images 54 species b. Single view M
9 2021 MacaquePose(Labuguen et al., 2021) Macaque Monkey a. Manually collected M [Uncaptioned image] 17 keypoint (2D pose)
13,000 images & Onlince
b. Single view
10 2022 Animal Kingdom (Ng et al., 2022b) 850 animal species a. Online Repositories S [Uncaptioned image] 23 keypoints (2D pose)
35k images b. Single view
11 2022 OpenMonkey 26 Monkey species a. Manually collected S [Uncaptioned image] 17 keypoints (2D pose)
Challenge (Yuan Yao and Park, 2022) & Online
111k images b. Single view
12 2023 OpenApePose (Desai et al., 2023) 6 Ape species a. Manually collected S [Uncaptioned image] 16 keypoints (2D pose)
71,868 images b. Single view
13 2023 Animal3D (Xu et al., 2024) 40 Mammal species a. PartImageNet (He et al., 2022) S [Uncaptioned image] 26 keypoints (2D pose),
3379 images &   COCO (Lin et al., 2014b) 3D mesh
b. Single view
2. Datasets
It is hard to obtain accurate ground truth poses for APE, especially for wild animals, due to the unpredictable natural environments, the wide range of animal behaviours, occlusions, and the lack of controlled setups for precise annotation. In this paper, we have categorized APE datasets into two types based on the data they provide and the sensor or modality they use: (1) RGB-based APE datasets, which rely only on RGB images or video sequences (often involving multiple frames) to provide animal pose annotations. (2) Other-sensor-based datasets, which utilize sensors and modalities from sources beyond RGB, such as MoCap systems, 3D scanners, infrared cameras, IMUs, and LiDAR to acquire pose information.
2.1. RGB-based Animal Pose Estimation Dataset
In RGB-based APE datasets, we can further divide them into two categories based on the data type: 1) image-based APE datasets, which rely on individual RGB images to provide pose annotations for animals (table 2); 2) video-based APE datasets, which utilize video sequences, consisting of multiple frames, to capture animal pose labels over time (table 3).
2.1.1. Images-based Animal Pose Estimation Dataset
Early datasets in this category lacked large-scale annotated animal pose labels. For example, the Stanford Dogs (2011) dataset (Khosla et al., 2011) includes 20k images of 120 dog breeds with class labels and bounding boxes but no pose keypoints. Similarly, the COCO (2014) dataset (Lin et al., 2014b) with 1k images only provides annotations such as bounding boxes, segmentation, and class labels for animals like cats, dogs, and horses. The Animals-10 (2018) dataset (C., 2018), consisting of 14k images of common animals such as dogs, cats, and horses, also lacks animal keypoint annotations.
As there are few pose-labeled animal datasets available, more 2D pose-labeled animal datasets have been developed to evaluate the performance of APE objectively. For instance, the Animal Pose (2019) (Cao et al., 2019) consists of 4k images across 10 animal categories, with each image annotated with 20 2D keypoints to represent the animal skeleton. The StanfordExtra (2020) (Biggs et al., 2020) offers 12k dogs images with 24 2D pose keypoints. The AP-10K (2021) (Yu et al., 2021b) provides 10k images from 23 animal families, with 17 2D keypoints labeled for each detected animal. The recent Animal Kingdom (2022) (Ng et al., 2022b), the OpenMonkeyChallenge (2022) (Yuan Yao and Park, 2022), and the OpenApePose (2023) (Desai et al., 2023) all include annotated 2D animal poses keypoints.
Beyond 2D keypoint pose labels, more recent datasets have started providing 3D pose annotations, including 3D keypoints and meshes. Examples include the Synthetic Grevy’s Zebra (2019) (Zuffi et al., 2019b), the Synthetic Animal (2020) (Mu et al., 2020), and the Animal3D (2023) (Xu et al., 2024), which offer both high-quality 2D keypoint annotations and 3D pose labels. This development reflects the growing diversity and depth of datasets available for APE research.
2.1.2. Video-based Animal Pose Estimation Dataset
Table 3. RGB-based APE Dataset (Video-based). Column [Uncaptioned image] Rng is camera range. NGD denotes National Geographic Documentaries; the symbol [Uncaptioned image] denotes cameras. All the other symbols and column meanings are defined in table 2.
Sl Year Datasets & Size Animal Classes Source & Experiment set-up [Uncaptioned image] Rng N I/W Annotation
RGB-based APE Dataset - Videos-based
14 2016 TigDog (Del Pero et al., 2017, 2015) Dogs, Horses, a. YouTube & NGD videos N/A S [Uncaptioned image] 19 keypoints
110k frames Tigers b. Single view (2D pose)
55 videos
15 2018 BADJA (Biggs et al., 2018) Bears, Camels, a. DAVIS (Pont-Tuset et al., 2017) videos 1-20m S [Uncaptioned image] 20 keypoints
11 videos Dogs, Horses, b. Single view (2D & 3D pose),
Tigers, Cats, Cows segmentation
16 2020 ATRW (Li et al., 2020b) Amur tigers a. Manually collected N/A M [Uncaptioned image] 15 keypoints
8k videos b. Single view (2D pose)
17 2020 OpenMonkeyPose Monkeys a. Manually collected 0-2.5m M [Uncaptioned image] 13 keypoints
(Bala et al., 2020) 195,228 frames b. Multi views (62 [Uncaptioned image] views) (3D pose)
18 2021 AcinoSet (Joska et al., 2021) Cheetahs a. Manually collected 3-12m S [Uncaptioned image] 20 keypoints
119k frames b. Multi views (2D & 3D pose)
93 videos (6 GoPro [Uncaptioned image] viewpoint)
19 2021 Horse-30 (Mathis et al., 2021) Horses a. Manually collected N/A S [Uncaptioned image] 22 keypoints
8114 frames b. Single view (2D pose)
20 2022 APT-36K (Yang et al., 2022b) Quadruped a. YouTube videos N/A M [Uncaptioned image] 17 keypoints
36k frames Mammals b. Single view (2D pose)
2.4k videos (30 Species)
21 2023 ChimpACT (Ma et al., 2023) Chimpanzees a. Manually collected N/A M [Uncaptioned image] & [Uncaptioned image] 16 keypoints
16,028 frames b. Single view (2D pose)
163 videos
We have listed several well-known datasets in this category, with most pose annotations being presented in 2D and 3D keypoint formats (table 3). In particular, most 2D keypoint pose labels in these datasets (e.g. Horse-30 (Mathis et al., 2021), APT-36K (Yang et al., 2022b), ChimpACT (Ma et al., 2023)) are provided by well-trained annotators. For instance, the TigDog (2016) (Del Pero et al., 2017, 2015) includes 110k frames of in-the-wild dogs, horses, and tigers, with 19 keypoint annotations for 2D poses. The Horse-30 (2021) (Mathis et al., 2021) provides 8k frames of horses in-the-wild, with 22 annotated 2D keypoints. While these datasets are valuable, they are limited to individual animals. However, there is a recent growing trend toward providing multi-animal pose annotations (e.g. ATRW (Li et al., 2020b), APT-36K (Yang et al., 2022b) and ChimpACT (Ma et al., 2023)) to account for the frequent interactions between quadruped animals.
There is also a small collection of publicly available 3D animal pose benchmark datasets. For example, the BADJA (2018) (Biggs et al., 2018) contains 11 in-the-wild animal videos annotated with 20 3D joints mapped onto a SMAL mesh. The AcinoSet (2021) (Joska et al., 2021) offers 7k frames of hand-labelled 3D keypoints from cheetahs, capturing their running behaviour in-the-wild. Additionally, the OpenMonkeyPose (2020) (Bala et al., 2020) includes 195k well-annotated 3D keypoint pose instances, spanning a large variation of poses and positions of monkeys across diverse activities. It is worth noting that both the OpenMonkeyPose (Bala et al., 2020) and AcinoSet (Joska et al., 2021) include synchronized multiple-camera-views footage and camera calibration, enabling multiple-camera-views geometry and triangulation for precise 3D pose estimation (e.g. OpenMonkeyPose offers 62-viewpoint, while AcinoSet uses 6). We also discuss using MoCap systems with markers to obtain ground truth 3D animal poses in the following subsection.
One common limitation across many datasets in this category is their focus on specific animal species. Due to the significant variations in appearance, behaviour, and joint distributions between species, APE models trained on these datasets often fail to generalize to unseen species, leading to poor performance in real-world scenarios. To fill this gap, the APT-36K (2022) (Yang et al., 2022b) was introduced, covering 2k video clips of 30 quadruped species with high-quality 2D keypoint annotations. This enables inter-species generalization testing and reflects the need for models that work across diverse species.
Table 4. Other-Sensor-based APE Dataset. [Uncaptioned image] is markers; vol. is volume. All the other symbols and column meanings are defined in table 2 and table 3.
Sl Year Datasets & Size Animal Source/Experiment [Uncaptioned image] Rng N I/W Annotation
Classes set-up
Other-Sensor-based APE Dataset - RGB+MoCap-based
22 2021 PAIR-R24M (Marshall et al., 2021) Rat a. Manually collected N/A M [Uncaptioned image] 12 keypoints
24.3M frames b. Multi views (6 RGB video [Uncaptioned image]) (3D pose)
c. 1 Mocap [Uncaptioned image] & 12 [Uncaptioned image]
23 2021 Rat7M (Dunn et al., 2021) Rat a. Manually collected N/A S [Uncaptioned image] 20 keypoints
7M frames b. Multi views (6 RGB video [Uncaptioned image]) (3D pose)
c. 12 Mocap [Uncaptioned image] & 20 [Uncaptioned image]
24 2023 3D-POP (Naik et al., 2023) Pigeons a. Manually collected 0-4.2m (RGB) M [Uncaptioned image] 9 keypoints
300k frames b. Multi views (4 RGB [Uncaptioned image]) 15x7x4m vol. (2D & 3D pose)
4 videos c. 30 Mocap [Uncaptioned image] & 8 [Uncaptioned image] (Mocap)
25 2024 PFERD (Li et al., 2024c) Horses a. Manually collected 16x20m S [Uncaptioned image] 3D Mesh,
55+ hrs video b. Multi views (10 RGB [Uncaptioned image]) from the 2D joints
c. 56 Mocap [Uncaptioned image] & over 100 [Uncaptioned image] arena center
Other-Sensor-based APE Dataset - RGB+Depth+MoCap-based
26 2020 RGBD-Dog (Kearney et al., 2020a) Dogs a. Manually collected 0.5-4.5m S [Uncaptioned image] 3D Mesh,
No size given b. Multi views (6 RGBD sensors) (Depth) 3D Skeletal data,
c. 20 infrared [Uncaptioned image] & MoCap suit N/A (infrared) RGBD video
Other-Sensor-based APE Dataset - RGB+LiDAR-based
27 2024 WildPose (Muramatsu et al., 2024) Lion, a. Manually collected 2–120m M [Uncaptioned image] 15 keypoints,
No size given Springbok b. Single view 3D point cloud
c. 1 RGB Camera + LiDAR
Other-Sensor-based APE Dataset - Infrared-based
28 2023 LoTE-Animal (Liu et al., 2023) 11 animal a. Manually collected & Online N/A S [Uncaptioned image] 17 keypoints
33k images species b. Single view (2D pose)
Other-Sensor-based APE Dataset - 3D scanner-based
29 2024 VAREN (Zuffi et al., 2024) Horses a. Manually collected N/A S [Uncaptioned image] 3D Mesh
4000 3D scans b. No. of scanner Not given
Other-Sensor-based APE Dataset - MoCap+IMU-based
30 2024 Yigit et al. (Yigit et al., 2022) Horse a. Manually collected S [Uncaptioned image] & [Uncaptioned image] 3D Motion data
No size given b. Multi views
c. MoCap system (8 [Uncaptioned image] & 37 [Uncaptioned image]) N/A (MoCap)
d. 4 IMUs 0-50m (IMU)
2.2. Other-Sensor-based Animal Pose Estimation Dataset
With developments in sensor technology, more APE datasets start to incorporate data from sources beyond traditional RGB cameras, such as depth sensors, MoCap systems, 3D scanners, LiDARs, and IMUs, offering increased accuracy and versatility (table 4).
As discussed in section 2.1, researchers have traditionally relied on manual annotations or triangulation techniques to obtain 3D ground truth pose labels. An alternative approach for achieving high accuracy and low-noise 3D ground-truth animal pose data is to use a maker-based MoCap system with RGB cameras. For example, the Rat7M (2021) (Dunn et al., 2021) and PFERD (2024) (Li et al., 2024c) utilize MoCap system with RGB cameras and a set of makers to generate 3D ground truth pose data for rats and horses, respectively. Although these are effective for tracking individual animals from multiple views with temporal consistency, these datasets are limited by their focus on single animals and dependence on a large number of markers. To address these limitations, recent datasets like PAIR-R24M (2021) (Marshall et al., 2021) and 3D-POP (2023) (Naik et al., 2023) have been developed. These datasets use MoCap systems with fewer markers (12 and 8 markers, respectively) to track the 3D postures of multiple animals in the same frame, offering more scalable solutions for multi-animal tracking. Similarly, a recent dataset (Kearney et al., 2020a) utilizes MoCap systems with 63–82 markers to generate highly accurate 3D ground truth pose labels for dogs. This dataset also integrates depth sensors (i.e. RGBD), which enable us to develop more flexible solutions for 3D APE without relying on physical markers. Although the MoCap-based dataset provides high-quality 3D ground truth pose labels, APE algorithms trained using these MoCap datasets may exhibit poor generalization to in-the-wild scenarios, as the data collection for the MoCap-based datasets is all done indoors in a controlled environment, as you can see in table 4.
WildPose (Muramatsu et al., 2024) is another promising multi-sensor system that combine RGB and LiDAR data. This system captures 2D RGB video and 3D point cloud data from a distance (up to 120 meters), making it particularly suited for studying free-ranging animals in natural habitats.
In addition to these technologies, infrared cameras for camera-trapping in the wild, which use infrared LEDs to illuminate the scene to enable nighttime observations, have also been employed for wild animal pose labelling. For instance, the LoTE-Animal (2023) (Liu et al., 2023) uses infrared trap cameras to collect large-scale video footage of endangered animals across 11 species. This dataset captures a wide range of variations—such as ecological seasons, weather conditions, viewpoints, and habitats—and provides both instance segmentation labels and 2D pose keypoint annotations. The LoTE-Animal dataset is especially valuable for advancing semi-supervised and self-supervised APE models, with potential applications in conservation efforts for endangered species.
Moreover, the VAREN (2024) dataset (Zuffi et al., 2024) focuses on 3D mesh representation, featuring 4k 3D scans of horses provided by the 3D scanner; this further advances 3D modelling and representation in APE research.
Lastly, the IMU-based animal pose dataset collected by (Yigit et al., 2022) consists of acceleration and gyroscope measurements from sensors attached to the horse’s limbs. The data captures each limb’s 3D motion characteristics (like linear acceleration and angular velocity). It provides 3D information on movement and orientation but lacks absolute coordinates, as IMUs only capture relative motion.
These emerging sensor-based datasets offer new opportunities for APE, enhancing the precision and breadth of APE research across diverse animal species and environments.
3. Error Metrics and Evaluation Methods
This section provides insight into the error (err.) metrics and evaluation (eval.) process of the APE methods based on different sensors and modalities.
3.1. RGB-based 2D Animal Pose Estimation Metric
This section discusses the common evaluation metrics used for RGB-based APE methods that extract poses in the 2D keypoint format, and more evaluation metrics are given in table 5 and table 6.
2D single-animal pose estimation methods like DeepLabCut (Mathis et al., 2018) and LEAP (Pereira et al., 2019) use a common metric like Root Means Square Error (RMSE) to evaluate their performance. RMSE is versatile and can be used for both 2D and 3D pose estimation and for evaluating various types of keypoints, regardless of the specific application domain. In addition to RMSE, many APE methods also use the Probability of Correct Keypoint (PCK), which measures the fraction of joints where the predicted joint positions are within a certain threshold distance from the ground truth joint positions. It is a normalised measure that indicates how well the keypoints are localised, and a higher PCK value indicates better localisation performance. Single-animal pose estimation methods like Synthetic-to-Real (Li and Lee, 2021b), Kim et al. (Kim et al., 2022), ScarceNet (Li and Lee, 2023), D-Gen (Li et al., 2023) and FSKD (Lu and Koniusz, 2022) use this metric for evaluation. Moreover, Li et al. (Li et al., 2020a) evaluate their single-animal APE methods with PCK, RMSE and Area Under Curve (AUC) along with an extension to these metrics with Permutation Invariant (PI). Li et al. (Li et al., 2020a) introduce PI metrics to evaluate the accuracy of pose estimation in challenging scenarios where it is difficult to uniquely identify limb identities (e.g. in fruit flies) with traditional metrics. PI metrics are designed to account for the ambiguity in identifying corresponding limbs or joints between the predicted pose and ground truth.
Object Keypoint Similarity (OKS) has also been used in evaluating 2D single-animal pose estimation methods like D-Gen (Li et al., 2023); 2D multiple-animal pose estimation methods like DepthFormer (Liu et al., 2022b), Straka et al. (Straka et al., 2024), SLEAP (Pereira et al., 2022), and also 2D video based semi-supervised method like SemiMultiPose (Blau et al., 2022). OKS quantifies how well the keypoints are located to spatial scale and keypoint visibility, with closeness defined based on ground truth. The OKS lies in the range of 0 to 1, and a value close to 1 indicates that the detected keypoints coincide with the ground-truth keypoints. The OKS scores for every keypoint are computed for all objects in the dataset, and the average OKS is used to evaluate a model. Another common metric used for 2D APE is mean Average Precision (mAP), which has been used in evaluating ScarceNet (Li and Lee, 2023), SIPEC (Marks et al., 2022), SuperAnimal-Quadruped (Shaokai Ye and Mathis, 2024), OpenPose (Cao et al., 2017), DLCRNet (Lauer et al., 2022), maDLC (Lauer et al., 2022), SLEAP (Pereira et al., 2022), SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SemiMultiPose (Blau et al., 2022). The mAP is an evaluation metric based on OKS and has become widely used for pose estimation, particularly following its adoption in the COCO keypoint challenge. OKS measures the similarity between the predicted keypoints and the ground truth, and Average Precision (AP) is then calculated to assess the prediction’s performance on a dataset by comparing it against different OKS thresholds. The mAP is the mean of several AP values, each calculated using a different threshold.
Last but not least, another metric that has been used for the 2D video-based APE method evaluation is Canonical Correlation Analysis (CCA), which is a statistical technique used to examine the relationship between two sets of variables. It determines linear combinations of the variables in each set that are maximally correlated; this metric has been used for evaluating the Lightning Pose (Biderman et al., 2023) method, however, we have not seen wide adoption of this evaluation metric within the APE community.
3.2. RGB-based 3D Animal Pose Estimation Metric
This section discusses the common evaluation metrics used for both keypoint-based and mesh-based 3D RGB models within the APE community, more evaluation metrics for the 3D APE methods are given in table 7.
3D APE metrics can also incorporate some of the metrics used in 2D APE. For example, PCK is employed in methods such as Three-D Safari (Zuffi et al., 2019a), SMBLD (Biggs et al., 2020), Li et al. (Li and Lee, 2021a), Birds of a Feather (Wang et al., 2021a), BITE (Rüegg et al., 2023), MagicPony (Wu et al., 2023b), DigiDogs (Shooter et al., 2024), Monet (Yao et al., 2019), DAN-NCE (Dunn et al., 2021), and Muramatsu et al. (Muramatsu et al., 2022). The well-known RMSE metric is also used in 3D APE methods such as Hu et al. (Hu et al., 2023), Muramatsu et al. (Joska et al., 2021), 3D-UPPER (Ebrahimi et al., 2023), and SBeA (Han et al., 2024).
Apart from PCK and RMSE, 3D pose estimation-specific metrics such as Mean Per Joint Position Error (MPJPE) have been used in 3D APE methods like DigiDogs (Shooter et al., 2024) and MAMMAL (An et al., 2023). MPJPE measures the average Euclidean distance between the predicted and ground truth joint positions across all joints in the skeleton, quantifying the accuracy of the joint position estimation. A lower MPJPE indicates that the predicted joint positions are closer to the actual joint locations. An extension of the MPJPE metric is N-MPJPE, which has been used by Dai et al. (Dai et al., 2023) for their 3D APE model. N-MPJPE involves normalising the ground truth 3D pose by subtracting the mean and dividing by the standard deviation, aligning the predicted 3D pose with the normalised ground truth using Procrustes alignment, and then calculating the MPJPE, which results in Normalized-MPJPE.
For mesh-based APE as discussed in section 4.3.1, a well-known evaluation metric is the 3D Chamfer Distance, which has been used in several model-free methods such as DOVE (Wu et al., 2023a), MagicPony (Wu et al., 2023b), and BANMo (Yang et al., 2022a). The 3D Chamfer Distance measures the average distance between the reconstructed surface points and the ground truth surface points. It finds the nearest neighbour matches between the two sets of points and computes the distance, which is sensitive to outliers. A lower value indicates better reconstruction quality.
Another approach in mesh-based APE is the model-based pose estimation methods, for example, SMAL (Zuffi et al., 2017). In these approaches, metrics like Scan-to-Mesh distance, keypoint reprojection error, and silhouette reprojection error are commonly used. The Scan-to-Mesh distance measures the accuracy of the reconstructed 3D shape by calculating the distance from each point on the scan to the closest point on the predicted mesh surface. Keypoint reprojection error evaluates the pose estimation by comparing the 2D projection of 3D keypoints from the predicted model with the annotated 2D keypoints in the images. It assesses how well the 3D pose model aligns with the 2D image data. Similarly, the silhouette reprojection error measures how accurately the 3D model silhouette matches the ground truth silhouette in the 2D image, indicating how well the 3D shape projection aligns with the visible boundaries of the animal in the image. Other model-based methods, such as SMBLD (Biggs et al., 2020) and Li et al. (Li and Lee, 2021a), use metrics like PCK and Intersection over Union (IoU) for performance evaluation.
Moreover, due to the scarcity of animal datasets that include 3D pose annotations, comprehensive quantitative evaluations of 3D pose predictions are not always feasible. Therefore, researchers often assess 3D pose predictions using qualitative means, as Sosa et al. (Sosa and Hogg, 2023) explained. Examples of qualitative evaluations can be found in Salem et al. (Salem et al., 2019), Sosa et al. (Sosa and Hogg, 2023), Monet (Yao et al., 2019), and Muramatsu et al. (Joska et al., 2021).
3.3. Other-Modality-based Animal Pose Estimation Metric
Different sensor modalities have different ways of evaluating their APE models. For example, an IMU-based APE method (Yigit et al., 2022) is evaluated by calculating the mean and standard deviation (Std) of the estimated joint angle errors and comparing the predictions with ground truth data obtained through a MoCap system. Additionally, the model is assessed for Intra-individual Variability (IAV) and Inter-individual Variability (IEV), which quantify the consistency of limb movement across different gait cycles and between different horses. Another metric used is the Variance Ratio (VR), which captures the repeatability of the gait cycle by evaluating how similar joint movements are within multiple cycles. This thorough approach ensures that this IMU-based model’s performance is evaluated for accuracy and consistency across different conditions.
On the other hand, one of the RGBD-based APE methods, RGBD-Dog (Kearney et al., 2020a), is evaluated using MPJPE and PCK. The evaluation is conducted across different body parts and for entire dog poses. Additionally, Procrustes Analysis (PA) is applied to align the predicted skeleton with the ground truth skeleton. The PA is a statistical shape analysis technique. In the context of pose estimation, it aligns the predicted 3D skeleton with the ground truth skeleton by minimising the sum of squared distances between corresponding points in the two sets. In the RGBD-Dog (Kearney et al., 2020a), PA is applied to refine the existing PCK and MPJPE evaluation metrics to PA-PCK, PA-PCK 3D and PA-MPJPE, which is then used to measure the error after aligning the predicted and ground truth skeletons to focus purely on pose accuracy. Other multi-modal systems, for example, Sound of Motion (RGB+acoustic) (Li et al., 2022) and CLHOP (audio+video) (Li et al., 2024a), also use PA-MPJPE for their APE evaluation in addition to PCK and IoU, which are common in the RGB-based APE evaluations.
This section briefly summarises some of the metrics for non-RGB-based APE models, more evaluation metrics are provided in table 8.
4. RGB-based Animal Pose Estimation
In recent years, APE has achieved significant progress. Among the various modalities utilised for APE, RGB-based approaches have emerged as the most developed and extensively studied research method (Pereira et al., 2019; Cao et al., 2019; Li and Lee, 2021b; Lauer et al., 2022). These approaches leverage visual information captured by RGB cameras, take animal images or videos as inputs, and subsequently produce the animal’s pose, which manifests as a 2D skeletal keypoints representation of pose, 3D skeletal keypoints representation of pose, or 3D animal mesh reconstructions, as exemplified in LABEL:fig:Skeleton_Pose_Representations and LABEL:fig:Body_Mesh_Reconstructions. In addition to skeletal keypoints and mesh representation for pose estimations, dense pose representations have also been used mainly for humans (Riza Alp Güler, 2018) and chimpanzees (Sanakoyeu et al., 2020). In this paper, skeletal keypoints and mesh representations APE have been mainly analysed in detail, while keypoint-based methods remain dominant due to their established presence, easier access to annotated data, and lower computational requirements.
Before we go into the details of RGB-based APE methods, let us first discuss the development of this branch. Traditionally, we use direct visual observation to extract the animal poses, which can offer real-time and context-rich data. However, it suffers from inherent drawbacks, including observer bias, mental fatigue (Graving et al., 2019), and long data collection and observation times. Later, researchers start to use video monitoring and recording to expedite data collection and minimise the influence of human involvement (Graving et al., 2019). However, manual scoring of videos by eyes still remains time-consuming and often fails to capture subtle and fast-paced movements that are integral to various behaviours. Moreover, the reliance on human observers restricts the dataset size, this may also lead to statistical errors that limit researchers’ ability to address scientific questions accurately.
Recently, automated body kinematics tracking techniques utilising advanced imaging software and hardware have been developed to tackle the challenges in manual scoring. Marker-based methods rely on attaching markers (e.g. MoCap suits (Kearney et al., 2020b), animal-borne cameras (Patel et al., 2017)) to animals for posture measurements. While these methods can achieve precise pose estimation, the reliance on physical markers can pose practical challenges, especially when investigating innate behaviours in natural habitats (Wu et al., 2023b) or when observing the movements of small creatures like mice (Lauer et al., 2022) and fish. Furthermore, most marker-based approaches are limited to highly controlled lab environments and specialised hardware setups. Such methodologies are often impractical when the target animals demonstrate varied behaviours within unfamiliar environments.
With the rapid advancement in Computer Vision (CV), many automated markerless APE methods (Karashchuk et al., 2021; Mathis et al., 2018; Lauer et al., 2022; Pereira et al., 2019; Graving et al., 2019) start to employ supervised learning to directly infer and estimate animals’ pose and movements from images or videos, removing the need for physical markers that may alter natural behaviour. For example, the markerless method DeepLabCut (Mathis et al., 2018) has been used for behaviour tracking in wild chimpanzees and bonobos (Wiltshire et al., 2023).
However, a significant obstacle in APE is the acquisition of accurate ground truth poses as discussed in section 2. To address the scarcity of large-scale, well-labelled animal pose datasets, many weakly and semi-supervised (Cao et al., 2019), semi-supervised (Kim et al., 2022; Blau et al., 2022; Biderman et al., 2023; Yao et al., 2019), self-supervised (Sosa and Hogg, 2023; Shooter et al., 2024; Yang et al., 2022a; Han et al., 2024) and unsupervised learning (Dai et al., 2023; Kim et al., 2022; Li and Lee, 2021b; Ebrahimi et al., 2023; Shaokai Ye and Mathis, 2024; Wu et al., 2023a, b; Li et al., 2024b) approaches have been developed. For instance, recent literature (Dai et al., 2023) proposes an unsupervised 3D animal canonical pose estimation method with geometric self-supervision; this method necessitates only 2D poses for training, which significantly alleviates the requirement for extensively labelled data. Current APE research trends show growing interest in unsupervised, semi-supervised, and self-supervised methods due to the mentioned challenges to reach the same performance level as supervised methods. However, supervised methods are still a significant focus in current APE studies, especially with the emergence of larger datasets and benchmarking platforms like AP-10K (Yu et al., 2021b), APT-36K (Yang et al., 2022b), TigDog (Del Pero et al., 2015, 2017), and AcinoSet (Joska et al., 2021), as we can see in Table 5-Table 7.
This review paper categorises the existing uni-modal RGB-based APE studies into two main groups: 1) 2D APE methods from images (Table 5, section 4.1) and from videos (Table 5 & Table 6, section 4.2), 2) 3D APE methods from images and videos (Table 7, section 4.3).
4.1. 2D Animal Pose Estimation from an Image
This section focuses on 2D APE methods that do not use temporal information to localize the 2D coordinates of an animal’s or animal’s joints from a single image frame (table 5). Note that the majority of the APE algorithms that can extract poses from videos (i.e. multiple frames) can also extract poses from a single image. In Table 5, column ”I,V” provides the information on APE methods that can be applied on both an image and videos; or only on an image.
Based on the number of target animals being processed in the scene, this pose estimation problem can be further divided into single-animal pose estimation (section 4.1.1) and multiple-animal pose estimation (section 4.1.2). While pose estimation in individual animals has received significant attention and serves as the foundation for most APE studies, extending it to multiple-animal pose estimation is an active area of research that is still evolving (Pereira et al., 2022).
4.1.1. Single Animal
2D single-frame single-animal pose estimation aims to accurately estimate the 2D pose of an animal from a single image. These methodologies share a conceptual foundation with HPE (Zheng et al., 2020) and can be broadly classified into two categories: heatmap-based thods. The former predicts the locations of an animal’s body joints using heatmaps for supervision, necessitating a subsequent postprocessing step to identify the final keypoints based on the predicted heatmaps. In contrast, the latter approach directly regresses the keypoint coordinates of the animal’s body joints, employing an end-to-end deep neural network to establish a mapping from the input image to the body keypoints. LABEL:fig:single_animal illustrates the general architecture of 2D single-frame single-animal pose estimation methods.
Heatmap-based frameworks such as (Pereira et al., 2019; Kim et al., 2022; Li and Lee, 2023; Li et al., 2020a; Li and Lee, 2021b; Graving et al., 2019; Li et al., 2023; Mathis et al., 2018) have become popular in 2D APE. These approaches are structured around creating a 2D heatmap, interpreted as a probabilistic representation of body part locations within the image. The architecture of a typical heatmap-based scheme includes i) a CNN backbone to generate heatmaps from input images, followed by ii) a post-processing step to finalize the pose estimations based on these predictions (LABEL:heatheat). There are some common architectures to input an image and output the corresponding heatmaps in step i); for instance, (Kim et al., 2022; Li and Lee, 2021b; Li et al., 2023; Mathis et al., 2018) uses the ResNet (He et al., 2016) and its variants to produce the heatmap; (Li and Lee, 2023) adapts a variant of the HRNet (Sun et al., 2019) as the backbone; (Li et al., 2020a) applies the stacked hourglass (Newell et al., 2016) architectures. Subsequently, in step ii), the predicted heatmaps are then processed to produce precise 2D spatial coordinates for each body joint using techniques such as peak detection (Graving et al., 2019; Mathis et al., 2018; Pereira et al., 2019). Although heatmap-based approaches are well known for their robustness against occlusions and variations in appearance, their computational complexity has been critiqued for impeding end-to-end training. Nonetheless, DeepPoseKit (Graving et al., 2019) proposes an efficient end-to-end heatmap-based approach.
Regression-based approaches directly regress the coordinates of joint locations as continuous values, employing CNNs to learn a mapping from the input image to the joint coordinates. The output layer of the network consists of multiple units, each representing the $x$ and $y$ coordinates of a joint, and is trained to minimize the distance between the predicted and ground truth coordinates. Although regression-based approaches are less common than heatmap-based methods, they have shown great promise (Lu and Koniusz, 2022). One notable work (Lu and Koniusz, 2022) adopts a regression-based approach to estimate keypoint positions directly, bypassing the intermediate heatmap representation (LABEL:RegressionRegression). Their Few-shot Keypoint Detection (FSKD) model is designed to detect keypoints on unseen species with minimal annotated data, which is particularly useful in challenging scenarios where the species have not been previously encountered or the available annotated data is limited. Regression-based approaches offer the advantage of direct joint location estimation without additional peak detection steps, which can benefit real-time inference. However, they may face long-range 2D offset regression challenges, which may lead to a notable performance gap compared to heatmap-based methods.
Table 5. RGB-based 2D APE Methods. In the column ”I,V”: ”I,V” is the methods that can perform pose estimation on both an image and videos, whereas, ”I” is the APE methods that do only on an image. Column ”L” indicates the learning types as follows: ”S” is supervised methods, ”U” is unsupervised methods, ”SF” is self-supervised methods, ”SM” is semi-supervised methods. [Uncaptioned image] is CNN architecture, [Uncaptioned image] is the use of the attention mechanism, [Uncaptioned image] is GAN, [Uncaptioned image] is underlying architecture as vision transformer, [Uncaptioned image] is Multi-Scale Domain Adaptation Module (MDAM). [Uncaptioned image] is the pose estimation of the candidate algorithm given in the keypoint pose format. [Uncaptioned image] is single-camera-view, [Uncaptioned image] is multi-camera-views. [Uncaptioned image] is ”indoor” experimental set-up and [Uncaptioned image] is ”in-the-wild” experimental set-up. Symbols inside the {} indicate that both categories are used in that method.
Sl Year Methods I,V L Features Evaluated On Metrics
2D Single-Animal Pose Estimation Methods
Heatmap-based:
1 2018 DeepLabCut (Mathis et al., 2018) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mammals RMSE (pixels)
2 2019 LEAP (Pereira et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse RMSE (pixels)
3 2019 DeepPoseKit (Graving et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Zebra (Zuffi et al., 2019b) ([Uncaptioned image]), Euclidean Error,
Desert locust, Fly ([Uncaptioned image]) Bayesian Inference (Graving et al., 2019)
4 2020 Li et al. (Li et al., 2020a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Worm, RMSE, PCK, AUC (pixels),
Zebrafish PI-AUC (4-45 pixels),
PI-PCK (5&15 pixels),
PI-RMSE (pixels)
5 2021 Synthetic-to-Real (Li and Lee, 2021b) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse, TigDog(Del Pero et al., 2017) PCK@0.05 (%)
6 2022 Kim et al. (Kim et al., 2022) I SM,U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] {[Uncaptioned image], [Uncaptioned image]} TigDog(Del Pero et al., 2017), PCK@0.05 (%)
Synthetic Animal (Mu et al., 2020),
Animal Pose (Cao et al., 2019)
7 2023 ScarceNet (Li and Lee, 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), mAP (%), PCK@0.05 (%)
TigDog(Del Pero et al., 2017)
8 2023 D-Gen (Li et al., 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), OKS, PCK@0.05 (%)
Animal Pose (Cao et al., 2019),
Animal Kingdom (Ng et al., 2022a)
Regression-based:
9 2022 FSKD (Lu and Koniusz, 2022) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animal Pose (Cao et al., 2019) PCK
2D Multi-Animal Pose Estimation Methods
Top-down:
10 2022 SIPEC (Marks et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Primates, Mice Pearson Correlation,
RMSE (pixels), mAP (%),
IoU, Dice Coefficient
11 2022 DepthFormer (Liu et al., 2022b) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b) OKS
12 2024 Straka et al. (Straka et al., 2024) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Lynx PCK, OKS, AP
13 2024 SuperAnimal-Quadruped I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), RMSE (pixels), mAP,
(Shaokai Ye and Mathis, 2024) Animal Pose (Cao et al., 2019) Normalized Error
Bottom-up:
14 2017 OpenPose (Cao et al., 2017) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} COCO (Lin et al., 2014b), MPII (Andriluka et al., 2014) mAP (%)
15 2021 Negrete et al. (Blanco Negrete et al., 2021) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] MacaquePose (Labuguen et al., 2021), AP
MSCOCO (Lin et al., 2014a)
16 2021 Farahnakian et al. (Farahnakian et al., 2021) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pig(Nath et al., 2019) Error (pixels), RMSE (pixels)
17 2022 DLCRNet, maDLC (Lauer et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} 3 Mice ([Uncaptioned image]), 1 Mice ([Uncaptioned image]), RMSE (pixels),
2 marmosets ([Uncaptioned image]) mAP, PCK
Bottom-up & Top-Down:
18 2022 SLEAP (Pereira et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, mAP (%), PCK,
Mice, Gerbils OKS, Localization
19 2024 SuperAnimal-TopViewMouse I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), RMSE (pixels), mAP,
(Shaokai Ye and Mathis, 2024) Animal Pose (Cao et al., 2019) Normalized Error
4.1.2. Multiple Animals
We have seen significant developments in markless pose estimation in individual animals. However, extending these methodologies to multiple animals poses unique challenges, particularly in studying social behaviours or animals in their natural environments (Yang et al., 2022b; Gong et al., 2022). For example, it is common for the body parts of one animal to be obscured by another in a multi-animal pose estimation. Another challenge arises from scale variation, where animals at different distances from the camera show differences in apparent sizes. Inter-species variation also introduces additional complexity due to differences in body structures and poses. However, given these challenges, 2D multiple-animal pose estimation research has still garnered significant attention due to its potential applications in various domains, such as animal behaviour analysis, wildlife monitoring, and ecological studies. Methodologies developed in this field are categorized into two main categories: top-down methods (Marks et al., 2022; Liu et al., 2022b; Straka et al., 2024; Shaokai Ye and Mathis, 2024) and bottom-up methods (Cao et al., 2017; Blanco Negrete et al., 2021; Farahnakian et al., 2021; Lauer et al., 2022) (LABEL:fig:multi_animal).
Top-down approaches for multiple-animal pose estimation (LABEL:aaa) generally involve a two-step process: first i) detecting each animal in the input image using pre-trained object detectors and then ii) applying 2D single-animal pose estimation techniques to estimate the pose for each detected instance. Common object detectors in step i) include Faster R-CNN (Ren et al., 2015), the Single Shot MultiBox Detector (SSD) (Liu et al., 2016), You Only Look Once (YOLO) v4 (Gong et al., 2022), and RetinaNet (Lin et al., 2017); common 2D APE methods in step ii) includes variants of ResNet, HRNets and transformer. In the early work, SIPEC (Marks et al., 2022) uses SIPEC:SegNet (a Mask R-CNN architecture (He et al., 2017)) to segment animal instances in the images or videos first, then it uses SIPEC:PoseNet to estimate the pose for each individual. Later, we can spot a trend of using HRNets and transformers as the 2D APE backbone architecture in step ii) to achieve better performance, especially with the limited animal pose data. For example, the top-down DepthFormer method (Liu et al., 2022b) has been proposed to achieve lightweight multiple-animal pose estimation by designing a new basic block based on the transformer structure and the similarities between self-attention and depthwise convolution. In 2024, SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) adapts Faster R-CNN for enhanced feature extraction to detect animal instances in the given image or videos first, and then also uses the transformer-based AnimalTokenPose models to achieve high zero-shot APE performance in OOD animal datasets. In the same paper, both SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SuperAnimal-Quadruped (Shaokai Ye and Mathis, 2024) models also enable us to use HRNet-w32 (Wang et al., 2020) as the underlying architecture to get the final animal 2D pose. Until now, SuperAnimal (Shaokai Ye and Mathis, 2024) has claimed to be the current state-of-the-art across a wide variety of animal species for both 2D images and videos. More information on how SuperAnimal is adopted to videos is discussed in section 4.2. Another recent top-down approach for 2D endangered species (i.e. lynx) pose estimation with limited data (Straka et al., 2024) also leverages the high-resolution network HRNet-w32 (Wang et al., 2020) architecture in step ii) to get accurate 2D animal pose, as this HRNet-w32 architecture helps to maintain detailed spatial information throughout the network. All of these methodologies prove the trend.
In contrast, bottom-up methods (LABEL:bbb) prioritize predicting all keypoint locations in an image first, then grouping them into individual instances. Initial work in this area includes methods like OpenPose (Cao et al., 2017), which introduced the concept of Part Affinity Fields (PAFs) to associate body parts with individuals coherently. This PAFs idea was further adapted in the DLCRNet and maDLC architecture (Lauer et al., 2022) to predict the location and orientation of animal limbs and link the specific keypoints to the correct individual. OpenPose (Cao et al., 2017) was also used in (Blanco Negrete et al., 2021) to associate detected keypoints to particular individuals by taking advantage of similarities in skeleton structure between humans and monkeys. Further refinement in the bottom-up approach was made by works such as (Farahnakian et al., 2021), which focused on pig body part association using DeepLabCut (Mathis et al., 2018).
Both top-down and bottom-up methods have limitations and features. For example, top-down approaches are generally less sensitive to scale variance of objects compared to bottom-up methods, as they use bounding boxes to identify each individual first, then estimate the joints of the detected body in that box region. However, they often face difficulties in scenarios involving occlusions, as the detection step may fail to identify all instances accurately. On the other hand, the primary shortcoming of the bottom-up approaches often stems from incorrect assembly. To allow the users to compare the two competing approaches directly, researchers have developed both bottom-up and top-down variants for maDLC (Lauer et al., 2022), SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SLEAP (Pereira et al., 2022). The decision between these two methods depends on the specific requirements and constraints of the application.
Before we move on to the next section, it is worth noting that some of the APE methods explained in section 4.1 can perform pose estimation on both images and videos, as can be seen in table 5 in column ”IV”. In section 4.2, we explain which of these video-based 2D APE methods extract pose frame by frame, and which uses the spatiotemporal signals of video data in an end-to-end manner for pose estimation.
4.2. 2D Animal Pose Estimation from Videos
2D multi-frame APE transforms consecutive image sequences or videos into probabilistic estimates of the 2D spatial coordinates of keypoint locations for target animals across frames. Based on how these approaches leverage the temporal dimension in video data to extract poses, we categorize 2D APE from videos into two main categories: (1) post-processing methods to incorporate temporal information; and (2) end-to-end methods to incorporate temporal information, as shown in table 6.
Table 6. RGB-based 2D APE methods from videos. All the symbols and column meanings are defined in table 5, and any additional symbols are defined here: [Uncaptioned image] is the underlying architecture as 3D-CNN, [Uncaptioned image]+[Uncaptioned image] is the use of CNN-based pose module followed by an Optical Flow module, [Uncaptioned image] is Bayesian Ensembling, [Uncaptioned image] is Probabilistic Graphical Model.
Sl Year Methods L Features Evaluated On Metrics
2D APE from Videos: End-to-End Methods to Incorporate Temporal Information
1 2021 OptiFlex (Liu et al., 2021) S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image] Mouse, Fish, Monkey, Fly aPCK error (%)
2 2022 T-LEAP (Russello et al., 2022) S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cow PCKh@0.2(%)
3 2022 SLEAP (Pereira et al., 2022) S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, Mice, Gerbils mAP (%), PCK,
OKS, Localization
4 2022 SemiMultiPose (Blau et al., 2022) SM [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mouse, HoneyBee, Fly mAP, OKS
5 2024 Lightning Pose (Biderman et al., 2023) SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse, Fish CCA, Pixel Error, AUC
6 2024 SuperAnimal (Shaokai Ye and Mathis, 2024) U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b) RMSE (pixels), mAP
Animal Pose (Cao et al., 2019) Normalized Error
4.2.1. Post-Processing Methods to Incorporate Temporal Information
Early APE frameworks process each video frame independently to predict pose keypoints, without inherently incorporating temporal features in an end-to-end manner during training. Temporal consistency is typically addressed through optional post-processing methods, such as smoothing or filtering, to improve performance. Some examples that follow this frame-by-frame approach include DeepLabCut (Mathis et al., 2018), DeepPoseKit (Graving et al., 2019), and LEAP (Pereira et al., 2019).
4.2.2. End-to-End Methods to Incorporate Temporal Information
Recently, researchers have achieved more robust predictions by utilising the spatiotemporal information embedded within video data during pose estimation model training in an end-to-end manner. By evaluating inherent temporal coherence and ensuring consistency across successive frames during training, models can produce more accurate and stable pose estimations over time. We highlight several methods that fall in this category in the following bullet points and table 6.
A. Extensions of LEAP method
A notable extension of the LEAP (Pereira et al., 2019) method is T-LEAP (Russello et al., 2022), which conceptualises temporal information as an auxiliary dimension within the CNN framework, thereby employing 3D convolutional layers to capture both spatial and temporal features concurrently to extract animal poses from video data. To leverage the temporal information embedded in videos, T-LEAP processes a sequence of frames of length $T$, and outputs the confidence maps linked to the last frame $t=T$. As for architectural modifications compared to a deeper version of LEAP, all convolution, pooling and transposed-convolution operations transition from 2D to 3D. Notably, these 3D convolutional operations are extended to not only the spatial dimension of the input but also to the temporal dimension, thus facilitating the extraction of spatiotemporal signals from the video data, and making T-LEAP robust for occlusions. For example, in a scenario where a cow’s leg is discernible in one frame but obscured in the subsequent frame, T-LEAP can employ visible information to predict the probable position of the leg in the occluded frame. Another extension of the LEAP model is SLEAP (Pereira et al., 2022), which also incorporates temporal information, but unlike LEAP, this extends the model to multi-animal pose estimation.
B. Optical Flow-based Methods
OptiFlex (Liu et al., 2021) is an optical flow-based methods that allow the estimation of motion vectors between consecutive frames to capture the temporal dynamics of animal poses. In particular, OptiFlex (Liu et al., 2021) combines a deep learning-based pose estimation module called FlexibleBaseline with an optical flow module named OpticalFlow to enhance its APE performance across multiple frames. The FlexibleBaseline module predicts the initial pose for each frame, and the OpticalFlow module converges initial predictions on a target frame and its adjacent frames into a final prediction for the target frame. By tracking the pixel-level displacements over time, optical flow estimation techniques, such as the Lucas-Kanade algorithms (Baker and Matthews, 2004), can compute dense motion fields between frames, thereby inferring the movement of body parts and joints. Suppose a certain body part is occluded or unclear in a frame. In that case, the model can infer its likely position based on the detected motion and poses in adjacent frames, showing its robustness against temporary obstruction of keypoint poses.
C. Future Trend using Semi-supervised and Unsupervised Video Adaptation Learning
Recently, there has been a growing shift towards using semi-supervised learning (e.g. SemiMultiPose (Blau et al., 2022), Lightning Pose (Biderman et al., 2023)) and unsupervised video adaptation learning (e.g. SuperAnimal (Shaokai Ye and Mathis, 2024)) to extract poses from videos. In 2D single-frame APE methods, supervised learning dominates, relying on many labelled frames for training. However, manually adding annotations for videos is costly and labour-intensive, especially for multiple instances. In real-world scenarios like wildlife videography, there are exponentially more abundant unlabeled frames than labelled frames. Given these deficiencies, SemiMultiPose (Blau et al., 2022) proposes a semi-supervised architecture to exploit the abundant structures pervasive in these unlabeled video frames to enhance the training process, which is particularly useful in the sparse labelling video-based scenarios. Similarly, Lightning Pose (2023) (Biderman et al., 2023) incorporates a semi-supervised approach to exploit the spatiotemporal statistics present in unlabeled videos. This methodology introduces unsupervised training objectives that penalise the network when its predictions deviate from the expected smoothness of physical motion, multiple-camera-views geometry, or stray from a low-dimensional subspace of plausible body configurations. Additionally, Lightning Pose predicts the pose for a specific frame by considering the temporal context from surrounding unlabeled frames. This incorporation of context is instrumental in resolving transient occlusions and ambiguities between proximal and morphologically similar body parts. Last, the SuperAnimal (2024) (Shaokai Ye and Mathis, 2024), an unsupervised video adaptation APE method, can be applied to images and videos. However, its approach differs depending on the data type. For images, SuperAnimal operates as a supervised method, relying on labelled data for training. In contrast, for videos, it functions as an unsupervised method by incorporating unsupervised techniques, such as video adaptation and pseudo-labelling, to improve performance on unseen videos. This method helps reduce temporal jitter and refine model outputs without further manual annotation. However, the unsupervised techniques are applied after the model has been initially trained in a supervised manner. These works demonstrate better APE performance with fewer labels, generalise better to unseen videos, and provide smoother and more reliable pose trajectories for downstream analysis, which proves the shift towards semi-supervised and unsupervised learning in APE.
4.3. 3D Animal Pose Estimation from Images and Videos
3D APE refers to the task of determining the locations of animal body joints in 3D space. This challenge extends beyond the 2D APE by adding depth information to the analysis, which allows for a deeper understanding of animal behaviour, biomechanics, and motion. Some 3D APE methods extract poses from monocular imagery and video data streams. However, these singular perspectives introduce an ill-posed nature to the problem, as the reduction from 3D reality to 2D representation inherently results in the loss of spatial data. Conversely, the integration of multiple-camera-views imaging facilitates a well-conditioned approach to interpolate the lost dimension for 3D APE using geometric consistency. Based on the number of camera views used to extract the animal poses, we categorize the 3D APE into two sections: (1) 3D APE from single-camera-view images and videos (Salem et al., 2019; Gosztolai et al., 2021; Dai et al., 2023; Sosa and Hogg, 2023; Shooter et al., 2024; Hu et al., 2023; Zuffi et al., 2017, 2019a; Biggs et al., 2020; Li and Lee, 2021a; Wang et al., 2021a; Rüegg et al., 2022, 2023; Wu et al., 2023a, b; Yang et al., 2022a; Li et al., 2024b), (2) 3D APE from multiple-camera-views images and videos (Yao et al., 2019; Bala et al., 2020; Zhang et al., 2021; Karashchuk et al., 2021; Dunn et al., 2021; Muramatsu et al., 2022; Ebrahimi et al., 2023; An et al., 2023; Han et al., 2024).
Table 7. 3D RGB-based APE methods. All the symbols and column meanings are defined in table 5 and table 6, and any additional symbols are defined here: [Uncaptioned image] is shape model (Zuffi et al., 2017), [Uncaptioned image] is Structured Forests architecture (Dollár and Zitnick, 2015), [Uncaptioned image] is Graph Convolutional Network (GCN), [Uncaptioned image] is Stacked Hourglass architecture based on CNN, [Uncaptioned image] is GAN, [Uncaptioned image] is the combination of Neural Radiance Fields (NeRF), Volumetric Signed Distance Functions (VolSDF) & Differentiable Physics Simulation, [Uncaptioned image] is a framework based on Epipolar Divergence, [Uncaptioned image] is Gaussian-based model, [Uncaptioned image] is Statistical Shape Model (SSM), [Uncaptioned image] is NeRF, [Uncaptioned image] is Semantic Bank of Skinned Models (SBSM), [Uncaptioned image] is feed-forward neural network. [Uncaptioned image] is the Body Mesh Recovery method
Sl Year Methods VI L Features Evaluated On Evaluation Metrics
3D APE from Image & Video - Single-Camera-View Methods:
Skeleton-Only:
1 2019 Salem et al. (Salem et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse Pose Distance Measure (Perona, 2010),
Reconstruction Err.
2 2021 LiftPose3D (Gosztolai et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice, Rats & Macaques MAE
3 2023 Hu et al. (Hu et al., 2023) V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse OKS for 2D Keypoints,
RMSE for 3D Keypoints
4 2023 Dai et al. (Dai et al., 2023) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AcinoSet (Joska et al., 2021), N-MPJPE
SMAL (Zuffi et al., 2017)
5 2023 Sosa et al. (Sosa and Hogg, 2023) I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse PCK@0.05 for 2D,
(Weizmann Data) (Borenstein et al., 2004) Qualitative for 3D
6 2024 DigiDogs/D-Pose (Shooter et al., 2024) I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dogs PCK for 2D&3D, MPJPE for 3D
AMR (Model-Based):
7 2017 SMAL (Zuffi et al., 2017) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Quadruped Scan-to-Mesh Distance,
Keypoint Reprojection Err.,
Silhouette Reprojection Err.
8 2019 Three-D Safari (Zuffi et al., 2019a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Zebra PCK, IoU, Geodesic Distance
9 2020 SMBLD (Biggs et al., 2020) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) IoU, PCK
10 2021 Li et al. (Li and Lee, 2021a) I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020), PCK, IoU
Animal Pose (Cao et al., 2019),
BADJA (Biggs et al., 2018)
11 2021 Birds of a Feather (Wang et al., 2021a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Birds, StanfordExtra (Biggs et al., 2020) PCK, IoU
12 2022 BARC (Rüegg et al., 2022) I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) PCK & IoU for 2D,
Breed Prototype Consistency &
Perceptual Shape Eval. for 3D
13 2023 BITE (Rüegg et al., 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) Scan-to-Mesh Distance,
Semi-Synthetic 3D (Rüegg et al., 2023) Mesh-to-Scan Distance,
PCK, IoU
AMR (Model-Free):
14 2023 DOVE (Wu et al., 2023a) I,V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Bird, Horse, mIoU,
3D Toy Bird Data 3D Chamfer Distance (cm)
15 2023 MagicPony (Wu et al., 2023b) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse-10 (Mathis et al., 2021), PCK,
DOVE Data (Wu et al., 2023a), 3D Chamfer Distance (cm)
COCO (Lin et al., 2014a), & others
16 2022 BANMo (Yang et al., 2022a) I,V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} , [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animals & Humans 3D Chamfer Distance (cm),
F-score
17 2024 3D-Fauna (Li et al., 2024b) I U [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Quadruped Animals PCK@0.1, KT-PCK@0.1
3D APE from Image & Video - Multi-Camera-Views Methods:
18 2019 Monet (Yao et al., 2019) I SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dog, Monkey PCK, Reprojection Err.
19 2020 OpenMonkeyStudio (Bala et al., 2020) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Monkey Median & Mean Err. (cm),
Standard Deviation (cm)
20 2021 GIMBAL (Zhang et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse MPE (mm), RPA-MPE (mm)
21 2021 DAN-NCE (Dunn et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rats, Marmosets Mice PCK, Reconstruction Err.
Rat Pups & Chickadees
22 2021 Anipose (Karashchuk et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse Joint Position Err. (mm)
23 2022 Muramatsu et al. (Muramatsu et al., 2022) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AcinoSet (Joska et al., 2021) RMSE, PCK,
$l^{2}$ norm in pixels
24 2023 3D-UPPER (Ebrahimi et al., 2023) V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice RMSE, Scalar Product,
Pearson’s Correlation
25 2024 MAMMAL (An et al., 2023) V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pigs, Mouse, Dogs MPJPE, IoU,
Avg. Err. on Each Keypoint (cm)
26 2024 SBeA (Han et al., 2024) V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mice, Dogs, Birds RMSE (pixels), Reprojection Err.
4.3.1. Single Camera View
Single-camera-view 3D APE methods reconstruct 3D animal poses from monocular images and videos, which presents a challenge due to the inherent ambiguity in deriving depth from 2D projections. Existing methodologies within this domain can be classified into two categories: (1) skeleton-only methods, which target reconstructing the animal’s skeletal structure. (2) Animal Mesh Recovery (AMR), which recovers the animal’s entire mesh.
Skeleton-only Methods
Skeleton-only methods are dedicated to estimating the 3D coordinates of animal joints for a skeletal representation as their result. Inspired by the advancements in 2D APE, the paradigm of inferring 3D poses from estimated 2D poses, termed 2D-to-3D lifting, has garnered considerable attention. This approach typically leverages strong priors for predicting 3D joint locations from a single-camera view. A notable example is the LiftPose3D framework (Gosztolai et al., 2021), which trains a supervised neural network utilizing 3D ground truth data to lift 2D poses from one camera. The training data is augmented with different camera angles and bone lengths, which allows the network to internalize camera parameters implicitly and accommodate variations in animal sizes. In contrast, recent innovations have endeavoured to estimate camera parameters and construct shape distributions explicitly (Hu et al., 2023). In this paper (Hu et al., 2023), 3D APE is formulated as an optimization problem utilizing a pre-defined mouse skeletal model. This 3D skeleton model imposes a strong prior, addressing the issue of occlusions and acting as a regularization on the over-parameterized joint space. These two methods hinge on extensive 3D annotated datasets, which are hard to apply to wild animals whose data are difficult to collect.
Recent attempts have been made to address this challenge to eliminate the dependence on 3D ground truth through unsupervised 2D-to-3D lifting techniques. For instance, Dai et al. (Dai et al., 2023) have introduced an unsupervised approach that, given a 2D pose, employs a pose generator network to generate a corresponding 3D posture, while a separate camera network estimates the camera rotation. During training, the generated 3D pose is subjected to random re-projections onto camera viewpoints to synthesize a new 2D pose. The synthesized 2D pose is then decomposed into a 3D pose and a camera rotation, based on which consistency losses are imposed in both 3D canonical poses and camera rotations for self-supervised training with geometric consistency. Complementing this, Sosa et al. (Sosa and Hogg, 2023) have proposed a similar methodology which relies on self-supervision with the same geometric consistency constraint for 3D pose estimation. Notably, while Dai et al. (Dai et al., 2023) assume the availability of actual 2D poses for each input image, Sosa et al. (Sosa and Hogg, 2023) demonstrate the feasibility of their approach with minimal assumptions, including unlabeled images and a small set of synthetic 2D poses. The minimalistic requirements and the availability of large volumes of unlabeled data suggest that this method (Sosa and Hogg, 2023) could be readily generalized across diverse animal species. Another 2024 approach to automatically extract the 3D pose from single-camera-view images using only synthetic data for training is DigiDogs (Shooter et al., 2024). This method takes DINOv2 (Oquab et al., 2023) as the base architecture, which is a foundation model known for its robust pre-training capabilities and superior domain adaptation performance. These studies exploit synthetic data to predict 3D joint locations from 2D images, reinforcing the potential of synthetic augmentation in the domain of monocular 3D APE.
Collectively, these methods in single-camera-view skeleton-only 3D APE underscore the dynamic evolution of the field, reflecting a shift towards more generalized and unsupervised learning paradigms that promise to enhance the versatility and applicability of pose estimation across a broader spectrum of animal species.
Animal Mesh Recovery
In the domain of 3D single-camera-view APE from monocular images and videos, this AMR subfield is further categorized into model-based AMR and model-free AMR. The former predominantly utilizes parametric body models to infer the detailed 3D mesh of animals. The latter eliminates the need for a pre-existing parametric template model to recover 3D animal meshes.
One of the most well-known model-based AMR is the Skinned Multi-Animal Linear (SMAL) model (Zuffi et al., 2017), which provides a template for various animal classes based on 3D scanning of deformable toy animals. The SMAL model represents animals using a linear combination of shape bases, which are adapted to capture the diversity of animal forms, alongside a set of pose parameters to articulate the model. This framework serves as a foundation for many model-based AMR works. A notable enhancement of the SMAL model is SMALR (Zuffi et al., 2018), which allows the recovery of detailed 3D shapes of unseen animals and creates textured 3D meshes. SMALST (Zuffi et al., 2019a) further amalgamates real textures with synthetically generated poses, shapes, and backgrounds. Beyond SMALST, the field has seen the development of breed-specific variants of SMAL, such as models for dogs (Rüegg et al., 2023, 2022; Biggs et al., 2020; Li and Lee, 2021a) and birds (Wang et al., 2021a). These models incorporate scaling parameters into the SMAL framework to enable breed-specific fitting. By using the pre-defined 3D shape and pose priors for the target animals, model-based AMR methods aid the animal shape reconstructions in ambiguous cases; the recovered mesh model further provides details for analyzing body size, posture, motion, behaviour, and health conditions. However, when we need to estimate the poses for new animal types without an existing template mesh, data collection for these unseen animals can be time-intensive, thus restricting model-based AMR’s applicability to the diverse animal kingdom.
On the other hand, model-free approaches (Kanazawa et al., 2018; Goel et al., 2020; Wu et al., 2023a, b; Yang et al., 2022a; Li et al., 2024b; Jakab et al., 2024; Yang et al., 2023b) do not rely on pre-existing parametric body models, which offers greater flexibility when we have rare and diverse animals. Many of these methods attempt to learn 3D priors from large-scale internet images and videos in an unsupervised or self-supervised manner, focusing on one category at a time (Goel et al., 2020; Wu et al., 2023a, b; Jakab et al., 2024; Yang et al., 2023b). In contrast, 3D-Fauna (2024) (Li et al., 2024b) introduces a pan-category deformable model capable of reconstructing any quadruped animal instance in a feed-forward manner, pushing towards a more generalizable and unconstrained understanding of animal forms from monocular images. However, the recovered 3D shapes of current model-free approaches often produce overly smooth surfaces, which do not preserve as much detail as the model-based methods; this severely limits their applicability for downstream applications.
4.3.2. Multiple Camera Views
Multiple-camera-views 3D APE methods utilize multiple synchronized cameras to capture different perspectives of animals and reconstruct their 3D poses.
The majority of the existing methods in multiple-camera-views 3D APE are designed to estimate the pose of a single animal in keypoints format (Yao et al., 2019; Günel et al., 2019; Nath et al., 2019; Zhang and Park, 2020; Bala et al., 2020; Zhang et al., 2021; Dunn et al., 2021; Karashchuk et al., 2021; Muramatsu et al., 2022; Ebrahimi et al., 2023). In particular, early methods estimate the 2D poses from multiple independent camera views first; then, these 2D predictions are triangulated to 3D poses (LABEL:fig:multiview). These post hoc triangulation approaches like (Nath et al., 2019) work well when all the joints are visible in all views. However, they are not suited to handle occlusions, which is a scenario that always happens in animal naturalistic environments. To address these limitations, some methods apply skeletal constraints and filtering (Günel et al., 2019; Karashchuk et al., 2021). In detail, Anipose (Karashchuk et al., 2021) benefits significantly from post-processing spatial-temporal filtering, especially in scenarios involving freely moving subjects where occlusion is frequent. However, these methods have yet to achieve robust 3D tracking of freely moving animals unless trained with large numbers of frames. A notable exception is OpenMonkeyStudio (Bala et al., 2020), which can estimate the 3D macaques pose in extensive, unconstrained environments with a network of 62 high-resolution video cameras. However, these methods triangulate 3D locations from 2D estimates; their 2D nature limits their ability to capture and quantify behaviours in 3D accurately. To incorporate advances in 3D, some approaches integrate multiple-camera-views geometric data directly within end-to-end training pipelines (LABEL:fig:multiview2). A notable system, DANNCE (Dunn et al., 2021), utilizes projective geometry to construct a 3D feature space from multiple-camera-views imagery and subsequently trains a 3D CNN to predict the positions of anatomical landmarks from the shared features across cameras. In addition, to tackle the data scarcity in 3D APE, 3D-Upper (2023) (Ebrahimi et al., 2023) uses an unsupervised algorithm based on a Statistical Shape Model (SSM) to estimate 3D poses of freely moving animals; this SSM estimator is robust even when many poses (up to 50%) are contaminated by outliers and missing data. The 3D-Upper APE method does not require labelled training data, which makes it highly scalable and more practical for applications in-the-wild or in situations where manual annotations are impractical.
All of the aforementioned 3D multiple-camera-views approaches are limited to estimating the pose of a single animal, and they only provide poses in keypoints format, which does not preserve as much animal shape information as mesh-based APE methods. MAMMAL (2024) (An et al., 2023) and SBeA (2024) (Han et al., 2024) extend this to enable multi-animal 3D social pose estimation using multi-camera-views. In particular, MAMMAL is designed for large mammals and provides a full 3D surface mesh model compared to the sparse 2D keypoint methods. It captures fine-level surface geometries to understand animal motion involving interactions and occlusions in their complex social behaviours. Moreover, the cross-view matching algorithm further allows MAMMAL to reconstruct 3D poses robustly, even in the presence of keypoint occlusion. Triangulation-based methods are weak in such scenarios and may lead to incomplete or incorrect pose estimations. Another recent method, SBeA, further pushes the trend towards using a few-shot learning strategy to accurately estimate multi-animal 3D poses and identities with minimal manual annotation. SBeA introduces a novel bidirectional transfer learning strategy that enables zero-shot multi-animal identity recognition. It reuses models trained on single-animal data to identify multiple animals, thus minimizing extra identity annotations. The SBeA model employs multiview camera arrays and geometrical constraints to reconstruct 3D multiple animals’ poses effectively, which is hard to attain through single-camera-view APE methods. Lastly, SBeA confers the capability for unsupervised behaviour classification by decomposing 3D trajectories into 2D representations and clustering social behaviour modules. This permits the discovery of subtle social behaviour phenotypes that supervised techniques might miss.
In summary, RGB-based APE methods have advanced significantly from 2017 to 2024. Early approaches, like DeepLabCut (Mathis et al., 2018), focused on 2D keypoint-based supervised methods adapted from HPE for tracking animal joints in controlled environments. As research progressed, interest grew in 3D APE (particularly mesh-based methods), which capture more detailed pose and shape data for downstream tasks like tracking and behaviour analysis. However, the limited annotated data, and variations in animal species and environmental set-up in APE have initially restricted the progress of supervised APE methods. To address these, recent methods have shifted to weakly-supervised, self-supervised, semi-supervised, and unsupervised learning, using synthetic data, transfer learning, and domain adaptation to improve model generalization and tackle data scarcity. Additionally, given that many animals are social creatures, there is a trend toward multi-animal tracking from video data to understand their communication and social structures. These trends culminated in 2024 with the release of SuperAnimal (Shaokai Ye and Mathis, 2024) (2D APE), MAMMAL (An et al., 2023) and SBeA (Han et al., 2024) (3D APE), which offer superior APE performance, generalization across species and diverse environments, reduced dependence on labelled data, and allow for tracking multiple animals simultaneously, setting a new benchmark in the field.
5. Non-RGB-based Unimodal Animal Pose Estimation
In the previous section, we comprehensively reviewed APE methods primarily based on RGB cameras. However, beyond RGB-based approaches, researchers have explored the use of other unimodal methods for analyzing animal behaviour, table 8 summarises the latest APE methods using both non-RGB unimodal (Yigit et al., 2022) and multi-modal approaches, which is discussed in detailed in section 6.3.
IMU
IMU is a small, lightweight, low-cost device that measures an object’s acceleration and angular velocity. Attaching multiple IMUs to limbs and the trunk makes it possible to capture the motion and orientation of the animal in 3D space. Recent work (Yigit et al., 2022) presents a lameness detection and pose estimation method for horse walk and trot locomotion based on the acceleration and gyroscope measurements from IMUs. Initially, they train the network offline using IMU measurements as input sequences, with labelled data provided by the MoCap system. Once the long short-term memory (LSTM) model for gait event and lameness detection is trained, it embeds the IMU measurements into high-dimensional joint angles, which are then fed into the Gaussian Process Dynamic Model (GPDM) to learn a low-dimensional latent representation of the horse’s limb motion. This learned latent manifold dynamics is parameterized with the gait phase variable to predict the horse’s 24 limb joints in real time.
Currently, non-RGB unimodal APE methods have yet to be widely adopted. However, using multi-modality approaches, such as combining RGB cameras with other sensors, has gained more popularity, as discussed in section 6.3.
Table 8. Non-RGB-based unimodal APE methods and multi-modal APE methods. All the symbols and column meanings are defined in table 5, table 6 and table 7, and any additional symbols are defined here: [Uncaptioned image] is GPDM architecture, [Uncaptioned image] is Hierarchical Gaussian Process Latent Variable Model (H-GPLVM), [Uncaptioned image]is Extended Kalman Smoother (EKS), [Uncaptioned image] is Context Models, [Uncaptioned image] is CLIP. [Uncaptioned image] is ego-view. Ther. is an abbreviation for thermal, and Lang. is an abbreviation for language.
Sl Year Methods Sensor Features Evaluated On Metric
Non-RGB-based Unimodal APE Method: IMU-based
1 2022 Yigit et al. (Yigit et al., 2022) IMU [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Horse Mean, Std, VR, IAE & IEV
Multi-Modal APE Methods
2 2017 Patel et al. (Patel et al., 2017) RGB+GPS+ [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cheetah Qualitative for 3D
IMU (video observations)
3 2020 Kearney et al. (Kearney et al., 2020a) RGBD [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] RGBD-Dog (Kearney et al., 2020a), PCK, MPJPE, PA-PCK,
BADJA (Biggs et al., 2018) PCK 3D, PA-PCK 3D,
PA-MPJPE
4 2023 Luo et al. (Luo et al., 2023) RGBD {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cattles, Pigs (Ruchay et al., 2020) MAE, MRE (Luo et al., 2023)
5 2023 OptiPose (Patel et al., 2023) RGBD+Ther. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rodent3D (Patel et al., 2023), PCK, MPJPE,
+Infrared Rat7M (Dunn et al., 2021), AcinoSet (Joska et al., 2021) Reprojection Error
6 2024 WildPose(Muramatsu et al., 2024) RGB+LiDAR [Uncaptioned image], N/A, [Uncaptioned image], [Uncaptioned image] Eagle, Giraffe, Lion Mean & Std
7 2022 Sound of Motion (Li et al., 2022) RGB+Acoustic [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse PCK, IoU, PA-MPJPE
8 2024 CLHOP (Li et al., 2024a) RGB+Acoustic [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image],[Uncaptioned image]} Horse (Li et al., 2024a) PCK@0.1, IoU, PA-MPJPE
9 2023 CLAMP (Zhang et al., 2023) RGB+Lang. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), AP, AR
Animal Pose (Cao et al., 2019)
10 2023 AmadeusGPT (Ye et al., 2023) RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] Mouse & others (Ye et al., 2023) F1 Score
11 2024 AWOL (Zuffi and Black, 2024) RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] StanfordExtra (Biggs et al., 2020), Perceptual Study,
3D Trees CLIP score, Qualitative
6. Multi-Modal Animal Pose Estimation
Researchers have begun to explore the integration of multi-sensor and multi-modality to improve the accuracy, robustness, and ecological validity of APE algorithms (Patel et al., 2017; Kearney et al., 2020a; Li et al., 2022; Luo et al., 2023; Patel et al., 2023; Zhang et al., 2023; Ye et al., 2023; Muramatsu et al., 2024; Deng, 2023). In this section, we examine the motivation behind multi-modal systems, review existing multi-modal HPE and APE techniques, discuss how they tackle the limitations in unimodal APE methods, and highlight applications demonstrating the potential of integrating diverse sensors and modalities for APE, as shown in table 8.
6.1. Multi-Modal System in Human Pose Estimation
To inspire the use of multi-modal systems for APE, we draw insights from recent multi-modal HPE studies. For instance, (Zheng et al., 2022; Zanfir et al., 2023) combine data from visual RGB cameras with depth information obtained from RGBD and LiDAR sensors in autonomous driving to increase the robustness of 3D HPE in complex environments. The mRI dataset (An et al., 2022) integrates mmWave radar, RGBD cameras, and IMUs and provides a benchmark to support 3D HPE sensor fusion research. Another study (An et al., 2022) introduces the mRI dataset, a multi-modal dataset for 3D HPE that integrates data from mmWave radar, RGBD cameras, and IMUs. These studies highlight the benefits of multi-modal integration in HPE and further inspire us to explore similar approaches for APE.
6.2. Motivation for Multi-Modal System in Animal Pose Estimation
Unimodal APE faces challenges like occlusions and limited perspectives, which multi-modal systems can overcome by combining complementary data to improve accuracy and understanding of animal behaviour (Deng, 2023). For example, multi-modal systems enhance APE resilience by compensating for sensor failures, such as using thermal in low-light conditions when RGB cameras struggle. In addition, RGB-based cameras excel in capturing detailed visual information but may falter with occlusions, while some mmWave sensors can see through walls, and LiDAR can provide valuable 3D spatial data; these significantly help us to expand the coverage area and reduce blind spots. Beyond visual data, integrating IMUs and audio sensors further enriches animal behaviour analysis by capturing motion and auditory cues, which makes multi-modal systems essential for a more robust APE and a better understanding of dynamic environments.
6.3. Existing Multi-Modal Animal Pose Estimation Techniques
In recent years, there has been growing interest in the development of multi-modal APE systems (Kearney et al., 2020a; Luo et al., 2023; Patel et al., 2017; Li et al., 2022; Patel et al., 2023; Zhang et al., 2023; Ye et al., 2023; Muramatsu et al., 2024; Li et al., 2024a; Zuffi and Black, 2024). It involves combining information from different sensors to derive a unified representation of animal pose. Compared to the significant progress in unimodal APE, there is limited research on using multiple sensors and modalities for enhanced accuracy in APE, and we aim to review all of the existing multi-modal APE studies in this section.
Animal-borne cameras + GPS + IMUs (Ego-view)
One notable example of a multi-modal system using animal-borne cameras, GPS, and IMU is the work on tracking the cheetah tail (Patel et al., 2017) in ego-view. Ego-view 3D APE, though less explored, presents a novel perspective by placing cameras on the animals themselves, offering a first-person viewpoint. In this study (Patel et al., 2017), researchers attached camera and sensor equipment to cheetahs to obtain motion data from the animal’s viewpoint. Using kinematic models, they can estimate the 3D pose of the cheetah’s tail, including its position, orientation, and movement in 3D space. This research demonstrates the potential of multi-modal ego-view APE methods in capturing the animal’s pose and movement directly from the animal’s perspective, which helps us gain insights into the animal’s behaviour, locomotion, and agility so we can have a better understanding of the cheetah’s hunting strategies, balance control, and communication signals. Please note that the ego-view 3D APE study is currently limited, and the cheetah tail tracking study opens a door for it.
RGB + Depth
One promising sensor modality is the RGBD-based pose estimation, as exemplified by (Kearney et al., 2020a). The pose estimation model is a two-stage approach for predicting 3D canine poses from RGBD data. In the first stage, a stacked hourglass network is employed to predict 2D joint locations from RGB images, with multiple stacked layers that allow for iterative refinement of joint positions. To generalise the model across various canine shapes and sizes, this study trains the network on both real and synthetic data. In the second stage, a Hierarchical Gaussian Process Latent Variable Model (H-GPLVM) is used to lift the predicted 2D joint positions to 3D, while ensuring the output poses are anatomically plausible by imposing constraints on the possible configurations. The H-GPLVM leverages the structure of the canine skeleton to refine the pose estimations, which further reduces errors and improves robustness in complex poses or challenging viewpoints. This model can extract dog poses effectively even when there is significant variability in appearance and motion. In 2023, there is another RGBD-based APE approach (Luo et al., 2023) that enhances livestock body measurements. It first preprocesses point cloud data to detect 3D keypoints, then aligns a template mesh (i.e. a statistical shape model) with the scan data, and uses non-rigid registration and poses normalisation to standardise the animals’ poses into a canonical form. These help to address the common challenges such as noise, missing data, and inconsistencies caused by livestock movement. Even though RGBD is a promising approach, its popularity is limited by dataset availability and hardware constraints; hence, it has not yet been widely adopted within the APE community compared to RGB-based methods.
RGB + Depth + Thermal infrared
The OptiPose study (Patel et al., 2023) introduces a novel approach for animal pose tracking that leverages a multi-modal dataset and token-based pose optimisation. To capture accurate APE, researchers collected a 3D multi-modal dataset comprising RGB, depth, and thermal infrared data. RGB data provides high-resolution colour imagery that helps identify detailed animal features, allowing for better 2D keypoint extraction. Depth data provides spatial information about the distance between the camera and the animal, which is important for reconstructing the 3D pose from multiple 2D views. Finally, thermal infrared data measures the heat emitted by the animal, which is particularly useful in low-light or occluded conditions where RGB data struggled. They propose a token-based optimisation method that uses multi-modal data to refine the pose estimation results. The study demonstrates the potential of integrating multiple sensor modalities for improved animal pose tracking and motivates further exploration of multi-modal approaches in animal behaviour analysis.
RGB + LiDAR
A novel approach to APE uses LiDAR and RGB sensors, as exemplified by the WildPose (Muramatsu et al., 2024) system. Unlike traditional methods that rely on RGB images, WildPose utilises a combination of a zoom-lens camera and a solid-state LiDAR to capture both 2D and 3D data of animals in their natural habitats. The WildPose system is non-invasive and can capture creatures in a long-range distance (ranging from 2-200 meters) without disturbing the animals. Moreover, LiDAR adds depth information about the animals and their surrounding environments to provide a more comprehensive understanding of the animal’s morphology and movements. Despite these advantages, the high cost and complexity of LiDAR systems may limit their accessibility and scalability. Additionally, the accuracy and quality of the LiDAR data can be affected by environmental factors such as weather conditions and lighting.
RGB + Acoustic
The Sound of Motion (Li et al., 2022) and CLHOP (Li et al., 2024a) present a multi-modal approach for estimating horse motion using video and audio data. Audio features are extracted from sounds like horse hoof strikes and are processed in parallel with visual features. These audio features are converted into Mel spectrograms and passed through a network that helps predict 3D pose parameters. These studies highlight the value of incorporating audio to capture fine-grained details of animal movements. Apart from APE, audiovisual data have also been applied in other animal motion analyses such as horse behaviour detection (Nunes et al., 2021) and primate action recognition (Bain et al., 2021).
RGB + Language
As we advance, the APE trend is moving towards using vision-language-based APE methods (Zhang et al., 2023; Zuffi and Black, 2024). The CLAMP (2023) (Zhang et al., 2023) method uses language to enhance learning processes through the description provided, thus increasing the capability of generalising models about new and unseen animal poses. Following that, the AWOL (2024) (Zuffi and Black, 2024) approach creates a model that takes input from Contrastive Language-Image Pretraining (CLIP) encodings of either text or images and maps them to 3D shape parameters. This learns the mapping between the CLIP latent space and the shape space of SMAL, allowing it to generate realistic 3D animal mesh as a final output. This exploits the strengths of vision-language models in blending visual and textual information into a more robust framework, advancing the APE field of work.
In conclusion, the above studies provide insights and inspirations for future work, which shows a trend towards integrating multiple sensors and modalities, such as RGB, depth, IMUs, LiDAR, thermal, acoustic, and language, to enhance the accuracy, robustness, and contextual understanding of APE.
7. Future Work
Can we reduce or eliminate the effort required for ground-truth annotation in APE using unsupervised learning and meta-learning?
Unsupervised learning and meta-learning hold great potential to significantly reduce or even eliminate the effort required for ground-truth annotation in APE. It will be worth exploring how unsupervised learning models can leverage inherent structures in the data, such as consistency across multiple views, temporal coherence in video sequences, or self-supervised signals like motion cues. By exploiting such patterns, unsupervised APE models can hopefully learn from vast amounts of unlabeled data, which is particularly useful in wildlife observation where labelled data is scarce. Future APE work will also benefit from meta-learning capabilities. Meta-learning focuses on teaching models to learn how to learn, enabling them to generalise quickly to new species or environments with minimal labelled data. In APE, meta-learning will be able to train models on a limited set of labelled animal poses to generalise to entirely new species or different environmental conditions using only a small number of annotated examples. Lastly, future work will also combine both approaches to leverage large amounts of unlabeled data for initial training and then rapidly adapt to specific tasks with minimal supervision. This hybrid approach would dramatically reduce the need for extensive annotations while improving the model’s robustness and generalisation to new tasks.
How can we extend multi-modal frameworks to perform multiple-animal pose estimation in the scene?
Currently, the use of multi-modal frameworks for a single animal is limited, and for multi-animal, minimal work exists (i.e. OptiPose (Patel et al., 2023) that can perform APE on two rodents from a multi-modal system). However, the future work will include multi-modal multi-animals pose estimation in-the-wild for various species. Going forward, we will see a fusion of sensors and modalities for handling occlusions and provide additional context when animals interact closely. For example, depth data will help to differentiate between overlapping animals, while thermal data will assist in distinguishing animals from their surroundings in low-light conditions. In addition, we may also use multi-modal data to combine animal appearance and other information such as their shape, weight, and gaits together to resolve individual animal identity, which is very important for agriculture and conservation. In future, work will move towards scene object interaction and reconstruction. By using multi-modal data, we will reconstruct both the animals and their surrounding objects by understanding context-specific interactions, such as animals navigating around obstacles, interacting with other animals, or using objects (e.g. chimpanzees nut cracking in-the-wild). For instance, LiDAR or depth sensors will capture the 3D geometry of the environment, and when combined with RGB data, the system will reconstruct both the animal’s poses and the objects they interact with. The future will also move towards having real-time processing, which will allow us to track pose and behaviour in real-time, enabling effective health monitoring to protect endangered species.
Can we perform language-driven animal articulation with advanced APE methods?
With the rise of models like CLIP, the trend is moving towards more vision-language hybrid approaches within the CV community. So far, there is no work available where it is possible to control and articulate 3D animal models using natural language commands in an unsupervised manner, but this will likely happen in the future. By integrating language-driven models with advanced APE techniques, we can achieve sophisticated text-based control for animal poses. Similar to how text prompts can control image generation in models like CLIP, future APE methods could allow users to articulate an animal’s pose using natural language instructions. For instance, commands like ”raise the horse’s front right leg” or ”make the chimpanzee sit” could be translated into precise joint movements or skeletal adjustments in a 3D model. Moreover, mesh-based APE methods, when integrated with language models, could allow for real-time adjustments in the animal’s anatomy and surface details with commands like ”increase muscle tension” or ”broaden the horse’s stance” and this could instantly alter both the pose and physical appearance of the 3D mesh model in real-time. These advancements will be beneficial in applications like virtual reality or animations, where controlling animal movements dynamically with text could simplify workflows. It also has the potential for behavioural simulations; for example, researchers or animators can generate lifelike animal behaviours through high-level language inputs without manually defining joint movements.
Acknowledgements.
We thank Andrew Zisserman for his excellent suggestions and for reviewing the paper. We would also like to thank Jiaxing Zhong, Ta-Ying Cheng, Xinyu Hou, Qian Xie, Shaokai Ye, Daniel Schofield, Eldar Insafutdinov, Tomas Jakab and Anjun Hu for the insightful discussion on pose estimation. We would like to thank (Yunus et al., 2024) for providing the symbols. Additionally, Qianyi would like to thank Jian Zhou for his loving support. During this project, Qianyi Deng was supported by the UKRI’s EPSRC ACE-OPS grant (EP/S030832/1), and Oishi Deb was supported by the UKRI’s EPSRC AIMS CDT grant (EP/S024050/1).

### v2

@@
-Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis
-Abstract.
-Animal pose estimation (APE) aims to locate the animal body parts using a diverse array of sensor and modality inputs, which is crucial for research across neuroscience, biomechanics, and veterinary medicine.
-By evaluating 178 papers since 2013, APE methods are categorised by sensor and modality types, learning paradigms, experimental setup, and application domains, presenting detailed analyses of current trends, challenges, and future directions in single- and multi-modality APE systems.
-The analysis also highlights the transition between human and animal pose estimation.
+
+Abstract
+Animal pose estimation (APE) aims to locate the animal body parts using a diverse array of sensor and modality inputs (e.g.
+RGB cameras, LiDAR, infrared, IMU, acoustic and language cues), which is crucial for research across neuroscience, biomechanics, and veterinary medicine.
+By evaluating 176 papers since 2011, APE methods are categorised by their input sensor and modality types, output forms, learning paradigms, experimental setup, and application domains, presenting detailed analyses of current trends, challenges, and future directions in single- and multi-modality APE systems.
+The analysis also highlights the transition between human and animal pose estimation, and how innovations in APE can reciprocally enrich human pose estimation and the broader machine learning paradigm.
 Additionally, 2D and 3D APE datasets and evaluation metrics based on different sensors and modalities are provided.
 A regularly updated project page is provided in this GitHub link.
-Survey of animal pose estimation, human pose estimation, 2D and 3D pose estimation, unimodal and multi-modal pose estimation, deep learning-based pose estimation, pose estimation datasets, pose estimation metrics
-copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; 2024; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: General and reference Surveys and overviews
-1. Introduction
-Animal pose estimation (APE) identifies the spatial configuration of an animal’s body parts using a diverse array of sensor and modality inputs, which is important in numerous disciplines (Karashchuk et al., 2021).
-For instance, neuroscientists rely on accurate measurements of animal movement to correlate with brain dynamics (Mathis et al., 2018; Krakauer et al., 2017; Pereira et al., 2020; Luxem et al., 2022).
-Biomechanists analyse the movement of specific animal body structures to comprehend biomechanical phenomena, such as mechanical properties, stress-strain relationships, and functionality of different body structures (Patel and Braae, 2013; Joska et al., 2021; Patel and Braae, 2014).
+keywords:
+Animal pose estimation, human pose estimation, 2D and 3D pose estimation, unimodal and multi-modal pose estimation, multi-modal learning, sensor fusion, deep learning, pose estimation datasets and metrics, literature survey.
+1 Introduction
+Animal pose estimation (APE) identifies the spatial configuration of an animal’s body parts using a diverse array of sensor and modality inputs, which is important in numerous disciplines [55].
+For instance, neuroscientists rely on accurate measurements of animal movement to correlate with brain dynamics [90, 62, 113, 85].
+Biomechanists analyse the movement of specific animal body structures to comprehend biomechanical phenomena, such as mechanical properties, stress-strain relationships, and functionality of different body structures [106, 53, 107].
 In addition, ecologists monitor wildlife behaviours to develop more effective conservation strategies.
-Similarly, ethologists examine detailed animal gestures and postures to study intra- and inter-species communication and societal structures (Zuffi et al., 2019a; Ng et al., 2022a).
-Moreover, veterinary practitioners employ precise APE techniques to diagnose health conditions (Farahnakian et al., 2021; Hu et al., 2023) and assess recovery post-treatment, thereby mitigating potential disorders (Fang et al., 2021).
+Similarly, ethologists examine detailed animal gestures and postures to study intra- and inter-species communication and societal structures [176, 103].
+Moreover, veterinary practitioners employ precise APE techniques to diagnose health conditions [34, 47] and assess recovery post-treatment, thereby mitigating potential disorders [32].
 In this paper, we analyse the current state of deep learning-based APE in both single- and multi-modality approaches.
@@
 To provide context, we first overview the extensive research conducted in Human Pose Estimation (HPE).
-1.1. Human Pose Estimation
+1.1 Human Pose Estimation
 The advancements in HPE provides insights to develop the APE methods due to the skeleton similarity between humans and animals.
@@
 Modern HPE methods are mainly RGB-based and predominantly employ convolutional neural networks (CNNs).
-In the early stages, models like VGG (Simonyan and Zisserman, 2014), ResNet-50 (He et al., 2016), ResNet-101 (He et al., 2016), and ResNet-152 (He et al., 2016) were employed as backbones for HPE from images or videos.
-The stacked hourglass network (Newell et al., 2016) then presented a novel architecture characterized by repeated bottom-up and top-down processing.
+In the early stages, models like VGG [129], ResNet-50 [46], ResNet-101 [46], and ResNet-152 [46] were employed as backbones for HPE from images or videos.
+The stacked hourglass network [102] then presented a novel architecture characterized by repeated bottom-up and top-down processing.
 This design aimed to capture and consolidate features at multiple resolutions, enabling precise localization of keypoints across scales.
-Subsequent innovations, such as the High-Resolution Network (HRNet) (Wang et al., 2020) and its variants (Cheng et al., 2020; Yu et al., 2021a), have further advanced spatial information preservation by maintaining high-resolution representations throughout the entire network.
-Recently, transformer-based methods (Yang et al., 2021; Yuan et al., 2021) have integrated attention mechanisms to capture long-range dependencies and spatial correlations among joints and appearances, expanding the possibilities of HPE.
-Beyond vision-based HPE, alternative sensors like inertial measurement units (IMUs) (Huang et al., 2020), millimeter-wave (mmWave) radar (Sengupta et al., 2020; Xue et al., 2021; Xie et al., 2023) and thermal cameras (Smith et al., 2023; Chen et al., 2020) have been explored, tailoring HPE to diverse use cases.
-Additionally, multi-sensor and multi-modal datasets (Ma et al., 2024; Yan et al., 2023; An et al., 2022) and approaches (Fürst et al., 2021; Zheng et al., 2022; An et al., 2022) have emerged as an innovative step over unimodal HPE to overcome individual sensor and modality limitations, offering a synergistic fusion of data that promises increased robustness and precision, especially in scenarios marked by occlusions, changing lighting conditions, and dynamic environments.
-1.2. Definition of Animal Pose Estimation
+Subsequent innovations, such as the High-Resolution Network (HRNet) [142] and its variants [18, 163], have further advanced spatial information preservation by maintaining high-resolution representations throughout the entire network.
+Recently, transformer-based methods [156, 165] have integrated attention mechanisms to capture long-range dependencies and spatial correlations among joints and appearances, expanding the possibilities of HPE.
+Beyond vision-based HPE, alternative sensors like inertial measurement units (IMUs) [48], millimeter-wave (mmWave) radar [126, 151, 149] and thermal cameras [130, 17] have been explored, tailoring HPE to diverse use cases.
+Additionally, multi-sensor and multi-modal datasets [86, 152, 3] and approaches [35, 172, 3] have emerged as an innovative step over unimodal HPE to overcome individual sensor and modality limitations, offering a synergistic fusion of data that promises increased robustness and precision, especially in scenarios marked by occlusions, changing lighting conditions, and dynamic environments.
+1.2 Definition of Animal Pose Estimation
 APE, as discussed in this review, is defined as the task of identifying and localizing body joints of terrestrial mammals using a diverse array of sensor inputs.
-This task results in the reconstruction of the animal’s pose, depicted as either a skeletal representation marked by keypoints (LABEL:fig:Skeleton_Pose_Representations) or as a comprehensive body mesh (LABEL:fig:Body_Mesh_Reconstructions).
-Figure 2. Transfering human pose estimation knowledge to proximal animal classes (e.g. Chimpanzees) for their pose estimation (Sanakoyeu et al., 2020).
-[…]
-Chiampanzee and Human
+This task results in the reconstruction of the animal’s pose, depicted as either a skeletal representation marked by keypoints (Figure 1(a)) or as a comprehensive body mesh (Figure 1(b)).
+(a) Keypoints Pose Representations. (a) A 2D keypoints pose representation example adapted from [100]. (b) A 3D keypoints pose representation example adapted from [53].
+(b) Body Mesh Reconstructions.
+(a) Original RGB image.
+(b) Reconstructed mesh for (a) using the 3D model-based body mesh reconstruction adapted from [178].
+(c) Target RGB image.
+(d) Reconstructed mesh for (c) applying the 3D model-free body mesh reconstruction method from MagicPony [148].
+Figure 1: Different animal pose estimation methodologies.
+Figure 2: Transfering human pose estimation knowledge to proximal animal classes (e.g. Chimpanzees) for their pose estimation [124].
 Exploring APE through various sensory and modality inputs is important as animals may reside in diverse habitats which require different sensors for data collection.
@@
 For example, rodents are more active in the dark, and standard RGB cameras become inadequate due to poor lighting conditions.
-Consequently, infrared sensors have emerged as suitable tools for capturing the nocturnal activities of animals (Wang et al., 2018; Patel et al., 2023).
+Consequently, infrared sensors have emerged as suitable tools for capturing the nocturnal activities of animals [145, 109].
 Additionally, mmWave radar can detect movements but suffers from the scarcity of its generated point clouds; LiDAR adds value by providing depth information.
-Inertial sensors (Patel et al., 2017) can also track movement and orientation, offering invaluable insights but requiring on-animal attachment.
-In specific scenarios, even auditory cues (Li et al., 2022) can indicate animal movement.
+Inertial sensors [108] can also track movement and orientation, offering invaluable insights but requiring on-animal attachment.
+In specific scenarios, even auditory cues [67] can indicate animal movement.
 This paper also explores the potential of using multiple sensors (i.e.
@@
 There are two representations for the final estimated animal poses regardless of our sensors or modalities of the APE input data.
-The first encompasses either two-dimensional (2D) or three-dimensional (3D) keypoint pose representations (Karashchuk et al., 2021; Li and Lee, 2023; Nath et al., 2019; Desai et al., 2023) (LABEL:fig:Skeleton_Pose_Representations).
+The first encompasses either two-dimensional (2D) or three-dimensional (3D) keypoint pose representations [55, 70, 100, 25] (Figure 1(a)).
 In this category, the animal’s anatomy is represented as a tree of interconnected keypoints.
 These keypoints present animal body joints, and their interconnections reflect the natural relationship between these joints, providing a skeletal representation of the animal’s posture.
-The second method, body mesh reconstruction (Zuffi et al., 2017; Rüegg et al., 2022; Wu et al., 2023b, a; Yang et al., 2022a), presents a 3D view of the animal’s morphology (LABEL:fig:Body_Mesh_Reconstructions).
-Within this framework, we differentiate between model-based and model-free reconstructions.
+The second method, body mesh reconstruction [178, 121, 148, 153, 147], presents a 3D view of the animal’s morphology (Figure 1(b)).
+Within this framework, we distinguish between model-based and model-free reconstructions.
 The model-based technique (e.g.
-SMAL (Zuffi et al., 2017)) utilizes supervision from the predefined template shape of the creature learned from 3D toy animal scans.
+SMAL [178]) uses supervision from the predefined template shape of the creature learned from 3D toy animal scans.
 This template is then refined by optimizing the shape and pose parameters.
 In particular, the shape parameters capture variations in body metrics (e.g.
-proportions, height, and weight), while the pose parameters address the observable deformations due to movement.
-The model-based approach can generate detailed models, but collecting the requisite data for new types of animals can be time-intensive, restricting their applicability to rare species.
+proportions, height, and weight), while the pose parameters address observable deformations due to movement.
+The model-based approach can generate detailed models, but collecting the necessary data for new types of animals can be time-intensive, restricting their applicability to rare species.
 In contrast, the model-free approach operates without any supervision from the predefined template shape of the creature (e.g.
-MagicPony (Wu et al., 2023b), DOVE (Wu et al., 2023a) and 3D-Fauna (Li et al., 2024b)), which are more capable when estimating the animal poses for the diverse animal species.
-The transition from HPE to APE builds on HPE frameworks.
-Due to skeletal similarities between humans and certain animals like quadrupeds, HPE models for skeletal keypoint detection (Wang et al., 2020; Insafutdinov et al., 2016; Deb et al., 2023) and body mesh reconstruction (Loper et al., 2023) can be pertinent to APE, with modifications to cater to distinct anatomical and dynamic variations across animal species (Li and Lee, 2023; Rashid et al., 2017; Zuffi et al., 2017; Rüegg et al., 2022).
-One notable example is (Sanakoyeu et al., 2020), which involves transferring human DensePose knowledge to proximal animal classes (fig.
-2).
+MagicPony [148], DOVE [147] and 3D-Fauna [74]), which are more capable when estimating the animal poses for the diverse animal species.
+The transition from HPE to APE is based on HPE frameworks.
+Due to skeletal similarities between humans and certain animals like quadrupeds, HPE models for skeletal keypoint detection [142, 49, 21] and body mesh reconstruction [82] can be pertinent to APE, with modifications to cater to distinct anatomical and dynamic variations across animal species [70, 116, 178, 121].
+One notable example is [124], which involves transferring human DensePose knowledge to proximal animal classes (Figure 2).
 This approach leverages the DensePose-COCO dataset, which contains dense human body annotations, and adapts it to animal species like chimpanzees.
@@
 In addition, innovations in APE that allow adaptation to unseen animals and environmental changes can also improve the robustness of HPE, particularly in challenging scenarios like emergency response, where the surrounding environment changes rapidly.
-However, bridging these two fields requires adaptation of knowledge and consideration of the distinct challenges and ethical issues in human and animal studies, which we will discuss next.
-1.3. Challenges in Human Pose Estimation and Animal Pose Estimation
+However, bridging these two fields requires adaptation of knowledge and consideration of the distinct challenges and ethical issues in human and animal studies.
+Ethical and responsible practices [59] are especially critical in HPE methods development, and similar approaches can be applied to the APE methods as well.
+1.3 Challenges in Human and Animal Pose Estimation
 HPE and APE encounter several common challenges that can impact the accuracy of associated algorithms.
@@
 Different animal species exhibit significant domain discrepancies in physical characteristics, body structures, and postures.
-Even within the same species, there can be a large variety in body size, shape, texture, motion, and posture, presenting difficulties in creating universally applicable models (Rüegg et al., 2022).
+Even within the same species, there can be a large variety in body size, shape, texture, motion, and posture, presenting difficulties in creating universally applicable models [121].
 For example, the pelage of porcupines can vary a lot, and this is often more pronounced and thicker in animals than in humans’ hair, which can make tracking skeleton segments tricky.
-Furthermore, many animals exhibit natural camouflage, travel in herds, interact frequently, or share similar appearances within their species (Zuffi et al., 2019a).
+Furthermore, many animals exhibit natural camouflage, travel in herds, interact frequently, or share similar appearances within their species [176].
 These characteristics make it hard to distinguish individual animals and accurately estimate their poses as they blend into the background or appear similar to one another.
@@
 Vegetation or terrain features in the natural environments may obstruct the view of animals.
-One example can be found in the AcinoSet dataset (Joska et al., 2021), where the cheetahs’ paws were excluded as the grass occluded them in most of the videos, making it difficult to estimate animal poses accurately.
+One example can be found in the AcinoSet dataset [53], where the cheetahs’ paws were excluded as the grass occluded them in most of the videos, making it difficult to estimate animal poses accurately.
 Moreover, changing lighting conditions can also heavily affect an animal’s appearance, especially when the sun’s position is moving in outdoor environments.
@@
 Also, nocturnal species may require specialized sensors like infrared or thermal sensors to capture their movements and poses.
-Moreover, there are very limited annotated APE datasets (Yu et al., 2021b).
-Unlike HPE, which predominantly leans on supervised (Wang et al., 2020; Cheng et al., 2020) and semi-supervised learning methodologies (Ukita and Uematsu, 2018; Pavllo et al., 2019; Mitra et al., 2020) sustained by large-scale human pose datasets such as MSCOCO (Lin et al., 2014a) and MPII (Andriluka et al., 2014), APE is hindered by a lack of such datasets, especially for endangered species and wildlife (Zuffi et al., 2019a).
+Moreover, there are very limited annotated APE datasets [164].
+Unlike HPE, which predominantly leans on supervised [142, 18] and semi-supervised learning methodologies [138, 110, 92] sustained by large-scale human pose datasets such as MSCOCO [75] and MPII [4], APE is hindered by a lack of such datasets, especially for endangered species and wildlife [176].
 Also, it is often impractical to place animals in controlled environments to obtain ground truth poses, although zoos provide a semi-structured environment towards data collection.
-Therefore, unsupervised approaches that do not rely on labelled data have become popular in APE (Wu et al., 2023b; Li and Lee, 2021b).
-1.4. Applications of Animal Pose Estimation
+Therefore, unsupervised approaches that do not rely on labelled data have become popular in APE [148, 69].
+1.4 Applications of Animal Pose Estimation
 HPE has gained significant attention recently due to its applications in many domains, such as human-computer interaction, sports analytics, and healthcare.
@@
 Despite this, APE is critically essential in various applications, as explained below and summarised in Table 1.
-1.4.1. Tracking
-APE has been applied in various tracking contexts; for example, in the CATER (Haalck et al., 2023) framework, APE is combined with environment reconstruction to track ants in complex natural environments, providing fine-scale motion trajectories even in visually challenging conditions.
-The SMART-BARN (Nagy et al., 2023) system uses APE to track multiple animal species simultaneously in a large 3D arena and achieve accurate tracking and behavioural analysis in indoor and naturalistic environments.
-In the FaceMap (Syeda et al., 2024) framework, APE is used for tracking the orofacial movements of mice, enabling detailed analysis of their facial expressions and related neural activities.
-For honey bees, APE tracked an entire colony over extended periods, capturing complex social behaviours and interactions without needing physical tags, enhancing our understanding of colony dynamics and individual behaviours (Bozek et al., 2021).
-Table 1. Table listing various applications of animal pose estimation. Behav. denotes Behaviour Analysis.
+Tracking: APE has been applied in various tracking contexts; for example, in the CATER [42] framework, APE is combined with environment reconstruction to track ants in complex natural environments, providing fine-scale motion trajectories even in visually challenging conditions.
+The SMART-BARN [98] system uses APE to track multiple animal species simultaneously in a large 3D arena and achieve accurate tracking and behavioural analysis in indoor and naturalistic environments.
+In the FaceMap [136] framework, APE is used for tracking the orofacial movements of mice, enabling detailed analysis of their facial expressions and related neural activities.
+For honey bees, APE tracked an entire colony over extended periods, capturing complex social behaviours and interactions without needing physical tags, enhancing our understanding of colony dynamics and individual behaviours [14].
+Table 1: Table listing various applications of animal pose estimation. The paper citation is available in column Year.
 Sl Year APE Paper Animal Classes Domain
-1 2022 I-MuPPET (Urs Waldmann and Goldlücke, 2022) Pigeon Tracking
-2 2023 Bozek et al. (Bozek et al., 2021) Honey Bee Tracking
-3 2023 CATER (Haalck et al., 2023) Desert Ants Tracking
-4 2023 SMART-BARN (Nagy et al., 2023) Insects, Birds, Tracking
-Mammals
-5 2023 3D-POP (Naik et al., 2023) Birds Tracking
-6 2023 Facemap (Syeda et al., 2024) Mice Tracking
-7 2023 DeepWild (Wiltshire et al., 2023) Chimpanzees, Tracking
+1 2019 Bain et al. Chimpanzee Tracking
+2 2021 Bozek et al. Honey Bee Tracking
+3 2022 I-MuPPET Pigeon Tracking
+4 2023 CATER Desert Ants Tracking
+5 2023 SMART-BARN Insects, Mammals, Tracking
+Birds
+6 2023 3D-POP Birds Tracking
+7 2023 DeepWild Chimpanzees, Tracking
 Bonobos
-8 2024 3D-MuPPET (Urs Waldmann and Kano, 2024) Pigeon Tracking
-9 2023 Stathopoulos et al. (Stathopoulos et al., 2023) Sheep, Cows, Bear Articulation
+8 2024 Facemap Mice Tracking
+9 2024 3D-MuPPET Pigeon Tracking
+10 2023 Stathopoulos et al. Sheep, Cows, Bear, Articulation
 Giraffe, Horses
-10 2022 Sheppard et al. (Sheppard et al., 2022) Mouse Behav.
-11 2023 Voloh et al. (Voloh et al., 2023) Macaque Behav.
-12 2023 BARN (Yang et al., 2023a) Macaque Behav.
-13 2023 Koger et al. (Koger et al., 2023) Gelada Monkeys Behav.
-14 2023 AmadeusGPT (Ye et al., 2023) Mammals Behav.
-15 2024 Testard et al. (Testard et al., 2024) Macaque Behav.
-1.4.2. Articulation
-The authors of (Stathopoulos et al., 2023) explore how APE can be used to understand animal articulation by predicting their 3D shapes from monocular images.
+11 2022 Sheppard et al. Mouse Behaviour
+12 2023 Voloh et al. Macaque Behaviour
+13 2023 BARN Macaque Behaviour
+14 2023 Koger et al. Gelada Monkeys Behaviour
+15 2023 AmadeusGPT Mammals Behaviour
+16 2024 Testard et al. Macaque Behaviour
+Articulation: The authors of [133] explore how APE can be used to understand animal articulation by predicting their 3D shapes from monocular images.
 They demonstrate that training a 2D keypoint estimator on a small set of labelled images (50-150) can generate pseudo-labels on a more extensive set of unlabeled web images.
 Then, these pseudo-labels are used to train models for 3D shape reconstruction of various articulated quadrupeds.
-1.4.3. Behaviour Analysis
-In agriculture, animal behaviour analysis based on APE plays an important role by enabling effective livestock management and health surveillance, which in turn provides continuous monitoring of movement and also early detection of animal’s illness, thereby ensuring optimal animal welfare (Fang et al., 2021; Russello et al., 2022; Farahnakian et al., 2021; Gong et al., 2022; Doornweerd et al., 2021; Fang et al., 2022).
-Recently, more APE methods have been applied to study the behaviour of more species, such as macaques (Voloh et al., 2023; Yang et al., 2023a; Testard et al., 2024), Gelada monkeys (Koger et al., 2023), and mice (Sheppard et al., 2022).
-Since 2023, there has been a notable shift towards integrating language prompts with APE in behaviour analysis research, exemplified by developing models like AmadeusGPT (Ye et al., 2023).
+Behaviour Analysis: In agriculture, animal behaviour analysis based on APE plays an important role by enabling effective livestock management and health surveillance, which in turn provides continuous monitoring of movement and also early detection of animal’s illness, thereby ensuring optimal animal welfare [32, 122, 34, 37, 28, 33].
+Recently, more APE methods have been applied to study the behaviour of more species, such as macaques [139, 155, 137], Gelada monkeys [61], and mice [127].
+Since 2023, there has been a notable shift towards integrating language prompts with APE in behaviour analysis research, exemplified by developing models like AmadeusGPT [161].
 This combination has opened up new possibilities for more comprehensive interpretations of animal behaviour.
-1.5. Scope of this Survey
+1.5 Scope of this Survey
 We aim to provide a systematic review of APE, focusing on developments in deep learning-based methodologies for both 2D and 3D scenarios.
@@
 Several existing reviews have been reported on HPE and APE separately.
-Prior surveys (Zheng et al., 2020; Ji and Liu, 2009; Sarafianos et al., 2016; Khan and Wan, 2018; Gong et al., 2016; Liu et al., 2022a; Song et al., 2021; Dubey and Dixit, 2023; Wang et al., 2021b; Moeslund and Granum, 2001; Moeslund et al., 2006; Poppe, 2007) have primarily concentrated on vision-based human motion capture, encompassing aspects of pose estimation, motion tracking, and action recognition.
+Prior surveys [171, 51, 125, 57, 38, 80, 131, 29, 143, 93, 94, 115, 173] have primarily concentrated on vision-based human motion capture, encompassing aspects of pose estimation, motion tracking, action recognition and motion generation.
 While these studies have significantly contributed to HPE, there remains a conspicuous absence of in-depth comparison and analysis between HPE and APE techniques.
-Reference (Jiang et al., 2022) discusses vision-based APE between 2013 and 2021, yet it falls short of offering an exhaustive discussion on the existing challenges and future directions in APE.
-Another limitation is that only a limited number of reviews (Zheng et al., 2020) discuss HPE or APE from sources other than vision-based inputs.
+Reference [52] discusses vision-based APE between 2013 and 2021, yet it falls short of offering an exhaustive discussion on the existing challenges and future directions in APE.
+Another limitation is that only a limited number of reviews [171] discuss HPE or APE from sources other than vision-based inputs.
 This paper aims to bridge these gaps in the literature by offering a detailed analysis that extends beyond the scope of prior surveys and reviews:
-  1. (1)
-    We meticulously reviewed 178 papers up to 2024, categorizing them based on different parameters such as their output forms (e.g.
+  • We meticulously reviewed 176 papers up to 2024, categorizing them based on different parameters such as their output forms (e.g.
 2D vs.
@@
 mesh pose representation), input sensor and modality types (monocular inputs vs.
-other sensory inputs; single-modal APE vs.
+other sensory inputs; uni-modal APE vs.
 multi-modal APE), learning paradigms (e.g.
@@
 self-supervised vs.
-unsupervised APE methods), experimental setup (e.g.
+unsupervised APE methods), and experimental setup (e.g.
 single-camera-view vs.
@@
 in-the-wild setup; the range of the sensors).
-  2. (2)
   • We summarised and discussed datasets and evaluation metrics for both 2D and 3D APEs not only based on RGB cameras but also include other sensors and modalities such as motion capture (MoCap) systems, RGBD, LiDAR, infrared, 3D Scanner, IMU, acoustic and language cues.
-  3. (3)
   • We delved into the potential and limitations of current multi-sensor and multi-modal APE, highlighting how these approaches can complement or enhance unimodal APE methodologies.
-  4. (4)
   • We presented an extensive comparison between HPE and APE.
@@
 It also explained the trend towards self-supervised and unsupervised methods for APE and compared the advantages and disadvantages of the existing supervised methods.
-  5. (5)
   • We also provided an overview of a wide array of applications of APE, ranging from wildlife conservation, biomechanics and neuroscience to 3D animal asset articulation, illustrating the practical implications and the interdisciplinary nature of APE.
-  6. (6)
   • Finally, we thoroughly discussed the key challenges in APE and pointed towards potential future research avenues to enhance performance and applicability.
-Table 2.
-RGB-based APE Dataset (Images-based).
+2 Datasets
+Table 2: RGB-based APE Dataset (Images-based).
 Column N indicates no.
-of animals: ”M” denotes multiple animals’ pose is annotated in one frame, while ”S” denotes a single animal’s pose is annotated.
-Column I/W is ”indoor” [Uncaptioned image] or ”in-the-wild” [Uncaptioned image] set-up.
-Sl Year Datasets & Size Animal Classes Source/Experiment N I/W Annotation
-set-up
+of animals: M denotes multiple animals’ pose is annotated in one frame, while S denotes a single animal’s pose is annotated.
+In I/W column, [Uncaptioned image] and [Uncaptioned image] represent indoor and in-the-wild setups, respectively.
+Exp.
+is Experiment.
+The dataset citation is available in column Year.
+Sl Year Datasets & Size Animal Classes Source & Exp. Set-up N I/W Annotation
 RGB-based APE Dataset - Images-based
-1 2011 Stanford Dogs (Khosla et al., 2011) Dogs (120 breeds) a. ImageNet data N/A [Uncaptioned image] Class labels,
-20k images b. Single view bounding box
-2 2014 COCO (Lin et al., 2014b) Cat, Dog, Horse, Bear, a. Online Repositories N/A [Uncaptioned image] & [Uncaptioned image] Bounding box, class
-1k images for animals Sheep, Zebra, Giraffe b. Single view label, segmentation
-3 2018 Animals-10 (C., 2018) Dog, Cat, Horse, a. Google Images N/A [Uncaptioned image] & [Uncaptioned image] N/A
-28k images Sheep, Cow, Elephant b. Single view
-4 2019 Animal Pose (Cao et al., 2019) Dogs, Cats, Cows, a. Animals-10 (C., 2018) & S [Uncaptioned image] & [Uncaptioned image] 20 keypoints (2D pose),
-4k images Horses, Sheep, COCO (Lin et al., 2014b) bounding box
-& other 7 categories b. Single view
-5 2019 Grevy’s Zebra (Zuffi et al., 2019b) Synthetic Zebra a. Computer generated S [Uncaptioned image] 28 keypoints (2D pose),
+1 2011 Stanford Dogs Dogs (120 breeds) a. ImageNet data N/A [Uncaptioned image] Class label, bounding box
+20k images b. Single view
+2 2014 COCO Bear, Sheep, Zebra, a. Online repositories N/A [Uncaptioned image] & [Uncaptioned image] Bounding box, class label,
+1k animals images Cat, Dog, Horse, b. Single view segmentation
+Giraffe
+3 2018 Animals-10 Dog, Horse, Sheep, a. Google images N/A [Uncaptioned image] & [Uncaptioned image] N/A
+28k images Cat, Cow, Elephant b. Single view
+4 2019 Animal Pose Dogs, Cats, Cows, a. Animals-10 (2018) & S [Uncaptioned image] & [Uncaptioned image] 20 keypoints (2D pose),
+4k images Horses, Sheep & COCO (2014) bounding box
+other 7 categories b. Single view
+5 2019 Grevy’s Zebra Synthetic Zebra a. Computer generated S [Uncaptioned image] 28 keypoints (2D pose),
 12,850 images b. Single view 3D mesh
-6 2020 Synthetic Animal (Mu et al., 2020) Synthetic Hound, a. COCO2017 (Lin et al., 2014b) S [Uncaptioned image] & [Uncaptioned image] 18 keypoints (2D &
-10k images Horse, Sheep, b. Single view 3D pose), depth maps,
-Elephant, Tiger segmentation
-7 2020 StanfordExtra (Biggs et al., 2020) Dogs a. Stanford Dogs (Khosla et al., 2011) S [Uncaptioned image] 24 keypoints (2D pose),
+6 2020 Synthetic Animal Synthetic Hound, a. COCO2017 (2014) S [Uncaptioned image] & [Uncaptioned image] 18 keypoints (2D&3D pose),
+10k images Elephant, Tiger, b. Single view depth maps, segmentation
+Horse, Sheep
+7 2021b AP-10K 23 animal families, a. Online repositories M [Uncaptioned image] 17 keypoints (2D pose)
+10,015 images 54 species b. Single view
+8 2020 StanfordExtra Dogs a. Stanford Dogs (2011) S [Uncaptioned image] 24 keypoints (2D pose),
 12k images b. Single view segmentation
-8 2021 AP-10K (Yu et al., 2021b) 23 animal families, a. Online Repositories [Uncaptioned image] 17 keypoints (2D pose)
-10,015 images 54 species b. Single view M
-9 2021 MacaquePose(Labuguen et al., 2021) Macaque Monkey a. Manually collected M [Uncaptioned image] 17 keypoint (2D pose)
-13,000 images & Onlince
+9 2021 MacaquePose Macaque Monkey a. Manually collected & M [Uncaptioned image] 17 keypoint (2D pose)
+13,000 images Online
 b. Single view
-10 2022 Animal Kingdom (Ng et al., 2022b) 850 animal species a. Online Repositories S [Uncaptioned image] 23 keypoints (2D pose)
+10 2022 Animal Kingdom 850 animal species a. Online repositories S [Uncaptioned image] 23 keypoints (2D pose)
 35k images b. Single view
-11 2022 OpenMonkey 26 Monkey species a. Manually collected S [Uncaptioned image] 17 keypoints (2D pose)
-Challenge (Yuan Yao and Park, 2022) & Online
+11 2022 OpenMonkey 26 Monkey species a. Manually collected & S [Uncaptioned image] 17 keypoints (2D pose)
+Challenge Online
 111k images b. Single view
-12 2023 OpenApePose (Desai et al., 2023) 6 Ape species a. Manually collected S [Uncaptioned image] 16 keypoints (2D pose)
+12 2023 OpenApePose 6 Ape species a. Manually collected S [Uncaptioned image] 16 keypoints (2D pose)
 71,868 images b. Single view
-13 2023 Animal3D (Xu et al., 2024) 40 Mammal species a. PartImageNet (He et al., 2022) S [Uncaptioned image] 26 keypoints (2D pose),
-3379 images &   COCO (Lin et al., 2014b) 3D mesh
+13 2023 Animal3D 40 Mammal species a. PartImageNet (2022) & S [Uncaptioned image] 26 keypoints (2D pose),
+3379 images COCO (2014) 3D mesh
 b. Single view
-2. Datasets
 It is hard to obtain accurate ground truth poses for APE, especially for wild animals, due to the unpredictable natural environments, the wide range of animal behaviours, occlusions, and the lack of controlled setups for precise annotation.
@@
 (2) Other-sensor-based datasets, which utilize sensors and modalities from sources beyond RGB, such as MoCap systems, 3D scanners, infrared cameras, IMUs, and LiDAR to acquire pose information.
-2.1. RGB-based Animal Pose Estimation Dataset
+2.1 RGB-based Animal Pose Estimation Dataset
 In RGB-based APE datasets, we can further divide them into two categories based on the data type: 1) image-based APE datasets, which rely on individual RGB images to provide pose annotations for animals (Table 2); 2) video-based APE datasets, which utilize video sequences, consisting of multiple frames, to capture animal pose labels over time (Table 3).
-2.1.1. Images-based Animal Pose Estimation Dataset
+2.1.1 Images-based Animal Pose Estimation Dataset
 Early datasets in this category lacked large-scale annotated animal pose labels.
-For example, the Stanford Dogs (2011) dataset (Khosla et al., 2011) includes 20k images of 120 dog breeds with class labels and bounding boxes but no pose keypoints.
-Similarly, the COCO (2014) dataset (Lin et al., 2014b) with 1k images only provides annotations such as bounding boxes, segmentation, and class labels for animals like cats, dogs, and horses.
-The Animals-10 (2018) dataset (C., 2018), consisting of 14k images of common animals such as dogs, cats, and horses, also lacks animal keypoint annotations.
+For example, the Stanford Dogs (2011) dataset [58] includes 20k images of 120 dog breeds with class labels and bounding boxes but no pose keypoints.
+Similarly, the COCO (2014) dataset [75] with 1k images only provides annotations such as bounding boxes, segmentation, and class labels for animals like cats, dogs, and horses.
+The Animals-10 (2018) dataset [1], consisting of 14k images of common animals such as dogs, cats, and horses, also lacks animal keypoint annotations.
 As there are few pose-labeled animal datasets available, more 2D pose-labeled animal datasets have been developed to evaluate the performance of APE objectively.
-For instance, the Animal Pose (2019) (Cao et al., 2019) consists of 4k images across 10 animal categories, with each image annotated with 20 2D keypoints to represent the animal skeleton.
-The StanfordExtra (2020) (Biggs et al., 2020) offers 12k dogs images with 24 2D pose keypoints.
-The AP-10K (2021) (Yu et al., 2021b) provides 10k images from 23 animal families, with 17 2D keypoints labeled for each detected animal.
-The recent Animal Kingdom (2022) (Ng et al., 2022b), the OpenMonkeyChallenge (2022) (Yuan Yao and Park, 2022), and the OpenApePose (2023) (Desai et al., 2023) all include annotated 2D animal poses keypoints.
+For instance, the Animal Pose (2019) [15] consists of 4k images across 10 animal categories, with each image annotated with 20 2D keypoints to represent the animal skeleton.
+The StanfordExtra (2020) [10] offers 12k dogs images with 24 2D pose keypoints.
+The AP-10K (2021) [164] provides 10k images from 23 animal families, with 17 2D keypoints labeled for each detected animal.
+The recent Animal Kingdom (2022) [103], the OpenMonkeyChallenge (2022) [158], and the OpenApePose (2023) [25] all include annotated 2D animal poses keypoints.
 Beyond 2D keypoint pose labels, more recent datasets have started providing 3D pose annotations, including 3D keypoints and meshes.
-Examples include the Synthetic Grevy’s Zebra (2019) (Zuffi et al., 2019b), the Synthetic Animal (2020) (Mu et al., 2020), and the Animal3D (2023) (Xu et al., 2024), which offer both high-quality 2D keypoint annotations and 3D pose labels.
+Examples include the Synthetic Grevy’s Zebra (2019) [176], the Synthetic Animal (2020) [95], and the Animal3D (2023) [150], which offer both high-quality 2D keypoint annotations and 3D pose labels.
 This development reflects the growing diversity and depth of datasets available for APE research.
-2.1.2. Video-based Animal Pose Estimation Dataset
-Table 3.
-RGB-based APE Dataset (Video-based).
+2.1.2 Video-based Animal Pose Estimation Dataset
+Table 3: RGB-based APE Dataset (Video-based).
+NGD denotes National Geographic Documentaries; the symbol [Uncaptioned image] denotes cameras.
 Column [Uncaptioned image] Rng is camera range.
-NGD denotes National Geographic Documentaries; the symbol [Uncaptioned image] denotes cameras.
+Exp.
+is Experiment.
+The dataset citation is available in column Year.
 All the other symbols and column meanings are defined in Table 2.
-Sl Year Datasets & Size Animal Classes Source & Experiment set-up [Uncaptioned image] Rng N I/W Annotation
+Sl Year Datasets & Size Animal Classes Source & Exp. Set-up [Uncaptioned image] Rng N I/W Annotation
 RGB-based APE Dataset - Videos-based
-14 2016 TigDog (Del Pero et al., 2017, 2015) Dogs, Horses, a. YouTube & NGD videos N/A S [Uncaptioned image] 19 keypoints
-110k frames Tigers b. Single view (2D pose)
-55 videos
-15 2018 BADJA (Biggs et al., 2018) Bears, Camels, a. DAVIS (Pont-Tuset et al., 2017) videos 1-20m S [Uncaptioned image] 20 keypoints
-11 videos Dogs, Horses, b. Single view (2D & 3D pose),
-Tigers, Cats, Cows segmentation
-16 2020 ATRW (Li et al., 2020b) Amur tigers a. Manually collected N/A M [Uncaptioned image] 15 keypoints
+14 2015 TigDog Dogs, Horses, Tigers a. YouTube & NGD videos N/A S [Uncaptioned image] 19 keypoints
+110k frames, 55 videos b. Single view (2D pose)
+15 2018 BADJA Bears, Camels, Dogs, a. DAVIS (2017) videos 1-20m S [Uncaptioned image] 20 keypoints
+11 videos Horses, Tigers, Cats, b. Single view (2D&3D pose),
+Cows segmentation
+16 2020 ATRW Amur tigers a. Manually collected N/A M [Uncaptioned image] 15 keypoints
 8k videos b. Single view (2D pose)
 17 2020 OpenMonkeyPose Monkeys a. Manually collected 0-2.5m M [Uncaptioned image] 13 keypoints
-(Bala et al., 2020) 195,228 frames b. Multi views (62 [Uncaptioned image] views) (3D pose)
-18 2021 AcinoSet (Joska et al., 2021) Cheetahs a. Manually collected 3-12m S [Uncaptioned image] 20 keypoints
-119k frames b. Multi views (2D & 3D pose)
-93 videos (6 GoPro [Uncaptioned image] viewpoint)
-19 2021 Horse-30 (Mathis et al., 2021) Horses a. Manually collected N/A S [Uncaptioned image] 22 keypoints
+195,228 frames b. Multi views (62 [Uncaptioned image] views) (3D pose)
+18 2021 AcinoSet Cheetahs a. Manually collected 3-12m S [Uncaptioned image] 20 keypoints
+119k frames, 93 videos b. Multi views (6 [Uncaptioned image] views) (2D&3D pose)
+19 2021 Horse-30 Horses a. Manually collected N/A S [Uncaptioned image] 22 keypoints
 8114 frames b. Single view (2D pose)
-20 2022 APT-36K (Yang et al., 2022b) Quadruped a. YouTube videos N/A M [Uncaptioned image] 17 keypoints
-36k frames Mammals b. Single view (2D pose)
-2.4k videos (30 Species)
-21 2023 ChimpACT (Ma et al., 2023) Chimpanzees a. Manually collected N/A M [Uncaptioned image] & [Uncaptioned image] 16 keypoints
-16,028 frames b. Single view (2D pose)
-163 videos
+20 2022 APT-36K Quadruped Mammals a. YouTube videos N/A M [Uncaptioned image] 17 keypoints
+36k frames, 2.4k videos (30 Species) b. Single view (2D pose)
+21 2023 ChimpACT Chimpanzees a. Manually collected N/A M [Uncaptioned image] & [Uncaptioned image] 16 keypoints
+16,028 frames, 163 videos b. Single view (2D pose)
 We have listed several well-known datasets in this category, with most pose annotations being presented in 2D and 3D keypoint formats (Table 3).
 In particular, most 2D keypoint pose labels in these datasets (e.g.
-Horse-30 (Mathis et al., 2021), APT-36K (Yang et al., 2022b), ChimpACT (Ma et al., 2023)) are provided by well-trained annotators.
-For instance, the TigDog (2016) (Del Pero et al., 2017, 2015) includes 110k frames of in-the-wild dogs, horses, and tigers, with 19 keypoint annotations for 2D poses.
-The Horse-30 (2021) (Mathis et al., 2021) provides 8k frames of horses in-the-wild, with 22 annotated 2D keypoints.
+Horse-30 [91], APT-36K [157], ChimpACT [87]) are provided by well-trained annotators.
+For instance, the TigDog (2016) [23, 22] includes 110k frames of in-the-wild dogs, horses, and tigers, with 19 keypoint annotations for 2D poses.
+The Horse-30 (2021) [91] provides 8k frames of horses in-the-wild, with 22 annotated 2D keypoints.
 While these datasets are valuable, they are limited to individual animals.
 However, there is a recent growing trend toward providing multi-animal pose annotations (e.g.
-ATRW (Li et al., 2020b), APT-36K (Yang et al., 2022b) and ChimpACT (Ma et al., 2023)) to account for the frequent interactions between quadruped animals.
+ATRW [73], APT-36K [157] and ChimpACT [87]) to account for the frequent interactions between quadruped animals.
 There is also a small collection of publicly available 3D animal pose benchmark datasets.
-For example, the BADJA (2018) (Biggs et al., 2018) contains 11 in-the-wild animal videos annotated with 20 3D joints mapped onto a SMAL mesh.
-The AcinoSet (2021) (Joska et al., 2021) offers 7k frames of hand-labelled 3D keypoints from cheetahs, capturing their running behaviour in-the-wild.
-Additionally, the OpenMonkeyPose (2020) (Bala et al., 2020) includes 195k well-annotated 3D keypoint pose instances, spanning a large variation of poses and positions of monkeys across diverse activities.
-It is worth noting that both the OpenMonkeyPose (Bala et al., 2020) and AcinoSet (Joska et al., 2021) include synchronized multiple-camera-views footage and camera calibration, enabling multiple-camera-views geometry and triangulation for precise 3D pose estimation (e.g.
+For example, the BADJA (2018) [11] contains 11 in-the-wild animal videos annotated with 20 3D joints mapped onto a SMAL mesh.
+The AcinoSet (2021) [53] offers 7k frames of hand-labelled 3D keypoints from cheetahs, capturing their running behaviour in-the-wild.
+Additionally, the OpenMonkeyPose (2020) [8] includes 195k well-annotated 3D keypoint pose instances, spanning a large variation of poses and positions of monkeys across diverse activities.
+It is worth noting that both the OpenMonkeyPose [8] and AcinoSet [53] include synchronized multiple-camera-views footage and camera calibration, enabling multiple-camera-views geometry and triangulation for precise 3D pose estimation (e.g.
 OpenMonkeyPose offers 62-viewpoint, while AcinoSet uses 6).
@@
 Due to the significant variations in appearance, behaviour, and joint distributions between species, APE models trained on these datasets often fail to generalize to unseen species, leading to poor performance in real-world scenarios.
-To fill this gap, the APT-36K (2022) (Yang et al., 2022b) was introduced, covering 2k video clips of 30 quadruped species with high-quality 2D keypoint annotations.
+To fill this gap, the APT-36K (2022) [157] was introduced, covering 2k video clips of 30 quadruped species with high-quality 2D keypoint annotations.
 This enables inter-species generalization testing and reflects the need for models that work across diverse species.
-Table 4. Other-Sensor-based APE Dataset. [Uncaptioned image] is markers; vol. is volume. All the other symbols and column meanings are defined in table 2 and table 3.
+Table 4: Other-Sensor-based APE Dataset.
+[Uncaptioned image] is markers; vol.
+is volume.
+The dataset citation is available in column Year.
+All the other symbols and column meanings are defined in Table 2 and Table 3.
 Sl Year Datasets & Size Animal Source/Experiment [Uncaptioned image] Rng N I/W Annotation
@@
 Other-Sensor-based APE Dataset - RGB+MoCap-based
-22 2021 PAIR-R24M (Marshall et al., 2021) Rat a. Manually collected N/A M [Uncaptioned image] 12 keypoints
+22 2021 PAIR-R24M Rat a. Manually collected N/A M [Uncaptioned image] 12 keypoints
 24.3M frames b. Multi views (6 RGB video [Uncaptioned image]) (3D pose)
 c. 1 MoCap [Uncaptioned image] & 12 [Uncaptioned image]
-23 2021 Rat7M (Dunn et al., 2021) Rat a. Manually collected N/A S [Uncaptioned image] 20 keypoints
+23 2021 Rat7M Rat a. Manually collected N/A S [Uncaptioned image] 20 keypoints
 7M frames b. Multi views (6 RGB video [Uncaptioned image]) (3D pose)
 c. 12 MoCap [Uncaptioned image] & 20 [Uncaptioned image]
-24 2023 3D-POP (Naik et al., 2023) Pigeons a. Manually collected 0-4.2m (RGB) M [Uncaptioned image] 9 keypoints
-300k frames b. Multi views (4 RGB [Uncaptioned image]) 15x7x4m vol. (2D & 3D pose)
-4 videos c. 30 Mocap [Uncaptioned image] & 8 [Uncaptioned image] (Mocap)
-25 2024 PFERD (Li et al., 2024c) Horses a. Manually collected 16x20m S [Uncaptioned image] 3D Mesh,
+24 2023 3D-POP Pigeons a. Manually collected M [Uncaptioned image] 9 keypoints
+300k frames, b. Multi views (4 RGB [Uncaptioned image]) 0-4.2m (RGB) (2D & 3D pose)
+4 videos c. 30 MoCap [Uncaptioned image] & 8 [Uncaptioned image] 15x7x4m (MoCap)
+25 2024a PFERD Horses a. Manually collected 16x20m S [Uncaptioned image] 3D Mesh,
 55+ hrs video b. Multi views (10 RGB [Uncaptioned image]) from the 2D joints
@@
 Other-Sensor-based APE Dataset - RGB+Depth+MoCap-based
-26 2020 RGBD-Dog (Kearney et al., 2020a) Dogs a. Manually collected 0.5-4.5m S [Uncaptioned image] 3D Mesh,
+26 2020 RGBD-Dog Dogs a. Manually collected 0.5-4.5m S [Uncaptioned image] 3D Mesh,
 No size given b. Multi views (6 RGBD sensors) (Depth) 3D Skeletal data,
@@
 Other-Sensor-based APE Dataset - RGB+LiDAR-based
-27 2024 WildPose (Muramatsu et al., 2024) Lion, a. Manually collected 2–120m M [Uncaptioned image] 15 keypoints,
+27 2024 WildPose Lion, a. Manually collected 2–120m M [Uncaptioned image] 15 keypoints,
 No size given Springbok b. Single view 3D point cloud
-c. 1 RGB Camera + LiDAR
 Other-Sensor-based APE Dataset - Infrared-based
-28 2023 LoTE-Animal (Liu et al., 2023) 11 animal a. Manually collected & Online N/A S [Uncaptioned image] 17 keypoints
+28 2023 LoTE-Animal 11 animal a. Manually collected & Online N/A S [Uncaptioned image] 17 keypoints
 33k images species b. Single view (2D pose)
 Other-Sensor-based APE Dataset - 3D scanner-based
-29 2024 VAREN (Zuffi et al., 2024) Horses a. Manually collected N/A S [Uncaptioned image] 3D Mesh
+29 2024 VAREN Horses a. Manually collected N/A S [Uncaptioned image] 3D Mesh
 4000 3D scans b. No. of scanner Not given
 Other-Sensor-based APE Dataset - MoCap+IMU-based
-30 2024 Yigit et al. (Yigit et al., 2022) Horse a. Manually collected S [Uncaptioned image] & [Uncaptioned image] 3D Motion data
+30 2022 Yigit et al. Horse a. Manually collected S [Uncaptioned image] & [Uncaptioned image] 3D Motion data
 No size given b. Multi views
@@
 d. 4 IMUs 0-50m (IMU)
-2.2. Other-Sensor-based Animal Pose Estimation Dataset
+2.2 Other-Sensor-based Animal Pose Estimation Dataset
 With developments in sensor technology, more APE datasets start to incorporate data from sources beyond traditional RGB cameras, such as depth sensors, MoCap systems, 3D scanners, LiDARs, and IMUs, offering increased accuracy and versatility (Table 4).
@@
 An alternative approach for achieving high accuracy and low-noise 3D ground-truth animal pose data is to use a maker-based MoCap system with RGB cameras.
-For example, the Rat7M (2021) (Dunn et al., 2021) and PFERD (2024) (Li et al., 2024c) utilize MoCap system with RGB cameras and a set of makers to generate 3D ground truth pose data for rats and horses, respectively.
+For example, the Rat7M (2021) [30] and PFERD (2024) [65] utilize MoCap system with RGB cameras and a set of makers to generate 3D ground truth pose data for rats and horses, respectively.
 Although these are effective for tracking individual animals from multiple views with temporal consistency, these datasets are limited by their focus on single animals and dependence on a large number of markers.
-To address these limitations, recent datasets like PAIR-R24M (2021) (Marshall et al., 2021) and 3D-POP (2023) (Naik et al., 2023) have been developed.
+To address these limitations, recent datasets like PAIR-R24M (2021) [89] and 3D-POP (2023) [99] have been developed.
 These datasets use MoCap systems with fewer markers (12 and 8 markers, respectively) to track the 3D postures of multiple animals in the same frame, offering more scalable solutions for multi-animal tracking.
-Similarly, a recent dataset (Kearney et al., 2020a) utilizes MoCap systems with 63–82 markers to generate highly accurate 3D ground truth pose labels for dogs.
+Similarly, a recent dataset [56] utilizes MoCap systems with 63–82 markers to generate highly accurate 3D ground truth pose labels for dogs.
 This dataset also integrates depth sensors (i.e.
@@
 Although the MoCap-based dataset provides high-quality 3D ground truth pose labels, APE algorithms trained using these MoCap datasets may exhibit poor generalization to in-the-wild scenarios, as the data collection for the MoCap-based datasets is all done indoors in a controlled environment, as you can see in Table 4.
-WildPose (Muramatsu et al., 2024) is another promising multi-sensor system that combine RGB and LiDAR data.
+WildPose [97] is another promising multi-sensor system that combine RGB and LiDAR data.
 This system captures 2D RGB video and 3D point cloud data from a distance (up to 120 meters), making it particularly suited for studying free-ranging animals in natural habitats.
 In addition to these technologies, infrared cameras for camera-trapping in the wild, which use infrared LEDs to illuminate the scene to enable nighttime observations, have also been employed for wild animal pose labelling.
-For instance, the LoTE-Animal (2023) (Liu et al., 2023) uses infrared trap cameras to collect large-scale video footage of endangered animals across 11 species.
+For instance, the LoTE-Animal (2023) [77] uses infrared trap cameras to collect large-scale video footage of endangered animals across 11 species.
 This dataset captures a wide range of variations—such as ecological seasons, weather conditions, viewpoints, and habitats—and provides both instance segmentation labels and 2D pose keypoint annotations.
 The LoTE-Animal dataset is especially valuable for advancing semi-supervised and self-supervised APE models, with potential applications in conservation efforts for endangered species.
-Moreover, the VAREN (2024) dataset (Zuffi et al., 2024) focuses on 3D mesh representation, featuring 4k 3D scans of horses provided by the 3D scanner; this further advances 3D modelling and representation in APE research.
-Lastly, the IMU-based animal pose dataset collected by (Yigit et al., 2022) consists of acceleration and gyroscope measurements from sensors attached to the horse’s limbs.
+Moreover, the VAREN (2024) dataset [174] focuses on 3D mesh representation, featuring 4k 3D scans of horses provided by the 3D scanner; this further advances 3D modelling and representation in APE research.
+Lastly, the IMU-based animal pose dataset collected by [162] consists of acceleration and gyroscope measurements from sensors attached to the horse’s limbs.
 The data captures each limb’s 3D motion characteristics (like linear acceleration and angular velocity).
@@
 These emerging sensor-based datasets offer new opportunities for APE, enhancing the precision and breadth of APE research across diverse animal species and environments.
-3. Error Metrics and Evaluation Methods
+3 Error Metrics and Evaluation Methods
 This section provides insight into the error (err.) metrics and evaluation (eval.) process of the APE methods based on different sensors and modalities.
-3.1. RGB-based 2D Animal Pose Estimation Metric
+3.1 RGB-based 2D Animal Pose Estimation Metric
 This section discusses the common evaluation metrics used for RGB-based APE methods that extract poses in the 2D keypoint format, and more evaluation metrics are given in Table 5 and Table 6.
-2D single-animal pose estimation methods like DeepLabCut (Mathis et al., 2018) and LEAP (Pereira et al., 2019) use a common metric like Root Means Square Error (RMSE) to evaluate their performance.
+2D single-animal pose estimation methods like DeepLabCut [90] and LEAP [111] use a common metric like Root Means Square Error (RMSE) to evaluate their performance.
 RMSE is versatile and can be used for both 2D and 3D pose estimation and for evaluating various types of keypoints, regardless of the specific application domain.
@@
 It is a normalised measure that indicates how well the keypoints are localised, and a higher PCK value indicates better localisation performance.
-Single-animal pose estimation methods like Synthetic-to-Real (Li and Lee, 2021b), Kim et al.
-(Kim et al., 2022), ScarceNet (Li and Lee, 2023), D-Gen (Li et al., 2023) and FSKD (Lu and Koniusz, 2022) use this metric for evaluation.
+Single-animal pose estimation methods like Synthetic-to-Real [69], Kim et al.
+[60], ScarceNet [70], D-Gen [71] and FSKD [83] use this metric for evaluation.
 Moreover, Li et al.
-(Li et al., 2020a) evaluate their single-animal APE methods with PCK, RMSE and Area Under Curve (AUC) along with an extension to these metrics with Permutation Invariant (PI).
+[72] evaluate their single-animal APE methods with PCK, RMSE and Area Under Curve (AUC) along with an extension to these metrics with Permutation Invariant (PI).
 Li et al.
-(Li et al., 2020a) introduce PI metrics to evaluate the accuracy of pose estimation in challenging scenarios where it is difficult to uniquely identify limb identities (e.g.
+[72] introduce PI metrics to evaluate the accuracy of pose estimation in challenging scenarios where it is difficult to uniquely identify limb identities (e.g.
 in fruit flies) with traditional metrics.
 PI metrics are designed to account for the ambiguity in identifying corresponding limbs or joints between the predicted pose and ground truth.
-Object Keypoint Similarity (OKS) has also been used in evaluating 2D single-animal pose estimation methods like D-Gen (Li et al., 2023); 2D multiple-animal pose estimation methods like DepthFormer (Liu et al., 2022b), Straka et al.
-(Straka et al., 2024), SLEAP (Pereira et al., 2022), and also 2D video based semi-supervised method like SemiMultiPose (Blau et al., 2022).
+Object Keypoint Similarity (OKS) has also been used in evaluating 2D single-animal pose estimation methods like D-Gen [71]; 2D multiple-animal pose estimation methods like DepthFormer [78], Straka et al.
+[134], SLEAP [112], and also 2D video based semi-supervised method like SemiMultiPose [12].
 OKS quantifies how well the keypoints are located to spatial scale and keypoint visibility, with closeness defined based on ground truth.
@@
 The OKS scores for every keypoint are computed for all objects in the dataset, and the average OKS is used to evaluate a model.
-Another common metric used for 2D APE is mean Average Precision (mAP), which has been used in evaluating ScarceNet (Li and Lee, 2023), SIPEC (Marks et al., 2022), SuperAnimal-Quadruped (Shaokai Ye and Mathis, 2024), OpenPose (Cao et al., 2017), DLCRNet (Lauer et al., 2022), maDLC (Lauer et al., 2022), SLEAP (Pereira et al., 2022), SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SemiMultiPose (Blau et al., 2022).
+Another common metric used for 2D APE is mean Average Precision (mAP), which has been used in evaluating ScarceNet [70], SIPEC [88], SuperAnimal-Quadruped [160], OpenPose [16], DLCRNet [64], maDLC [64], SLEAP [112], SuperAnimal-TopViewMouse [160] and SemiMultiPose [12].
 The mAP is an evaluation metric based on OKS and has become widely used for pose estimation, particularly following its adoption in the COCO keypoint challenge.
@@
 Last but not least, another metric that has been used for the 2D video-based APE method evaluation is Canonical Correlation Analysis (CCA), which is a statistical technique used to examine the relationship between two sets of variables.
-It determines linear combinations of the variables in each set that are maximally correlated; this metric has been used for evaluating the Lightning Pose (Biderman et al., 2023) method, however, we have not seen wide adoption of this evaluation metric within the APE community.
-3.2. RGB-based 3D Animal Pose Estimation Metric
+It determines linear combinations of the variables in each set that are maximally correlated; this metric has been used for evaluating the Lightning Pose [9] method, however, we have not seen wide adoption of this evaluation metric within the APE community.
+3.2 RGB-based 3D Animal Pose Estimation Metric
 This section discusses the common evaluation metrics used for both keypoint-based and mesh-based 3D RGB models within the APE community, more evaluation metrics for the 3D APE methods are given in Table 7.
 3D APE metrics can also incorporate some of the metrics used in 2D APE.
-For example, PCK is employed in methods such as Three-D Safari (Zuffi et al., 2019a), SMBLD (Biggs et al., 2020), Li et al.
-(Li and Lee, 2021a), Birds of a Feather (Wang et al., 2021a), BITE (Rüegg et al., 2023), MagicPony (Wu et al., 2023b), DigiDogs (Shooter et al., 2024), Monet (Yao et al., 2019), DAN-NCE (Dunn et al., 2021), and Muramatsu et al.
-(Muramatsu et al., 2022).
+For example, PCK is employed in methods such as Three-D Safari [176], SMBLD [10], Li et al.
+[68], Birds of a Feather [144], BITE [120], MagicPony [148], DigiDogs [128], Monet [159], DAN-NCE [30], and Muramatsu et al.
+[96].
 The well-known RMSE metric is also used in 3D APE methods such as Hu et al.
-(Hu et al., 2023), Muramatsu et al.
-(Joska et al., 2021), 3D-UPPER (Ebrahimi et al., 2023), and SBeA (Han et al., 2024).
-Apart from PCK and RMSE, 3D pose estimation-specific metrics such as Mean Per Joint Position Error (MPJPE) have been used in 3D APE methods like DigiDogs (Shooter et al., 2024) and MAMMAL (An et al., 2023).
+[47], Muramatsu et al.
+[53], 3D-UPPER [31], and SBeA [43].
+Apart from PCK and RMSE, 3D pose estimation-specific metrics such as Mean Per Joint Position Error (MPJPE) have been used in 3D APE methods like DigiDogs [128] and MAMMAL [2].
 MPJPE measures the average Euclidean distance between the predicted and ground truth joint positions across all joints in the skeleton, quantifying the accuracy of the joint position estimation.
@@
 An extension of the MPJPE metric is N-MPJPE, which has been used by Dai et al.
-(Dai et al., 2023) for their 3D APE model.
+[19] for their 3D APE model.
 N-MPJPE involves normalising the ground truth 3D pose by subtracting the mean and dividing by the standard deviation, aligning the predicted 3D pose with the normalised ground truth using Procrustes alignment, and then calculating the MPJPE, which results in Normalized-MPJPE.
-For mesh-based APE as discussed in section 4.3.1, a well-known evaluation metric is the 3D Chamfer Distance, which has been used in several model-free methods such as DOVE (Wu et al., 2023a), MagicPony (Wu et al., 2023b), and BANMo (Yang et al., 2022a).
+For mesh-based APE, a well-known evaluation metric is the 3D Chamfer Distance, which has been used in several model-free methods such as DOVE [147], MagicPony [148], and BANMo [153].
 The 3D Chamfer Distance measures the average distance between the reconstructed surface points and the ground truth surface points.
@@
 A lower value indicates better reconstruction quality.
-Another approach in mesh-based APE is the model-based pose estimation methods, for example, SMAL (Zuffi et al., 2017).
+Another approach in mesh-based APE is the model-based pose estimation methods, for example, SMAL [178].
 In these approaches, metrics like Scan-to-Mesh distance, keypoint reprojection error, and silhouette reprojection error are commonly used.
@@
 Similarly, the silhouette reprojection error measures how accurately the 3D model silhouette matches the ground truth silhouette in the 2D image, indicating how well the 3D shape projection aligns with the visible boundaries of the animal in the image.
-Other model-based methods, such as SMBLD (Biggs et al., 2020) and Li et al.
-(Li and Lee, 2021a), use metrics like PCK and Intersection over Union (IoU) for performance evaluation.
+Other model-based methods, such as SMBLD [10] and Li et al.
+[68], use metrics like PCK and Intersection over Union (IoU) for performance evaluation.
 Moreover, due to the scarcity of animal datasets that include 3D pose annotations, comprehensive quantitative evaluations of 3D pose predictions are not always feasible.
 Therefore, researchers often assess 3D pose predictions using qualitative means, as Sosa et al.
-(Sosa and Hogg, 2023) explained.
+[132] explained.
 Examples of qualitative evaluations can be found in Salem et al.
-(Salem et al., 2019), Sosa et al.
-(Sosa and Hogg, 2023), Monet (Yao et al., 2019), and Muramatsu et al.
-(Joska et al., 2021).
-3.3. Other-Modality-based Animal Pose Estimation Metric
+[123], Sosa et al.
+[132], Monet [159], and Muramatsu et al.
+[53].
+3.3 Other-Modality-based Animal Pose Estimation Metric
 Different sensor modalities have different ways of evaluating their APE models.
-For example, an IMU-based APE method (Yigit et al., 2022) is evaluated by calculating the mean and standard deviation (Std) of the estimated joint angle errors and comparing the predictions with ground truth data obtained through a MoCap system.
+For example, an IMU-based APE method [162] is evaluated by calculating the mean and standard deviation (Std) of the estimated joint angle errors and comparing the predictions with ground truth data obtained through a MoCap system.
 Additionally, the model is assessed for Intra-individual Variability (IAV) and Inter-individual Variability (IEV), which quantify the consistency of limb movement across different gait cycles and between different horses.
@@
 This thorough approach ensures that this IMU-based model’s performance is evaluated for accuracy and consistency across different conditions.
-On the other hand, one of the RGBD-based APE methods, RGBD-Dog (Kearney et al., 2020a), is evaluated using MPJPE and PCK.
+On the other hand, one of the RGBD-based APE methods, RGBD-Dog [56], is evaluated using MPJPE and PCK.
 The evaluation is conducted across different body parts and for entire dog poses.
@@
 In the context of pose estimation, it aligns the predicted 3D skeleton with the ground truth skeleton by minimising the sum of squared distances between corresponding points in the two sets.
-In the RGBD-Dog (Kearney et al., 2020a), PA is applied to refine the existing PCK and MPJPE evaluation metrics to PA-PCK, PA-PCK 3D and PA-MPJPE, which is then used to measure the error after aligning the predicted and ground truth skeletons to focus purely on pose accuracy.
-Other multi-modal systems, for example, Sound of Motion (RGB+acoustic) (Li et al., 2022) and CLHOP (audio+video) (Li et al., 2024a), also use PA-MPJPE for their APE evaluation in addition to PCK and IoU, which are common in the RGB-based APE evaluations.
+In the RGBD-Dog [56], PA is applied to refine the existing PCK and MPJPE evaluation metrics to PA-PCK, PA-PCK 3D and PA-MPJPE, which is then used to measure the error after aligning the predicted and ground truth skeletons to focus purely on pose accuracy.
+Other multi-modal systems, for example, Sound of Motion (RGB+acoustic) [67] and CLHOP (audio+video) [66], also use PA-MPJPE for their APE evaluation in addition to PCK and IoU, which are common in the RGB-based APE evaluations.
 This section briefly summarises some of the metrics for non-RGB-based APE models, more evaluation metrics are provided in Table 8.
-4. RGB-based Animal Pose Estimation
+4 RGB-based Animal Pose Estimation
 In recent years, APE has achieved significant progress.
-Among the various modalities utilised for APE, RGB-based approaches have emerged as the most developed and extensively studied research method (Pereira et al., 2019; Cao et al., 2019; Li and Lee, 2021b; Lauer et al., 2022).
-These approaches leverage visual information captured by RGB cameras, take animal images or videos as inputs, and subsequently produce the animal’s pose, which manifests as a 2D skeletal keypoints representation of pose, 3D skeletal keypoints representation of pose, or 3D animal mesh reconstructions, as exemplified in LABEL:fig:Skeleton_Pose_Representations and LABEL:fig:Body_Mesh_Reconstructions.
-In addition to skeletal keypoints and mesh representation for pose estimations, dense pose representations have also been used mainly for humans (Riza Alp Güler, 2018) and chimpanzees (Sanakoyeu et al., 2020).
+Among the various modalities utilised for APE, RGB-based approaches have emerged as the most developed and extensively studied research method [111, 15, 69, 64].
+These approaches leverage visual information captured by RGB cameras, take animal images or videos as inputs, and subsequently produce the animal’s pose, which manifests as a 2D skeletal keypoints representation of pose, 3D skeletal keypoints representation of pose, or 3D animal mesh reconstructions, as exemplified in Figure 1(a) and Figure 1(b).
+In addition to skeletal keypoints and mesh representation for pose estimations, dense pose representations have also been used mainly for humans [118] and chimpanzees [124].
 In this paper, skeletal keypoints and mesh representations APE have been mainly analysed in detail, while keypoint-based methods remain dominant due to their established presence, easier access to annotated data, and lower computational requirements.
@@
 Traditionally, we use direct visual observation to extract the animal poses, which can offer real-time and context-rich data.
-However, it suffers from inherent drawbacks, including observer bias, mental fatigue (Graving et al., 2019), and long data collection and observation times.
-Later, researchers start to use video monitoring and recording to expedite data collection and minimise the influence of human involvement (Graving et al., 2019).
+However, it suffers from inherent drawbacks, including observer bias, mental fatigue [40], and long data collection and observation times.
+Later, researchers start to use video monitoring and recording to expedite data collection and minimise the influence of human involvement [40].
 However, manual scoring of videos by eyes still remains time-consuming and often fails to capture subtle and fast-paced movements that are integral to various behaviours.
@@
 Marker-based methods rely on attaching markers (e.g.
-MoCap suits (Kearney et al., 2020b), animal-borne cameras (Patel et al., 2017)) to animals for posture measurements.
-While these methods can achieve precise pose estimation, the reliance on physical markers can pose practical challenges, especially when investigating innate behaviours in natural habitats (Wu et al., 2023b) or when observing the movements of small creatures like mice (Lauer et al., 2022) and fish.
+MoCap suits [56], animal-borne cameras [108]) to animals for posture measurements.
+While these methods can achieve precise pose estimation, the reliance on physical markers can pose practical challenges, especially when investigating innate behaviours in natural habitats [148] or when observing the movements of small creatures like mice [64] and fish.
 Furthermore, most marker-based approaches are limited to highly controlled lab environments and specialised hardware setups.
 Such methodologies are often impractical when the target animals demonstrate varied behaviours within unfamiliar environments.
-With the rapid advancement in Computer Vision (CV), many automated markerless APE methods (Karashchuk et al., 2021; Mathis et al., 2018; Lauer et al., 2022; Pereira et al., 2019; Graving et al., 2019) start to employ supervised learning to directly infer and estimate animals’ pose and movements from images or videos, removing the need for physical markers that may alter natural behaviour.
-For example, the markerless method DeepLabCut (Mathis et al., 2018) has been used for behaviour tracking in wild chimpanzees and bonobos (Wiltshire et al., 2023).
+With the rapid advancement in Computer Vision (CV), many automated markerless APE methods [55, 90, 64, 111, 40] start to employ supervised learning to directly infer and estimate animals’ pose and movements from images or videos, removing the need for physical markers that may alter natural behaviour.
+For example, the markerless method DeepLabCut [90] has been used for behaviour tracking in wild chimpanzees and bonobos [146].
 However, a significant obstacle in APE is the acquisition of accurate ground truth poses as discussed in Section 2.
-To address the scarcity of large-scale, well-labelled animal pose datasets, many weakly and semi-supervised (Cao et al., 2019), semi-supervised (Kim et al., 2022; Blau et al., 2022; Biderman et al., 2023; Yao et al., 2019), self-supervised (Sosa and Hogg, 2023; Shooter et al., 2024; Yang et al., 2022a; Han et al., 2024) and unsupervised learning (Dai et al., 2023; Kim et al., 2022; Li and Lee, 2021b; Ebrahimi et al., 2023; Shaokai Ye and Mathis, 2024; Wu et al., 2023a, b; Li et al., 2024b) approaches have been developed.
-For instance, recent literature (Dai et al., 2023) proposes an unsupervised 3D animal canonical pose estimation method with geometric self-supervision; this method necessitates only 2D poses for training, which significantly alleviates the requirement for extensively labelled data.
+To address the scarcity of large-scale, well-labelled animal pose datasets, many weakly and semi-supervised [15], semi-supervised [60, 12, 9, 159], self-supervised [132, 128, 153, 43] and unsupervised learning [19, 60, 69, 147, 31, 148, 160, 74] approaches have been developed.
+For instance, recent literature [19] proposes an unsupervised 3D animal canonical pose estimation method with geometric self-supervision; this method necessitates only 2D poses for training, which significantly alleviates the requirement for extensively labelled data.
 Current APE research trends show growing interest in unsupervised, semi-supervised, and self-supervised methods due to the mentioned challenges to reach the same performance level as supervised methods.
-However, supervised methods are still a significant focus in current APE studies, especially with the emergence of larger datasets and benchmarking platforms like AP-10K (Yu et al., 2021b), APT-36K (Yang et al., 2022b), TigDog (Del Pero et al., 2015, 2017), and AcinoSet (Joska et al., 2021), as we can see in Table 5-Table 7.
+However, supervised methods are still a significant focus in current APE studies, especially with the emergence of larger datasets and benchmarking platforms like AP-10K [164], APT-36K [157], TigDog [22, 23], and AcinoSet [53], as we can see in Table 5-Table 7.
 This review paper categorises the existing uni-modal RGB-based APE studies into two main groups: 1) 2D APE methods from images (Table 5, Section 4.1) and from videos (Table 5 & Table 6, Section 4.2), 2) 3D APE methods from images and videos (Table 7, Section 4.3).
-4.1. 2D Animal Pose Estimation from an Image
+4.1 2D Animal Pose Estimation from an Image
 This section focuses on 2D APE methods that do not use temporal information to localize the 2D coordinates of an animal’s or animal’s joints from a single image frame (Table 5).
@@
 multiple frames) can also extract poses from a single image.
-In Table 5, column ”I,V” provides the information on APE methods that can be applied on both an image and videos; or only on an image.
+In Table 5, column “I,V” provides the information on APE methods that can be applied on both an image and videos; or only on an image.
 Based on the number of target animals being processed in the scene, this pose estimation problem can be further divided into single-animal pose estimation (Section 4.1.1) and multiple-animal pose estimation (Section 4.1.2).
-While pose estimation in individual animals has received significant attention and serves as the foundation for most APE studies, extending it to multiple-animal pose estimation is an active area of research that is still evolving (Pereira et al., 2022).
-4.1.1. Single Animal
+While pose estimation in individual animals has received significant attention and serves as the foundation for most APE studies, extending it to multiple-animal pose estimation is an active area of research that is still evolving [112].
+4.1.1 Single Animal
 2D single-frame single-animal pose estimation aims to accurately estimate the 2D pose of an animal from a single image.
-These methodologies share a conceptual foundation with HPE (Zheng et al., 2020) and can be broadly classified into two categories: heatmap-based thods.
+These methodologies share a conceptual foundation with HPE [171] and can be broadly classified into two categories: heatmap-based and regression methods.
 The former predicts the locations of an animal’s body joints using heatmaps for supervision, necessitating a subsequent postprocessing step to identify the final keypoints based on the predicted heatmaps.
 In contrast, the latter approach directly regresses the keypoint coordinates of the animal’s body joints, employing an end-to-end deep neural network to establish a mapping from the input image to the body keypoints.
-LABEL:fig:single_animal illustrates the general architecture of 2D single-frame single-animal pose estimation methods.
-Heatmap-based frameworks such as (Pereira et al., 2019; Kim et al., 2022; Li and Lee, 2023; Li et al., 2020a; Li and Lee, 2021b; Graving et al., 2019; Li et al., 2023; Mathis et al., 2018) have become popular in 2D APE.
+Figure 3 illustrates the general architecture of 2D single-frame single-animal pose estimation methods.
+(a) Heatmap-based Method [122].
+(b) Regression Method using Few-Shot Keypoint Detection [83].
+Figure 3: The general architectures of 2D single-frame single-animal pose estimation methods.
+Heatmap-based frameworks such as [111, 60, 70, 72, 69, 40, 71, 90] have become popular in 2D APE.
 These approaches are structured around creating a 2D heatmap, interpreted as a probabilistic representation of body part locations within the image.
-The architecture of a typical heatmap-based scheme includes i) a CNN backbone to generate heatmaps from input images, followed by ii) a post-processing step to finalize the pose estimations based on these predictions (LABEL:heatheat).
-There are some common architectures to input an image and output the corresponding heatmaps in step i); for instance, (Kim et al., 2022; Li and Lee, 2021b; Li et al., 2023; Mathis et al., 2018) uses the ResNet (He et al., 2016) and its variants to produce the heatmap; (Li and Lee, 2023) adapts a variant of the HRNet (Sun et al., 2019) as the backbone; (Li et al., 2020a) applies the stacked hourglass (Newell et al., 2016) architectures.
-Subsequently, in step ii), the predicted heatmaps are then processed to produce precise 2D spatial coordinates for each body joint using techniques such as peak detection (Graving et al., 2019; Mathis et al., 2018; Pereira et al., 2019).
+The architecture of a typical heatmap-based scheme includes i) a CNN backbone to generate heatmaps from input images, followed by ii) a post-processing step to finalize the pose estimations based on these predictions (Figure 3(a)).
+There are some common architectures to input an image and output the corresponding heatmaps in step i); for instance, [60, 69, 71, 90] uses the ResNet [46] and its variants to produce the heatmap; [70] adapts a variant of the HRNet [135] as the backbone; [72] applies the stacked hourglass [102] architectures.
+Subsequently, in step ii), the predicted heatmaps are then processed to produce precise 2D spatial coordinates for each body joint using techniques such as peak detection [40, 90, 111].
 Although heatmap-based approaches are well known for their robustness against occlusions and variations in appearance, their computational complexity has been critiqued for impeding end-to-end training.
-Nonetheless, DeepPoseKit (Graving et al., 2019) proposes an efficient end-to-end heatmap-based approach.
+Nonetheless, DeepPoseKit [40] proposes an efficient end-to-end heatmap-based approach.
 Regression-based approaches directly regress the coordinates of joint locations as continuous values, employing CNNs to learn a mapping from the input image to the joint coordinates.
 The output layer of the network consists of multiple units, each representing the $x$ and $y$ coordinates of a joint, and is trained to minimize the distance between the predicted and ground truth coordinates.
-Although regression-based approaches are less common than heatmap-based methods, they have shown great promise (Lu and Koniusz, 2022).
-One notable work (Lu and Koniusz, 2022) adopts a regression-based approach to estimate keypoint positions directly, bypassing the intermediate heatmap representation (LABEL:RegressionRegression).
+Although regression-based approaches are less common than heatmap-based methods, they have shown great promise [83].
+One notable work [83] adopts a regression-based approach to estimate keypoint positions directly, bypassing the intermediate heatmap representation (Figure 3(b)).
 Their Few-shot Keypoint Detection (FSKD) model is designed to detect keypoints on unseen species with minimal annotated data, which is particularly useful in challenging scenarios where the species have not been previously encountered or the available annotated data is limited.
@@
 However, they may face long-range 2D offset regression challenges, which may lead to a notable performance gap compared to heatmap-based methods.
-Table 5.
-RGB-based 2D APE Methods.
-In the column ”I,V”: ”I,V” is the methods that can perform pose estimation on both an image and videos, whereas, ”I” is the APE methods that do only on an image.
-Column ”L” indicates the learning types as follows: ”S” is supervised methods, ”U” is unsupervised methods, ”SF” is self-supervised methods, ”SM” is semi-supervised methods.
+Table 5: RGB-based 2D APE Methods.
+In the column I,V: I,V is the methods that can perform pose estimation on both an image and videos, whereas, I is the APE methods that do only on an image.
+Column L indicates the learning types as follows: S is supervised methods, U is unsupervised methods, SF is self-supervised methods, SM is semi-supervised methods.
 [Uncaptioned image] is CNN architecture, [Uncaptioned image] is the use of the attention mechanism, [Uncaptioned image] is GAN, [Uncaptioned image] is underlying architecture as vision transformer, [Uncaptioned image] is Multi-Scale Domain Adaptation Module (MDAM).
@@
 [Uncaptioned image] is single-camera-view, [Uncaptioned image] is multi-camera-views.
-[Uncaptioned image] is ”indoor” experimental set-up and [Uncaptioned image] is ”in-the-wild” experimental set-up.
+[Uncaptioned image] is indoor experimental set-up and [Uncaptioned image] is in-the-wild experimental set-up.
 Symbols inside the {} indicate that both categories are used in that method.
+The method citation is available in column Year.
 Sl Year Methods I,V L Features Evaluated On Metrics
@@
 Heatmap-based:
-1 2018 DeepLabCut (Mathis et al., 2018) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mammals RMSE (pixels)
-2 2019 LEAP (Pereira et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse RMSE (pixels)
-3 2019 DeepPoseKit (Graving et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Zebra (Zuffi et al., 2019b) ([Uncaptioned image]), Euclidean Error,
-Desert locust, Fly ([Uncaptioned image]) Bayesian Inference (Graving et al., 2019)
-4 2020 Li et al. (Li et al., 2020a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Worm, RMSE, PCK, AUC (pixels),
-Zebrafish PI-AUC (4-45 pixels),
-PI-PCK (5&15 pixels),
-PI-RMSE (pixels)
-5 2021 Synthetic-to-Real (Li and Lee, 2021b) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse, TigDog(Del Pero et al., 2017) PCK@0.05 (%)
-6 2022 Kim et al. (Kim et al., 2022) I SM,U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] {[Uncaptioned image], [Uncaptioned image]} TigDog(Del Pero et al., 2017), PCK@0.05 (%)
-Synthetic Animal (Mu et al., 2020),
-Animal Pose (Cao et al., 2019)
-7 2023 ScarceNet (Li and Lee, 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), mAP (%), PCK@0.05 (%)
-TigDog(Del Pero et al., 2017)
-8 2023 D-Gen (Li et al., 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), OKS, PCK@0.05 (%)
-Animal Pose (Cao et al., 2019),
-Animal Kingdom (Ng et al., 2022a)
+1 2018 DeepLabCut I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mammals RMSE (pixels)
+2 2019 LEAP I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse RMSE (pixels)
+3 2019 DeepPoseKit I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Zebra (2019) ([Uncaptioned image]); Euclidean Error,
+Desert locust, Fly ([Uncaptioned image]) Bayesian Inference 2019
+4 2020 Li et al. I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Worm, Zebrafish RMSE, PI-RMSE, PCK, AUC,
+PI-PCK, PI-AUC
+5 2021b Synthetic-to-Real I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse, TigDog (2017) PCK@0.05 (%)
+6 2022 Kim et al. I SM,U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] {[Uncaptioned image], [Uncaptioned image]} TigDog (2017), PCK@0.05 (%)
+Animal Pose (2019),
+Synthetic Animal (2020)
+7 2023 ScarceNet I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] TigDog (2017), mAP (%), PCK@0.05 (%)
+AP-10K (2021b)
+8 2023 D-Gen I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), OKS, PCK@0.05 (%)
+AP-10K (2021b),
+Animal Kingdom (2022)
 Regression-based:
-9 2022 FSKD (Lu and Koniusz, 2022) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animal Pose (Cao et al., 2019) PCK
+9 2022 FSKD I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animal Pose (2019) PCK
 2D Multi-Animal Pose Estimation Methods
 Top-down:
-10 2022 SIPEC (Marks et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Primates, Mice Pearson Correlation,
-RMSE (pixels), mAP (%),
-IoU, Dice Coefficient
-11 2022 DepthFormer (Liu et al., 2022b) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b) OKS
-12 2024 Straka et al. (Straka et al., 2024) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Lynx PCK, OKS, AP
-13 2024 SuperAnimal-Quadruped I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), RMSE (pixels), mAP,
-(Shaokai Ye and Mathis, 2024) Animal Pose (Cao et al., 2019) Normalized Error
+10 2022 SIPEC I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Primates, Mice Pearson Correlation, RMSE (pixels),
+mAP (%), IoU, Dice Coefficient
+11 2022 DepthFormer I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (2021b) OKS
+12 2024 Straka et al. I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Lynx PCK, OKS, AP
+13 2024 SuperAnimal- I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), RMSE (pixels), Normalized Error,
+Quadruped AP-10K (2021b) mAP
 Bottom-up:
-14 2017 OpenPose (Cao et al., 2017) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} COCO (Lin et al., 2014b), MPII (Andriluka et al., 2014) mAP (%)
-15 2021 Negrete et al. (Blanco Negrete et al., 2021) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] MacaquePose (Labuguen et al., 2021), AP
-MSCOCO (Lin et al., 2014a)
-16 2021 Farahnakian et al. (Farahnakian et al., 2021) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pig(Nath et al., 2019) Error (pixels), RMSE (pixels)
-17 2022 DLCRNet, maDLC (Lauer et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} 3 Mice ([Uncaptioned image]), 1 Mice ([Uncaptioned image]), RMSE (pixels),
-2 marmosets ([Uncaptioned image]) mAP, PCK
-Bottom-up & Top-Down:
-18 2022 SLEAP (Pereira et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, mAP (%), PCK,
-Mice, Gerbils OKS, Localization
-19 2024 SuperAnimal-TopViewMouse I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), RMSE (pixels), mAP,
-(Shaokai Ye and Mathis, 2024) Animal Pose (Cao et al., 2019) Normalized Error
-4.1.2. Multiple Animals
+14 2017 OpenPose I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} COCO (2014), MPII (2014) mAP (%)
+15 2021 Negrete et al. I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] MSCOCO (2014), AP
+MacaquePose (2021)
+16 2021 Farahnakian et al. I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pig (2019) Error (pixels), RMSE (pixels)
+17 2022 DLCRNet, maDLC I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} 3 Mice, 2 Marmosets ([Uncaptioned image]); RMSE (pixels), mAP, PCK
+1 Mice ([Uncaptioned image])
+Bottom-up &
+Top-down:
+18 2022 SLEAP I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, Mice, Gerbils mAP (%), PCK, OKS, Localization
+19 2024 SuperAnimal- I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), RMSE (pixels), Normalized Error,
+TopViewMouse AP-10K (2021b) mAP
+4.1.2 Multiple Animals
 We have seen significant developments in markless pose estimation in individual animals.
-However, extending these methodologies to multiple animals poses unique challenges, particularly in studying social behaviours or animals in their natural environments (Yang et al., 2022b; Gong et al., 2022).
+However, extending these methodologies to multiple animals poses unique challenges, particularly in studying social behaviours or animals in their natural environments [157, 37].
 For example, it is common for the body parts of one animal to be obscured by another in a multi-animal pose estimation.
@@
 However, given these challenges, 2D multiple-animal pose estimation research has still garnered significant attention due to its potential applications in various domains, such as animal behaviour analysis, wildlife monitoring, and ecological studies.
-Methodologies developed in this field are categorized into two main categories: top-down methods (Marks et al., 2022; Liu et al., 2022b; Straka et al., 2024; Shaokai Ye and Mathis, 2024) and bottom-up methods (Cao et al., 2017; Blanco Negrete et al., 2021; Farahnakian et al., 2021; Lauer et al., 2022) (LABEL:fig:multi_animal).
-Top-down approaches for multiple-animal pose estimation (LABEL:aaa) generally involve a two-step process: first i) detecting each animal in the input image using pre-trained object detectors and then ii) applying 2D single-animal pose estimation techniques to estimate the pose for each detected instance.
-Common object detectors in step i) include Faster R-CNN (Ren et al., 2015), the Single Shot MultiBox Detector (SSD) (Liu et al., 2016), You Only Look Once (YOLO) v4 (Gong et al., 2022), and RetinaNet (Lin et al., 2017); common 2D APE methods in step ii) includes variants of ResNet, HRNets and transformer.
-In the early work, SIPEC (Marks et al., 2022) uses SIPEC:SegNet (a Mask R-CNN architecture (He et al., 2017)) to segment animal instances in the images or videos first, then it uses SIPEC:PoseNet to estimate the pose for each individual.
+Methodologies developed in this field are categorized into two main categories: top-down methods [88, 78, 134, 160] and bottom-up methods [16, 101, 34, 64] (Figure 4).
+(a) Top-Down Approach: SLEAP Method [112] Instances - Parts.
+(b) Bottom-Up Approach: SLEAP Method [112] Parts - Instances.
+Figure 4: 2D single-frame multiple-animal pose estimation using SLEAP method [112].
+Top-down approaches for multiple-animal pose estimation (Figure 4(a)) generally involve a two-step process: first i) detecting each animal in the input image using pre-trained object detectors and then ii) applying 2D single-animal pose estimation techniques to estimate the pose for each detected instance.
+Common object detectors in step i) include Faster R-CNN [117], the Single Shot MultiBox Detector (SSD) [79], You Only Look Once (YOLO) v4 [37], and RetinaNet [76]; common 2D APE methods in step ii) includes variants of ResNet, HRNets and transformer.
+In the early work, SIPEC [88] uses SIPEC:SegNet (a Mask R-CNN architecture [45]) to segment animal instances in the images or videos first, then it uses SIPEC:PoseNet to estimate the pose for each individual.
 Later, we can spot a trend of using HRNets and transformers as the 2D APE backbone architecture in step ii) to achieve better performance, especially with the limited animal pose data.
-For example, the top-down DepthFormer method (Liu et al., 2022b) has been proposed to achieve lightweight multiple-animal pose estimation by designing a new basic block based on the transformer structure and the similarities between self-attention and depthwise convolution.
-In 2024, SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) adapts Faster R-CNN for enhanced feature extraction to detect animal instances in the given image or videos first, and then also uses the transformer-based AnimalTokenPose models to achieve high zero-shot APE performance in OOD animal datasets.
-In the same paper, both SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SuperAnimal-Quadruped (Shaokai Ye and Mathis, 2024) models also enable us to use HRNet-w32 (Wang et al., 2020) as the underlying architecture to get the final animal 2D pose.
-Until now, SuperAnimal (Shaokai Ye and Mathis, 2024) has claimed to be the current state-of-the-art across a wide variety of animal species for both 2D images and videos.
+For example, the top-down DepthFormer method [78] has been proposed to achieve lightweight multiple-animal pose estimation by designing a new basic block based on the transformer structure and the similarities between self-attention and depthwise convolution.
+In 2024, SuperAnimal-TopViewMouse [160] adapts Faster R-CNN for enhanced feature extraction to detect animal instances in the given image or videos first, and then also uses the transformer-based AnimalTokenPose models to achieve high zero-shot APE performance in OOD animal datasets.
+In the same paper, both SuperAnimal-TopViewMouse [160] and SuperAnimal-Quadruped [160] models also enable us to use HRNet-w32 [142] as the underlying architecture to get the final animal 2D pose.
+Until now, SuperAnimal [160] has claimed to be the current state-of-the-art across a wide variety of animal species for both 2D images and videos.
 More information on how SuperAnimal is adopted to videos is discussed in Section 4.2.
 Another recent top-down approach for 2D endangered species (i.e.
-lynx) pose estimation with limited data (Straka et al., 2024) also leverages the high-resolution network HRNet-w32 (Wang et al., 2020) architecture in step ii) to get accurate 2D animal pose, as this HRNet-w32 architecture helps to maintain detailed spatial information throughout the network.
+lynx) pose estimation with limited data [134] also leverages the high-resolution network HRNet-w32 [142] architecture in step ii) to get accurate 2D animal pose, as this HRNet-w32 architecture helps to maintain detailed spatial information throughout the network.
 All of these methodologies prove the trend.
-In contrast, bottom-up methods (LABEL:bbb) prioritize predicting all keypoint locations in an image first, then grouping them into individual instances.
-Initial work in this area includes methods like OpenPose (Cao et al., 2017), which introduced the concept of Part Affinity Fields (PAFs) to associate body parts with individuals coherently.
-This PAFs idea was further adapted in the DLCRNet and maDLC architecture (Lauer et al., 2022) to predict the location and orientation of animal limbs and link the specific keypoints to the correct individual.
-OpenPose (Cao et al., 2017) was also used in (Blanco Negrete et al., 2021) to associate detected keypoints to particular individuals by taking advantage of similarities in skeleton structure between humans and monkeys.
-Further refinement in the bottom-up approach was made by works such as (Farahnakian et al., 2021), which focused on pig body part association using DeepLabCut (Mathis et al., 2018).
+In contrast, bottom-up methods (Figure 4(b)) prioritize predicting all keypoint locations in an image first, then grouping them into individual instances.
+Initial work in this area includes methods like OpenPose [16], which introduced the concept of Part Affinity Fields (PAFs) to associate body parts with individuals coherently.
+This PAFs idea was further adapted in the DLCRNet and maDLC architecture [64] to predict the location and orientation of animal limbs and link the specific keypoints to the correct individual.
+OpenPose [16] was also used in [101] to associate detected keypoints to particular individuals by taking advantage of similarities in skeleton structure between humans and monkeys.
+Further refinement in the bottom-up approach was made by works such as [34], which focused on pig body part association using DeepLabCut [90].
 Both top-down and bottom-up methods have limitations and features.
@@
 On the other hand, the primary shortcoming of the bottom-up approaches often stems from incorrect assembly.
-To allow the users to compare the two competing approaches directly, researchers have developed both bottom-up and top-down variants for maDLC (Lauer et al., 2022), SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SLEAP (Pereira et al., 2022).
+To allow the users to compare the two competing approaches directly, researchers have developed both bottom-up and top-down variants for maDLC [64], SuperAnimal-TopViewMouse [160] and SLEAP [112].
 The decision between these two methods depends on the specific requirements and constraints of the application.
-Before we move on to the next section, it is worth noting that some of the APE methods explained in section 4.1 can perform pose estimation on both images and videos, as can be seen in table 5 in column ”IV”.
+Before we move on to the next section, it is worth noting that some of the APE methods explained in Section 4.1 can perform pose estimation on both images and videos, as can be seen in Table 5 in column IV.
 In Section 4.2, we explain which of these video-based 2D APE methods extract pose frame by frame, and which uses the spatiotemporal signals of video data in an end-to-end manner for pose estimation.
-4.2. 2D Animal Pose Estimation from Videos
+4.2 2D Animal Pose Estimation from Videos
 2D multi-frame APE transforms consecutive image sequences or videos into probabilistic estimates of the 2D spatial coordinates of keypoint locations for target animals across frames.
 Based on how these approaches leverage the temporal dimension in video data to extract poses, we categorize 2D APE from videos into two main categories: (1) post-processing methods to incorporate temporal information; and (2) end-to-end methods to incorporate temporal information, as shown in Table 6.
-Table 6.
-RGB-based 2D APE methods from videos.
+Table 6: RGB-based 2D APE methods from videos.
 All the symbols and column meanings are defined in Table 5, and any additional symbols are defined here: [Uncaptioned image] is the underlying architecture as 3D-CNN, [Uncaptioned image]+[Uncaptioned image] is the use of CNN-based pose module followed by an Optical Flow module, [Uncaptioned image] is Bayesian Ensembling, [Uncaptioned image] is Probabilistic Graphical Model.
+The method citation is available in column Year.
 Sl Year Methods L Features Evaluated On Metrics
 2D APE from Videos: End-to-End Methods to Incorporate Temporal Information
-1 2021 OptiFlex (Liu et al., 2021) S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image] Mouse, Fish, Monkey, Fly aPCK error (%)
-2 2022 T-LEAP (Russello et al., 2022) S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cow PCKh@0.2(%)
-3 2022 SLEAP (Pereira et al., 2022) S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, Mice, Gerbils mAP (%), PCK,
-OKS, Localization
-4 2022 SemiMultiPose (Blau et al., 2022) SM [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mouse, HoneyBee, Fly mAP, OKS
-5 2024 Lightning Pose (Biderman et al., 2023) SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse, Fish CCA, Pixel Error, AUC
-6 2024 SuperAnimal (Shaokai Ye and Mathis, 2024) U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b) RMSE (pixels), mAP
-Animal Pose (Cao et al., 2019) Normalized Error
-4.2.1. Post-Processing Methods to Incorporate Temporal Information
+1 2021 OptiFlex S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image] Mouse, Fish, Monkey, Fly aPCK error (%)
+2 2022 T-LEAP S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cow PCKh@0.2(%)
+3 2022 SLEAP S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, Mice, Gerbils mAP (%), PCK, OKS, Localization
+4 2022 SemiMultiPose SM [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mouse, HoneyBee, Fly mAP, OKS
+5 2023 Lightning Pose SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse, Fish CCA, Pixel Error, AUC
+6 2024 SuperAnimal U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), RMSE (pixels), Normalized Error,
+AP-10K (2021b) mAP
+4.2.1 Post-Processing Methods to Incorporate Temporal Information
 Early APE frameworks process each video frame independently to predict pose keypoints, without inherently incorporating temporal features in an end-to-end manner during training.
 Temporal consistency is typically addressed through optional post-processing methods, such as smoothing or filtering, to improve performance.
-Some examples that follow this frame-by-frame approach include DeepLabCut (Mathis et al., 2018), DeepPoseKit (Graving et al., 2019), and LEAP (Pereira et al., 2019).
-4.2.2. End-to-End Methods to Incorporate Temporal Information
+Some examples that follow this frame-by-frame approach include DeepLabCut [90], DeepPoseKit [40], and LEAP [111].
+4.2.2 End-to-End Methods to Incorporate Temporal Information
 Recently, researchers have achieved more robust predictions by utilising the spatiotemporal information embedded within video data during pose estimation model training in an end-to-end manner.
@@
 We highlight several methods that fall in this category in the following bullet points and Table 6.
-A. Extensions of LEAP method
-A notable extension of the LEAP (Pereira et al., 2019) method is T-LEAP (Russello et al., 2022), which conceptualises temporal information as an auxiliary dimension within the CNN framework, thereby employing 3D convolutional layers to capture both spatial and temporal features concurrently to extract animal poses from video data.
+Extensions of LEAP Method: A notable extension of the LEAP [111] method is T-LEAP [122], which conceptualises temporal information as an auxiliary dimension within the CNN framework, thereby employing 3D convolutional layers to capture both spatial and temporal features concurrently to extract animal poses from video data.
 To leverage the temporal information embedded in videos, T-LEAP processes a sequence of frames of length $T$, and outputs the confidence maps linked to the last frame $t=T$.
@@
 For example, in a scenario where a cow’s leg is discernible in one frame but obscured in the subsequent frame, T-LEAP can employ visible information to predict the probable position of the leg in the occluded frame.
-Another extension of the LEAP model is SLEAP (Pereira et al., 2022), which also incorporates temporal information, but unlike LEAP, this extends the model to multi-animal pose estimation.
-B. Optical Flow-based Methods
-OptiFlex (Liu et al., 2021) is an optical flow-based methods that allow the estimation of motion vectors between consecutive frames to capture the temporal dynamics of animal poses.
-In particular, OptiFlex (Liu et al., 2021) combines a deep learning-based pose estimation module called FlexibleBaseline with an optical flow module named OpticalFlow to enhance its APE performance across multiple frames.
+Another extension of the LEAP model is SLEAP [112], which also incorporates temporal information, but unlike LEAP, this extends the model to multi-animal pose estimation.
+Optical Flow-based Methods: OptiFlex [81] is an optical flow-based methods that allow the estimation of motion vectors between consecutive frames to capture the temporal dynamics of animal poses.
+In particular, OptiFlex [81] combines a deep learning-based pose estimation module called FlexibleBaseline with an optical flow module named OpticalFlow to enhance its APE performance across multiple frames.
 The FlexibleBaseline module predicts the initial pose for each frame, and the OpticalFlow module converges initial predictions on a target frame and its adjacent frames into a final prediction for the target frame.
-By tracking the pixel-level displacements over time, optical flow estimation techniques, such as the Lucas-Kanade algorithms (Baker and Matthews, 2004), can compute dense motion fields between frames, thereby inferring the movement of body parts and joints.
+By tracking the pixel-level displacements over time, optical flow estimation techniques, such as the Lucas-Kanade algorithms [7], can compute dense motion fields between frames, thereby inferring the movement of body parts and joints.
 Suppose a certain body part is occluded or unclear in a frame.
 In that case, the model can infer its likely position based on the detected motion and poses in adjacent frames, showing its robustness against temporary obstruction of keypoint poses.
-C. Future Trend using Semi-supervised and Unsupervised Video Adaptation Learning
-Recently, there has been a growing shift towards using semi-supervised learning (e.g.
-SemiMultiPose (Blau et al., 2022), Lightning Pose (Biderman et al., 2023)) and unsupervised video adaptation learning (e.g.
-SuperAnimal (Shaokai Ye and Mathis, 2024)) to extract poses from videos.
+Future Trend using Semi-supervised and Unsupervised Video Adaptation Learning: Recently, there has been a growing shift towards using semi-supervised learning (e.g.
+SemiMultiPose [12], Lightning Pose [9]) and unsupervised video adaptation learning (e.g.
+SuperAnimal [160]) to extract poses from videos.
 In 2D single-frame APE methods, supervised learning dominates, relying on many labelled frames for training.
@@
 In real-world scenarios like wildlife videography, there are exponentially more abundant unlabeled frames than labelled frames.
-Given these deficiencies, SemiMultiPose (Blau et al., 2022) proposes a semi-supervised architecture to exploit the abundant structures pervasive in these unlabeled video frames to enhance the training process, which is particularly useful in the sparse labelling video-based scenarios.
-Similarly, Lightning Pose (2023) (Biderman et al., 2023) incorporates a semi-supervised approach to exploit the spatiotemporal statistics present in unlabeled videos.
+Given these deficiencies, SemiMultiPose [12] proposes a semi-supervised architecture to exploit the abundant structures pervasive in these unlabeled video frames to enhance the training process, which is particularly useful in the sparse labelling video-based scenarios.
+Similarly, Lightning Pose (2023) [9] incorporates a semi-supervised approach to exploit the spatiotemporal statistics present in unlabeled videos.
 This methodology introduces unsupervised training objectives that penalise the network when its predictions deviate from the expected smoothness of physical motion, multiple-camera-views geometry, or stray from a low-dimensional subspace of plausible body configurations.
@@
 This incorporation of context is instrumental in resolving transient occlusions and ambiguities between proximal and morphologically similar body parts.
-Last, the SuperAnimal (2024) (Shaokai Ye and Mathis, 2024), an unsupervised video adaptation APE method, can be applied to images and videos.
+Last, the SuperAnimal (2024) [160], an unsupervised video adaptation APE method, can be applied to images and videos.
 However, its approach differs depending on the data type.
@@
 These works demonstrate better APE performance with fewer labels, generalise better to unseen videos, and provide smoother and more reliable pose trajectories for downstream analysis, which proves the shift towards semi-supervised and unsupervised learning in APE.
-4.3. 3D Animal Pose Estimation from Images and Videos
+4.3 3D Animal Pose Estimation from Images and Videos
 3D APE refers to the task of determining the locations of animal body joints in 3D space.
@@
 Conversely, the integration of multiple-camera-views imaging facilitates a well-conditioned approach to interpolate the lost dimension for 3D APE using geometric consistency.
-Based on the number of camera views used to extract the animal poses, we categorize the 3D APE into two sections: (1) 3D APE from single-camera-view images and videos (Salem et al., 2019; Gosztolai et al., 2021; Dai et al., 2023; Sosa and Hogg, 2023; Shooter et al., 2024; Hu et al., 2023; Zuffi et al., 2017, 2019a; Biggs et al., 2020; Li and Lee, 2021a; Wang et al., 2021a; Rüegg et al., 2022, 2023; Wu et al., 2023a, b; Yang et al., 2022a; Li et al., 2024b), (2) 3D APE from multiple-camera-views images and videos (Yao et al., 2019; Bala et al., 2020; Zhang et al., 2021; Karashchuk et al., 2021; Dunn et al., 2021; Muramatsu et al., 2022; Ebrahimi et al., 2023; An et al., 2023; Han et al., 2024).
-Table 7.
-3D RGB-based APE methods.
-All the symbols and column meanings are defined in table 5 and table 6, and any additional symbols are defined here: [Uncaptioned image] is shape model (Zuffi et al., 2017), [Uncaptioned image] is Structured Forests architecture (Dollár and Zitnick, 2015), [Uncaptioned image] is Graph Convolutional Network (GCN), [Uncaptioned image] is Stacked Hourglass architecture based on CNN, [Uncaptioned image] is GAN, [Uncaptioned image] is the combination of Neural Radiance Fields (NeRF), Volumetric Signed Distance Functions (VolSDF) & Differentiable Physics Simulation, [Uncaptioned image] is a framework based on Epipolar Divergence, [Uncaptioned image] is Gaussian-based model, [Uncaptioned image] is Statistical Shape Model (SSM), [Uncaptioned image] is NeRF, [Uncaptioned image] is Semantic Bank of Skinned Models (SBSM), [Uncaptioned image] is feed-forward neural network.
-[Uncaptioned image] is the Body Mesh Recovery method
+Based on the number of camera views used to extract the animal poses, we categorize the 3D APE into two sections: (1) 3D APE from single-camera-view images and videos [123, 39, 19, 132, 128, 47, 178, 176, 10, 68, 144, 121, 147, 120, 148, 153, 74], (2) 3D APE from multiple-camera-views images and videos [159, 8, 168, 55, 30, 96, 31, 2, 43].
+Table 7: 3D RGB-based APE methods.
+All the symbols and column meanings are defined in Table 5 and Table 6, and any additional symbols are defined here: [Uncaptioned image] is shape model [178], [Uncaptioned image] is Structured Forests architecture [27], [Uncaptioned image] is Graph Convolutional Network (GCN), [Uncaptioned image] is Stacked Hourglass architecture based on CNN, [Uncaptioned image] is GAN, [Uncaptioned image] is the combination of Neural Radiance Fields (NeRF), Volumetric Signed Distance Functions (VolSDF) & Differentiable Physics Simulation, [Uncaptioned image] is a framework based on Epipolar Divergence, [Uncaptioned image] is Gaussian-based model, [Uncaptioned image] is Statistical Shape Model (SSM), [Uncaptioned image] is NeRF, [Uncaptioned image] is Semantic Bank of Skinned Models (SBSM), [Uncaptioned image] is feed-forward neural network.
+[Uncaptioned image] is the Body Mesh Recovery method.
+The method citation is available in column Year.
 Sl Year Methods VI L Features Evaluated On Evaluation Metrics
-3D APE from Image & Video - Single-Camera-View Methods:
+3D APE from Image & Video - Single-Camera-View Methods
 Skeleton-Only:
-1 2019 Salem et al. (Salem et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse Pose Distance Measure (Perona, 2010),
+1 2019 Salem et al. I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse Pose Distance Measure (2010),
 Reconstruction Err.
-2 2021 LiftPose3D (Gosztolai et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice, Rats & Macaques MAE
-3 2023 Hu et al. (Hu et al., 2023) V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse OKS for 2D Keypoints,
+2 2021 LiftPose3D V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice, Rats, Macaques MAE
+3 2023 Hu et al. V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse OKS for 2D Keypoints,
 RMSE for 3D Keypoints
-4 2023 Dai et al. (Dai et al., 2023) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AcinoSet (Joska et al., 2021), N-MPJPE
-SMAL (Zuffi et al., 2017)
-5 2023 Sosa et al. (Sosa and Hogg, 2023) I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse PCK@0.05 for 2D,
-(Weizmann Data) (Borenstein et al., 2004) Qualitative for 3D
-6 2024 DigiDogs/D-Pose (Shooter et al., 2024) I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dogs PCK for 2D&3D, MPJPE for 3D
-AMR (Model-Based):
-7 2017 SMAL (Zuffi et al., 2017) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Quadruped Scan-to-Mesh Distance,
-Keypoint Reprojection Err.,
-Silhouette Reprojection Err.
-8 2019 Three-D Safari (Zuffi et al., 2019a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Zebra PCK, IoU, Geodesic Distance
-9 2020 SMBLD (Biggs et al., 2020) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) IoU, PCK
-10 2021 Li et al.
-(Li and Lee, 2021a) I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020), PCK, IoU
-Animal Pose (Cao et al., 2019),
-BADJA (Biggs et al., 2018)
-11 2021 Birds of a Feather (Wang et al., 2021a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Birds, StanfordExtra (Biggs et al., 2020) PCK, IoU
-12 2022 BARC (Rüegg et al., 2022) I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) PCK & IoU for 2D,
+4 2023 Dai et al. I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] SMAL (2017), N-MPJPE
+AcinoSet (2021)
+5 2023 Sosa et al. I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse - Weizmann (2004) PCK@0.05 for 2D,
+Qualitative for 3D
+6 2024 DigiDogs/D-Pose I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dogs PCK for 2D & 3D,
+MPJPE for 3D
+AMR
+(Model-Based):
+7 2017 SMAL I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Quadruped Scan-to-Mesh Distance,
+Keypoint and Silhouette
+Reprojection Err.
+8 2019 Three-D Safari I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Zebra PCK, IoU, Geodesic Distance
+9 2020 SMBLD I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (2020) IoU, PCK
+10 2021a Li et al. I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} BADJA (2018), PCK, IoU
+Animal Pose (2019),
+StanfordExtra (2020)
+11 2021 Birds of a Feather I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Birds, PCK, IoU
+StanfordExtra (2020)
+12 2022 BARC I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (2020) PCK & IoU for 2D,
 Breed Prototype Consistency &
 Perceptual Shape Eval. for 3D
-13 2023 BITE (Rüegg et al., 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) Scan-to-Mesh Distance,
-Semi-Synthetic 3D (Rüegg et al., 2023) Mesh-to-Scan Distance,
+13 2023 BITE I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (2020), Scan-to-Mesh Distance,
+Semi-Synthetic 3D (2023) Mesh-to-Scan Distance,
 PCK, IoU
-AMR (Model-Free):
-14 2023 DOVE (Wu et al., 2023a) I,V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Bird, Horse, mIoU,
+AMR
+(Model-Free):
+14 2022 BANMo I,V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} , [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animals & Humans F-score,
+3D Chamfer Distance (cm)
+15 2023 DOVE I,V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Bird, Horse, mIoU,
 3D Toy Bird Data 3D Chamfer Distance (cm)
-15 2023 MagicPony (Wu et al., 2023b) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse-10 (Mathis et al., 2021), PCK,
-DOVE Data (Wu et al., 2023a), 3D Chamfer Distance (cm)
-COCO (Lin et al., 2014a), & others
-16 2022 BANMo (Yang et al., 2022a) I,V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} , [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animals & Humans 3D Chamfer Distance (cm),
-F-score
-17 2024 3D-Fauna (Li et al., 2024b) I U [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Quadruped Animals PCK@0.1, KT-PCK@0.1
-3D APE from Image & Video - Multi-Camera-Views Methods:
-18 2019 Monet (Yao et al., 2019) I SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dog, Monkey PCK, Reprojection Err.
-19 2020 OpenMonkeyStudio (Bala et al., 2020) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Monkey Median & Mean Err. (cm),
-Standard Deviation (cm)
-20 2021 GIMBAL (Zhang et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse MPE (mm), RPA-MPE (mm)
-21 2021 DAN-NCE (Dunn et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rats, Marmosets Mice PCK, Reconstruction Err.
+16 2023 MagicPony I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] COCO (2014), PCK,
+Horse-10 (2021), 3D Chamfer Distance (cm)
+DOVE (2023), others
+17 2024b 3D-Fauna I U [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Quadruped Animals PCK@0.1, KT-PCK@0.1
+3D APE from Image & Video - Multi-Camera-Views Methods
+18 2019 Monet I SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dog, Monkey PCK, Reprojection Err.
+19 2020 OpenMonkeyStudio I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Monkey Median & Mean Err. (cm),
+Std (cm)
+20 2021 GIMBAL V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse MPE (mm), RPA-MPE (mm)
+21 2021 DAN-NCE V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rats, Marmosets Mice, PCK, Reconstruction Err.
 Rat Pups & Chickadees
-22 2021 Anipose (Karashchuk et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse Joint Position Err. (mm)
-23 2022 Muramatsu et al. (Muramatsu et al., 2022) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AcinoSet (Joska et al., 2021) RMSE, PCK,
-$l^{2}$ norm in pixels
-24 2023 3D-UPPER (Ebrahimi et al., 2023) V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice RMSE, Scalar Product,
+22 2021 Anipose V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse Joint Position Err. (mm)
+23 2022 Muramatsu et al. V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AcinoSet (2021) RMSE, PCK, $l^{2}$ norm in pixels
+24 2023 3D-UPPER V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice RMSE, Scalar Product,
 Pearson’s Correlation
-25 2024 MAMMAL (An et al., 2023) V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pigs, Mouse, Dogs MPJPE, IoU,
+25 2023 MAMMAL V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pigs, Mouse, Dogs MPJPE, IoU,
 Avg. Err. on each keypoint (cm)
-26 2024 SBeA (Han et al., 2024) V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mice, Dogs, Birds RMSE (pixels), Reprojection Err.
-4.3.1. Single Camera View
+26 2024 SBeA V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mice, Dogs, Birds RMSE (pixels), Reprojection Err.
+4.3.1 Single Camera View
 Single-camera-view 3D APE methods reconstruct 3D animal poses from monocular images and videos, which presents a challenge due to the inherent ambiguity in deriving depth from 2D projections.
@@
 (2) Animal Mesh Recovery (AMR), which recovers the animal’s entire mesh.
-Skeleton-only Methods
 Skeleton-only methods are dedicated to estimating the 3D coordinates of animal joints for a skeletal representation as their result.
@@
 This approach typically leverages strong priors for predicting 3D joint locations from a single-camera view.
-A notable example is the LiftPose3D framework (Gosztolai et al., 2021), which trains a supervised neural network utilizing 3D ground truth data to lift 2D poses from one camera.
+A notable example is the LiftPose3D framework [39], which trains a supervised neural network utilizing 3D ground truth data to lift 2D poses from one camera.
 The training data is augmented with different camera angles and bone lengths, which allows the network to internalize camera parameters implicitly and accommodate variations in animal sizes.
-In contrast, recent innovations have endeavoured to estimate camera parameters and construct shape distributions explicitly (Hu et al., 2023).
-In this paper (Hu et al., 2023), 3D APE is formulated as an optimization problem utilizing a pre-defined mouse skeletal model.
+In contrast, recent innovations have endeavoured to estimate camera parameters and construct shape distributions explicitly [47].
+In this paper [47], 3D APE is formulated as an optimization problem utilizing a pre-defined mouse skeletal model.
 This 3D skeleton model imposes a strong prior, addressing the issue of occlusions and acting as a regularization on the over-parameterized joint space.
@@
 For instance, Dai et al.
-(Dai et al., 2023) have introduced an unsupervised approach that, given a 2D pose, employs a pose generator network to generate a corresponding 3D posture, while a separate camera network estimates the camera rotation.
+[19] have introduced an unsupervised approach that, given a 2D pose, employs a pose generator network to generate a corresponding 3D posture, while a separate camera network estimates the camera rotation.
 During training, the generated 3D pose is subjected to random re-projections onto camera viewpoints to synthesize a new 2D pose.
@@
 Complementing this, Sosa et al.
-(Sosa and Hogg, 2023) have proposed a similar methodology which relies on self-supervision with the same geometric consistency constraint for 3D pose estimation.
+[132] have proposed a similar methodology which relies on self-supervision with the same geometric consistency constraint for 3D pose estimation.
 Notably, while Dai et al.
-(Dai et al., 2023) assume the availability of actual 2D poses for each input image, Sosa et al.
-(Sosa and Hogg, 2023) demonstrate the feasibility of their approach with minimal assumptions, including unlabeled images and a small set of synthetic 2D poses.
-The minimalistic requirements and the availability of large volumes of unlabeled data suggest that this method (Sosa and Hogg, 2023) could be readily generalized across diverse animal species.
-Another 2024 approach to automatically extract the 3D pose from single-camera-view images using only synthetic data for training is DigiDogs (Shooter et al., 2024).
-This method takes DINOv2 (Oquab et al., 2023) as the base architecture, which is a foundation model known for its robust pre-training capabilities and superior domain adaptation performance.
+[19] assume the availability of actual 2D poses for each input image, Sosa et al.
+[132] demonstrate the feasibility of their approach with minimal assumptions, including unlabeled images and a small set of synthetic 2D poses.
+The minimalistic requirements and the availability of large volumes of unlabeled data suggest that this method [132] could be readily generalized across diverse animal species.
+Another 2024 approach to automatically extract the 3D pose from single-camera-view images using only synthetic data for training is DigiDogs [128].
+This method takes DINOv2 [105] as the base architecture, which is a foundation model known for its robust pre-training capabilities and superior domain adaptation performance.
 These studies exploit synthetic data to predict 3D joint locations from 2D images, reinforcing the potential of synthetic augmentation in the domain of monocular 3D APE.
 Collectively, these methods in single-camera-view skeleton-only 3D APE underscore the dynamic evolution of the field, reflecting a shift towards more generalized and unsupervised learning paradigms that promise to enhance the versatility and applicability of pose estimation across a broader spectrum of animal species.
-Animal Mesh Recovery
-In the domain of 3D single-camera-view APE from monocular images and videos, this AMR subfield is further categorized into model-based AMR and model-free AMR.
+On the other hand, AMR for 3D APE from monocular images and videos is further categorized into model-based AMR and model-free AMR.
 The former predominantly utilizes parametric body models to infer the detailed 3D mesh of animals.
 The latter eliminates the need for a pre-existing parametric template model to recover 3D animal meshes.
-One of the most well-known model-based AMR is the Skinned Multi-Animal Linear (SMAL) model (Zuffi et al., 2017), which provides a template for various animal classes based on 3D scanning of deformable toy animals.
+One of the most well-known model-based AMR is the Skinned Multi-Animal Linear (SMAL) model [178], which provides a template for various animal classes based on 3D scanning of deformable toy animals.
 The SMAL model represents animals using a linear combination of shape bases, which are adapted to capture the diversity of animal forms, alongside a set of pose parameters to articulate the model.
 This framework serves as a foundation for many model-based AMR works.
-A notable enhancement of the SMAL model is SMALR (Zuffi et al., 2018), which allows the recovery of detailed 3D shapes of unseen animals and creates textured 3D meshes.
-SMALST (Zuffi et al., 2019a) further amalgamates real textures with synthetically generated poses, shapes, and backgrounds.
-Beyond SMALST, the field has seen the development of breed-specific variants of SMAL, such as models for dogs (Rüegg et al., 2023, 2022; Biggs et al., 2020; Li and Lee, 2021a) and birds (Wang et al., 2021a).
+A notable enhancement of the SMAL model is SMALR [177], which allows the recovery of detailed 3D shapes of unseen animals and creates textured 3D meshes.
+SMALST [176] further amalgamates real textures with synthetically generated poses, shapes, and backgrounds.
+Beyond SMALST, the field has seen the development of breed-specific variants of SMAL, such as models for dogs [120, 121, 10, 68] and birds [144].
 These models incorporate scaling parameters into the SMAL framework to enable breed-specific fitting.
@@
 However, when we need to estimate the poses for new animal types without an existing template mesh, data collection for these unseen animals can be time-intensive, thus restricting model-based AMR’s applicability to the diverse animal kingdom.
-On the other hand, model-free approaches (Kanazawa et al., 2018; Goel et al., 2020; Wu et al., 2023a, b; Yang et al., 2022a; Li et al., 2024b; Jakab et al., 2024; Yang et al., 2023b) do not rely on pre-existing parametric body models, which offers greater flexibility when we have rare and diverse animals.
-Many of these methods attempt to learn 3D priors from large-scale internet images and videos in an unsupervised or self-supervised manner, focusing on one category at a time (Goel et al., 2020; Wu et al., 2023a, b; Jakab et al., 2024; Yang et al., 2023b).
-In contrast, 3D-Fauna (2024) (Li et al., 2024b) introduces a pan-category deformable model capable of reconstructing any quadruped animal instance in a feed-forward manner, pushing towards a more generalizable and unconstrained understanding of animal forms from monocular images.
+On the other hand, model-free AMR approaches [54, 36, 153, 147, 154, 148, 74, 50] do not rely on pre-existing parametric body models, which offers greater flexibility when we have rare and diverse animals.
+Many of these methods attempt to learn 3D priors from large-scale internet images and videos in an unsupervised or self-supervised manner, focusing on one category at a time [36, 147, 154, 148, 50].
+In contrast, 3D-Fauna (2024) [74] introduces a pan-category deformable model capable of reconstructing any quadruped animal instance in a feed-forward manner, pushing towards a more generalizable and unconstrained understanding of animal forms from monocular images.
 However, the recovered 3D shapes of current model-free approaches often produce overly smooth surfaces, which do not preserve as much detail as the model-based methods; this severely limits their applicability for downstream applications.
-4.3.2. Multiple Camera Views
+4.3.2 Multiple Camera Views
 Multiple-camera-views 3D APE methods utilize multiple synchronized cameras to capture different perspectives of animals and reconstruct their 3D poses.
-The majority of the existing methods in multiple-camera-views 3D APE are designed to estimate the pose of a single animal in keypoints format (Yao et al., 2019; Günel et al., 2019; Nath et al., 2019; Zhang and Park, 2020; Bala et al., 2020; Zhang et al., 2021; Dunn et al., 2021; Karashchuk et al., 2021; Muramatsu et al., 2022; Ebrahimi et al., 2023).
-In particular, early methods estimate the 2D poses from multiple independent camera views first; then, these 2D predictions are triangulated to 3D poses (LABEL:fig:multiview).
-These post hoc triangulation approaches like (Nath et al., 2019) work well when all the joints are visible in all views.
+(a) An example of multiple camera views 3D APE method [30] that encompasses multi-views approach based on triangulation of 2D keypoint estimations.
+(b) An example of multiple camera views 3D APE method [30] that leverages multi-views geometric information during end-to-end training.
+Figure 5: Multiple camera views 3D animal pose estimation using “DAN-NCE” [30].
+The majority of the existing methods in multiple-camera-views 3D APE are designed to estimate the pose of a single animal in keypoints format [159, 41, 100, 170, 8, 168, 30, 55, 96, 31].
+In particular, early methods estimate the 2D poses from multiple independent camera views first; then, these 2D predictions are triangulated to 3D poses (Figure 5(a)).
+These post hoc triangulation approaches like [100] work well when all the joints are visible in all views.
 However, they are not suited to handle occlusions, which is a scenario that always happens in animal naturalistic environments.
-To address these limitations, some methods apply skeletal constraints and filtering (Günel et al., 2019; Karashchuk et al., 2021).
-In detail, Anipose (Karashchuk et al., 2021) benefits significantly from post-processing spatial-temporal filtering, especially in scenarios involving freely moving subjects where occlusion is frequent.
+To address these limitations, some methods apply skeletal constraints and filtering [41, 55].
+In detail, Anipose [55] benefits significantly from post-processing spatial-temporal filtering, especially in scenarios involving freely moving subjects where occlusion is frequent.
 However, these methods have yet to achieve robust 3D tracking of freely moving animals unless trained with large numbers of frames.
-A notable exception is OpenMonkeyStudio (Bala et al., 2020), which can estimate the 3D macaques pose in extensive, unconstrained environments with a network of 62 high-resolution video cameras.
+A notable exception is OpenMonkeyStudio [8], which can estimate the 3D macaques pose in extensive, unconstrained environments with a network of 62 high-resolution video cameras.
 However, these methods triangulate 3D locations from 2D estimates; their 2D nature limits their ability to capture and quantify behaviours in 3D accurately.
-To incorporate advances in 3D, some approaches integrate multiple-camera-views geometric data directly within end-to-end training pipelines (LABEL:fig:multiview2).
-A notable system, DANNCE (Dunn et al., 2021), utilizes projective geometry to construct a 3D feature space from multiple-camera-views imagery and subsequently trains a 3D CNN to predict the positions of anatomical landmarks from the shared features across cameras.
-In addition, to tackle the data scarcity in 3D APE, 3D-Upper (2023) (Ebrahimi et al., 2023) uses an unsupervised algorithm based on a Statistical Shape Model (SSM) to estimate 3D poses of freely moving animals; this SSM estimator is robust even when many poses (up to 50%) are contaminated by outliers and missing data.
+To incorporate advances in 3D, some approaches integrate multiple-camera-views geometric data directly within end-to-end training pipelines (Figure 5(b)).
+A notable system, DANNCE [30], utilizes projective geometry to construct a 3D feature space from multiple-camera-views imagery and subsequently trains a 3D CNN to predict the positions of anatomical landmarks from the shared features across cameras.
+In addition, to tackle the data scarcity in 3D APE, 3D-Upper (2023) [31] uses an unsupervised algorithm based on a Statistical Shape Model (SSM) to estimate 3D poses of freely moving animals; this SSM estimator is robust even when many poses (up to 50%) are contaminated by outliers and missing data.
 The 3D-Upper APE method does not require labelled training data, which makes it highly scalable and more practical for applications in-the-wild or in situations where manual annotations are impractical.
 All of the aforementioned 3D multiple-camera-views approaches are limited to estimating the pose of a single animal, and they only provide poses in keypoints format, which does not preserve as much animal shape information as mesh-based APE methods.
-MAMMAL (2024) (An et al., 2023) and SBeA (2024) (Han et al., 2024) extend this to enable multi-animal 3D social pose estimation using multi-camera-views.
+MAMMAL [2] and SBeA [43] extend this to enable multi-animal 3D social pose estimation using multi-camera-views.
 In particular, MAMMAL is designed for large mammals and provides a full 3D surface mesh model compared to the sparse 2D keypoint methods.
@@
 This permits the discovery of subtle social behaviour phenotypes that supervised techniques might miss.
-In summary, RGB-based APE methods have advanced significantly from 2017 to 2024.
-Early approaches, like DeepLabCut (Mathis et al., 2018), focused on 2D keypoint-based supervised methods adapted from HPE for tracking animal joints in controlled environments.
+In summary, RGB-based APE methods have advanced significantly from 2011 to 2024.
+Early approaches, like DeepLabCut [90], focused on 2D keypoint-based supervised methods adapted from HPE for tracking animal joints in controlled environments.
 As research progressed, interest grew in 3D APE (particularly mesh-based methods), which capture more detailed pose and shape data for downstream tasks like tracking and behaviour analysis.
@@
 Additionally, given that many animals are social creatures, there is a trend toward multi-animal tracking from video data to understand their communication and social structures.
-These trends culminated in 2024 with the release of SuperAnimal (Shaokai Ye and Mathis, 2024) (2D APE), MAMMAL (An et al., 2023) and SBeA (Han et al., 2024) (3D APE), which offer superior APE performance, generalization across species and diverse environments, reduced dependence on labelled data, and allow for tracking multiple animals simultaneously, setting a new benchmark in the field.
-5. Non-RGB-based Unimodal Animal Pose Estimation
+These trends culminated in 2024 with the release of SuperAnimal [160] (2D APE), MAMMAL [2] and SBeA [43] (3D APE), which offer superior APE performance, generalization across species and diverse environments, reduced dependence on labelled data, and allow for tracking multiple animals simultaneously, setting a new benchmark in the field.
+5 Non-RGB-based Unimodal Animal Pose Estimation
 In the previous section, we comprehensively reviewed APE methods primarily based on RGB cameras.
-However, beyond RGB-based approaches, researchers have explored the use of other unimodal methods for analyzing animal behaviour, table 8 summarises the latest APE methods using both non-RGB unimodal (Yigit et al., 2022) and multi-modal approaches, which is discussed in detailed in section 6.3.
-IMU
-IMU is a small, lightweight, low-cost device that measures an object’s acceleration and angular velocity.
+However, beyond RGB-based approaches, researchers have explored the use of other unimodal methods for analyzing animal behaviour, Table 8 summarises the latest APE methods using both non-RGB unimodal [162] and multi-modal approaches, which is discussed in detailed in Section 6.3.
+IMU: IMU is a small, lightweight, low-cost device that measures an object’s acceleration and angular velocity.
 Attaching multiple IMUs to limbs and the trunk makes it possible to capture the motion and orientation of the animal in 3D space.
-Recent work (Yigit et al., 2022) presents a lameness detection and pose estimation method for horse walk and trot locomotion based on the acceleration and gyroscope measurements from IMUs.
+Recent work [162] presents a lameness detection and pose estimation method for horse walk and trot locomotion based on the acceleration and gyroscope measurements from IMUs.
 Initially, they train the network offline using IMU measurements as input sequences, with labelled data provided by the MoCap system.
@@
 However, using multi-modality approaches, such as combining RGB cameras with other sensors, has gained more popularity, as discussed in Section 6.3.
-Table 8.
-Non-RGB-based unimodal APE methods and multi-modal APE methods.
+Table 8: Non-RGB-based unimodal APE methods and multi-modal APE methods.
 All the symbols and column meanings are defined in Table 5, Table 6 and Table 7, and any additional symbols are defined here: [Uncaptioned image] is GPDM architecture, [Uncaptioned image] is Hierarchical Gaussian Process Latent Variable Model (H-GPLVM), [Uncaptioned image]is Extended Kalman Smoother (EKS), [Uncaptioned image] is Context Models, [Uncaptioned image] is CLIP.
@@
 is an abbreviation for language.
+The method citation is available in column Year.
 Sl Year Methods Sensor Features Evaluated On Metric
 Non-RGB-based Unimodal APE Method: IMU-based
-1 2022 Yigit et al. (Yigit et al., 2022) IMU [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Horse Mean, Std, VR, IAE & IEV
+1 2022 Yigit et al. IMU [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Horse Mean, Std, VR, IAE & IEV
 Multi-Modal APE Methods
-2 2017 Patel et al. (Patel et al., 2017) RGB+GPS+ [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cheetah Qualitative for 3D
-IMU (video observations)
-3 2020 Kearney et al.
-(Kearney et al., 2020a) RGBD [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] RGBD-Dog (Kearney et al., 2020a), PCK, MPJPE, PA-PCK,
-BADJA (Biggs et al., 2018) PCK 3D, PA-PCK 3D,
-PA-MPJPE
-4 2023 Luo et al.
-(Luo et al., 2023) RGBD {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cattles, Pigs (Ruchay et al., 2020) MAE, MRE (Luo et al., 2023)
-5 2023 OptiPose (Patel et al., 2023) RGBD+Ther. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rodent3D (Patel et al., 2023), PCK, MPJPE,
-+Infrared Rat7M (Dunn et al., 2021), AcinoSet (Joska et al., 2021) Reprojection Error
-6 2024 WildPose(Muramatsu et al., 2024) RGB+LiDAR [Uncaptioned image], N/A, [Uncaptioned image], [Uncaptioned image] Eagle, Giraffe, Lion Mean & Std
-7 2022 Sound of Motion (Li et al., 2022) RGB+Acoustic [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse PCK, IoU, PA-MPJPE
-8 2024 CLHOP (Li et al., 2024a) RGB+Acoustic [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image],[Uncaptioned image]} Horse (Li et al., 2024a) PCK@0.1, IoU, PA-MPJPE
-9 2023 CLAMP (Zhang et al., 2023) RGB+Lang. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), AP, AR
-Animal Pose (Cao et al., 2019)
-10 2023 AmadeusGPT (Ye et al., 2023) RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] Mouse & others (Ye et al., 2023) F1 Score
-11 2024 AWOL (Zuffi and Black, 2024) RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] StanfordExtra (Biggs et al., 2020), Perceptual Study,
-3D Trees CLIP score, Qualitative
-6. Multi-Modal Animal Pose Estimation
-Researchers have begun to explore the integration of multi-sensor and multi-modality to improve the accuracy, robustness, and ecological validity of APE algorithms (Patel et al., 2017; Kearney et al., 2020a; Li et al., 2022; Luo et al., 2023; Patel et al., 2023; Zhang et al., 2023; Ye et al., 2023; Muramatsu et al., 2024; Deng, 2023).
+2 2017 Patel et al. RGB+GPS+IMU [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cheetah Qualitative for 3D
+(video observations)
+3 2020 Kearney et al. RGBD [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] BADJA (2018), PCK, MPJPE, PA-PCK,
+RGBD-Dog (2020) PA-MPJPE, PCK 3D,
+PA-PCK 3D
+4 2023 Luo et al. RGBD {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cattles, Pigs (2020) MAE, MRE (2023)
+5 2023 OptiPose RGBD+Ther. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rat7M (2021), PCK, MPJPE,
+Infrared AcinoSet (2021), Reprojection Error
+Infrared Rodent3D (2023)
+6 2024 WildPose RGB+LiDAR [Uncaptioned image], N/A, [Uncaptioned image], [Uncaptioned image] Eagle, Giraffe, Lion Mean & Std
+7 2022 Sound of Motion RGB+Acoustic [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse PCK, IoU, PA-MPJPE
+8 2024 CLHOP RGB+Acoustic [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image],[Uncaptioned image]} Horse (2024) PCK@0.1, IoU, PA-MPJPE
+9 2023 CLAMP RGB+Lang. [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), AP, AR
+AP-10K (2021b)
+10 2023 AmadeusGPT RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] Mouse & others (2023) F1 Score
+11 2024 AWOL RGB+Lang. [Uncaptioned image], [Uncaptioned image], N/A, [Uncaptioned image] StanfordExtra (2020), CLIP score, Qualitative,
+3D Trees Perceptual Study
+6 Multi-Modal Animal Pose Estimation
+Researchers have begun to explore the integration of multi-sensor and multi-modality to improve the accuracy, robustness, and ecological validity of APE algorithms [108, 56, 67, 84, 109, 169, 161, 97, 24].
 In this section, we examine the motivation behind multi-modal systems, review existing multi-modal HPE and APE techniques, discuss how they tackle the limitations in unimodal APE methods, and highlight applications demonstrating the potential of integrating diverse sensors and modalities for APE, as shown in Table 8.
-6.1. Multi-Modal System in Human Pose Estimation
+6.1 Multi-Modal System in Human Pose Estimation
 To inspire the use of multi-modal systems for APE, we draw insights from recent multi-modal HPE studies.
-For instance, (Zheng et al., 2022; Zanfir et al., 2023) combine data from visual RGB cameras with depth information obtained from RGBD and LiDAR sensors in autonomous driving to increase the robustness of 3D HPE in complex environments.
-The mRI dataset (An et al., 2022) integrates mmWave radar, RGBD cameras, and IMUs and provides a benchmark to support 3D HPE sensor fusion research.
-Another study (An et al., 2022) introduces the mRI dataset, a multi-modal dataset for 3D HPE that integrates data from mmWave radar, RGBD cameras, and IMUs.
+For instance, [172, 167] combine data from visual RGB cameras with depth information obtained from RGBD and LiDAR sensors in autonomous driving to increase the robustness of 3D HPE in complex environments.
+The mRI dataset [3] integrates mmWave radar, RGBD cameras, and IMUs and provides a benchmark to support 3D HPE sensor fusion research.
+Another study [3] introduces the mRI dataset, a multi-modal dataset for 3D HPE that integrates data from mmWave radar, RGBD cameras, and IMUs.
 These studies highlight the benefits of multi-modal integration in HPE and further inspire us to explore similar approaches for APE.
-6.2. Motivation for Multi-Modal System in Animal Pose Estimation
-Unimodal APE faces challenges like occlusions and limited perspectives, which multi-modal systems can overcome by combining complementary data to improve accuracy and understanding of animal behaviour (Deng, 2023).
+6.2 Motivation for Multi-Modal System in Animal Pose Estimation
+Unimodal APE faces challenges like occlusions and limited perspectives, which multi-modal systems can overcome by combining complementary data to improve accuracy and understanding of animal behaviour [24].
 For example, multi-modal systems enhance APE resilience by compensating for sensor failures, such as using thermal in low-light conditions when RGB cameras struggle.
@@
 Beyond visual data, integrating IMUs and audio sensors further enriches animal behaviour analysis by capturing motion and auditory cues, which makes multi-modal systems essential for a more robust APE and a better understanding of dynamic environments.
-6.3. Existing Multi-Modal Animal Pose Estimation Techniques
-In recent years, there has been growing interest in the development of multi-modal APE systems (Kearney et al., 2020a; Luo et al., 2023; Patel et al., 2017; Li et al., 2022; Patel et al., 2023; Zhang et al., 2023; Ye et al., 2023; Muramatsu et al., 2024; Li et al., 2024a; Zuffi and Black, 2024).
+6.3 Existing Multi-Modal Animal Pose Estimation Techniques
+In recent years, there has been growing interest in the development of multi-modal APE systems [56, 84, 108, 67, 109, 169, 161, 97, 66, 175].
 It involves combining information from different sensors to derive a unified representation of animal pose.
 Compared to the significant progress in unimodal APE, there is limited research on using multiple sensors and modalities for enhanced accuracy in APE, and we aim to review all of the existing multi-modal APE studies in this section.
-Animal-borne cameras + GPS + IMUs (Ego-view)
-One notable example of a multi-modal system using animal-borne cameras, GPS, and IMU is the work on tracking the cheetah tail (Patel et al., 2017) in ego-view.
+Animal-borne cameras + GPS + IMUs (Ego-view): One notable example of a multi-modal system using animal-borne cameras, GPS, and IMU is the work on tracking the cheetah tail [108] in ego-view.
 Ego-view 3D APE, though less explored, presents a novel perspective by placing cameras on the animals themselves, offering a first-person viewpoint.
-In this study (Patel et al., 2017), researchers attached camera and sensor equipment to cheetahs to obtain motion data from the animal’s viewpoint.
+In this study [108], researchers attached camera and sensor equipment to cheetahs to obtain motion data from the animal’s viewpoint.
 Using kinematic models, they can estimate the 3D pose of the cheetah’s tail, including its position, orientation, and movement in 3D space.
@@
 Please note that the ego-view 3D APE study is currently limited, and the cheetah tail tracking study opens a door for it.
-RGB + Depth
-One promising sensor modality is the RGBD-based pose estimation, as exemplified by (Kearney et al., 2020a).
+RGB + Depth: One promising sensor modality is the RGBD-based pose estimation, as exemplified by [56].
 The pose estimation model is a two-stage approach for predicting 3D canine poses from RGBD data.
@@
 This model can extract dog poses effectively even when there is significant variability in appearance and motion.
-In 2023, there is another RGBD-based APE approach (Luo et al., 2023) that enhances livestock body measurements.
+In 2023, there is another RGBD-based APE approach [84] that enhances livestock body measurements.
 It first preprocesses point cloud data to detect 3D keypoints, then aligns a template mesh (i.e.
@@
 Even though RGBD is a promising approach, its popularity is limited by dataset availability and hardware constraints; hence, it has not yet been widely adopted within the APE community compared to RGB-based methods.
-RGB + Depth + Thermal infrared
-The OptiPose study (Patel et al., 2023) introduces a novel approach for animal pose tracking that leverages a multi-modal dataset and token-based pose optimisation.
+RGB + Depth + Thermal infrared: The OptiPose study [109] introduces a novel approach for animal pose tracking that leverages a multi-modal dataset and token-based pose optimisation.
 To capture accurate APE, researchers collected a 3D multi-modal dataset comprising RGB, depth, and thermal infrared data.
@@
 The study demonstrates the potential of integrating multiple sensor modalities for improved animal pose tracking and motivates further exploration of multi-modal approaches in animal behaviour analysis.
-RGB + LiDAR
-A novel approach to APE uses LiDAR and RGB sensors, as exemplified by the WildPose (Muramatsu et al., 2024) system.
+RGB + LiDAR: A novel approach to APE uses LiDAR and RGB sensors, as exemplified by the WildPose [97] system.
 Unlike traditional methods that rely on RGB images, WildPose utilises a combination of a zoom-lens camera and a solid-state LiDAR to capture both 2D and 3D data of animals in their natural habitats.
@@
 Additionally, the accuracy and quality of the LiDAR data can be affected by environmental factors such as weather conditions and lighting.
-RGB + Acoustic
-The Sound of Motion (Li et al., 2022) and CLHOP (Li et al., 2024a) present a multi-modal approach for estimating horse motion using video and audio data.
+RGB + Acoustic: The Sound of Motion [67] and CLHOP [66] present a multi-modal approach for estimating horse motion using video and audio data.
 Audio features are extracted from sounds like horse hoof strikes and are processed in parallel with visual features.
@@
 These studies highlight the value of incorporating audio to capture fine-grained details of animal movements.
-Apart from APE, audiovisual data have also been applied in other animal motion analyses such as horse behaviour detection (Nunes et al., 2021) and primate action recognition (Bain et al., 2021).
-RGB + Language
-As we advance, the APE trend is moving towards using vision-language-based APE methods (Zhang et al., 2023; Zuffi and Black, 2024).
-The CLAMP (2023) (Zhang et al., 2023) method uses language to enhance learning processes through the description provided, thus increasing the capability of generalising models about new and unseen animal poses.
-Following that, the AWOL (2024) (Zuffi and Black, 2024) approach creates a model that takes input from Contrastive Language-Image Pretraining (CLIP) encodings of either text or images and maps them to 3D shape parameters.
+Apart from APE, audiovisual data have also been applied in other animal motion analyses such as horse behaviour detection [104] and primate action recognition [5].
+RGB + Language: As we advance, the APE trend is moving towards using vision-language-based APE methods [169, 175].
+The CLAMP (2023) [169] method uses language to enhance learning processes through the description provided, thus increasing the capability of generalising models about new and unseen animal poses.
+Following that, the AWOL (2024) [175] approach creates a model that takes input from Contrastive Language-Image Pretraining (CLIP) encodings of either text or images and maps them to 3D shape parameters.
 This learns the mapping between the CLIP latent space and the shape space of SMAL, allowing it to generate realistic 3D animal mesh as a final output.
@@
 In conclusion, the above studies provide insights and inspirations for future work, which shows a trend towards integrating multiple sensors and modalities, such as RGB, depth, IMUs, LiDAR, thermal, acoustic, and language, to enhance the accuracy, robustness, and contextual understanding of APE.
-7. Future Work
-Can we reduce or eliminate the effort required for ground-truth annotation in APE using unsupervised learning and meta-learning?
+7 Future Work
+7.1 Can we reduce or eliminate the effort required for ground-truth annotation in APE using unsupervised learning and meta-learning?
 Unsupervised learning and meta-learning hold great potential to significantly reduce or even eliminate the effort required for ground-truth annotation in APE.
@@
 This hybrid approach would dramatically reduce the need for extensive annotations while improving the model’s robustness and generalisation to new tasks.
-How can we extend multi-modal frameworks to perform multiple-animal pose estimation in the scene?
+7.2 How can we extend multi-modal frameworks to perform multiple-animal pose estimation in the scene?
 Currently, the use of multi-modal frameworks for a single animal is limited, and for multi-animal, minimal work exists (i.e.
-OptiPose (Patel et al., 2023) that can perform APE on two rodents from a multi-modal system).
+OptiPose [109] that can perform APE on two rodents from a multi-modal system).
 However, the future work will include multi-modal multi-animals pose estimation in-the-wild for various species.
@@
 The future will also move towards having real-time processing, which will allow us to track pose and behaviour in real-time, enabling effective health monitoring to protect endangered species.
-Can we perform language-driven animal articulation with advanced APE methods?
+7.3 Can we perform language-driven animal articulation with advanced APE methods?
 With the rise of models like CLIP, the trend is moving towards more vision-language hybrid approaches within the CV community.
@@
 Similar to how text prompts can control image generation in models like CLIP, future APE methods could allow users to articulate an animal’s pose using natural language instructions.
-For instance, commands like ”raise the horse’s front right leg” or ”make the chimpanzee sit” could be translated into precise joint movements or skeletal adjustments in a 3D model.
-Moreover, mesh-based APE methods, when integrated with language models, could allow for real-time adjustments in the animal’s anatomy and surface details with commands like ”increase muscle tension” or ”broaden the horse’s stance” and this could instantly alter both the pose and physical appearance of the 3D mesh model in real-time.
+For instance, commands like “raise the horse’s front right leg” or “make the chimpanzee sit” could be translated into precise joint movements or skeletal adjustments in a 3D model.
+Moreover, mesh-based APE methods, when integrated with language models, could allow for real-time adjustments in the animal’s anatomy and surface details with commands like “increase muscle tension” or “broaden the horse’s stance” and this could instantly alter both the pose and physical appearance of the 3D mesh model in real-time.
 These advancements will be beneficial in applications like virtual reality or animations, where controlling animal movements dynamically with text could simplify workflows.
 It also has the potential for behavioural simulations; for example, researchers or animators can generate lifelike animal behaviours through high-level language inputs without manually defining joint movements.
-Acknowledgements.
-We thank Andrew Zisserman for his excellent suggestions and for reviewing the paper.
-We would also like to thank Jiaxing Zhong, Ta-Ying Cheng, Xinyu Hou, Qian Xie, Shaokai Ye, Daniel Schofield, Eldar Insafutdinov, Tomas Jakab and Anjun Hu for the insightful discussion on pose estimation.
-We would like to thank (Yunus et al., 2024) for providing the symbols.
+Declarations
+We thank Professor Andrew Zisserman for his excellent suggestions and for reviewing the paper.
+We would also like to thank Jiaxing Zhong, Ta-Ying Cheng, Xinyu Hou, Qian Xie, Shaokai Ye, Daniel Schofield, Eldar Insafutdinov, Niki Amini-Naieni, Tomas Jakab and Anjun Hu for insightful technical discussion.
+We would like to thank [166, 20] for using some of their symbols, templates and scripts.
 Additionally, Qianyi would like to thank Jian Zhou for his loving support.